### v1

# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

Abstract.
We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision an ideal remote social connection. Using the Fictional Inquiry (FI) method within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled as “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord). Participants envisioned an immersive 3D platform that would bridge this gap by prioritizing meaningful social connections, enabling time that feels well spent through presence and immersion, natural individual expression, intuitive social navigation that leverages physical-world norms, and playful, low-stakes opportunities for gradual friendship development. We introduce the design framework of spatial integrity, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. The FI method proved effective in generating innovative ideas while empowering youth by fostering a sense of hope and agency over the future of social media through their own design contributions.
social media; fictional inquiry; youth; design
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Human-centered computing Collaborative and social computing
1. Introduction
Most youth today use social media daily and spend significant time on platforms such as Instagram and TikTok (Anderson, 2023). However, these same youth often express deep dissatisfaction with how they spend this time—not because they want less social connection, but because they desire more meaningful online interactions (Kim et al., 2024c). They often feel constrained by the design patterns of archetypal social media platforms, which prioritize uninteresting and attention-grabbing content consumption over genuine connection (Landesman et al., 2024). This disconnect has contributed to broader societal concerns about social media’s impact on youth well-being (Oth, 2023).
Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide: 41 U.S. states suing Meta (Lima-Strong and Nix, 2023), proposals for “warning labels” on social media use (Murthy, 2024), restrictive legislation like the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act” (Oth, [n. d.]), and outright bans on social media use for those under 16 in countries like Australia (Kim, 2024). The EU’s GDPR (gdp, [n. d.]) similarly limits teen access.
However, such alarmist narratives oversimplify a complex reality: the effectiveness of these narratives and policies (as well as the causal role of social media in youth mental health) remains unclear. Despite recognizing the issues raised in these narratives, the youth report more positive than negative personal experiences (Vogels, 2023), even as they acknowledge potential risks to others (Vogels, 2023). Rather than perpetuating fear-based narratives or attempting to completely rebuild existing platforms, researchers and teens themselves emphasize the importance of learning to navigate social media effectively (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025). Furthermore, such alarmist responses and outright restrictions may be misguided—they impede the development of resilience (Wisniewski, 2018b) and induce extreme and unhelpful anxiety and mistrust without fostering constructive actions (Kim et al., 2024a), ultimately leading to lower psychological well-being associated with social media use (Lee and Hancock, 2024).
This misalignment suggests that the solution lies not in the wholesale rejection of social media or attempts to limit youth access, but in expanding our understanding of what social platforms can be. Rather than trying to “fix” existing platforms or promote fear-based narratives, we need to broaden our vision of what social media could be. To explore alternative visions for social technologies that prioritize meaningful connection, we employed the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) in co-design workshops with 23 participants (ages 15-24) experienced in 3D gaming environments (e.g., Minecraft (min, [n. d.]), Roblox (Roblox Corporation, [n. d.])). Participants were asked to envision a social media platform for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This magical context freed participants from current platform constraints while maintaining familiar social dynamics.
Our findings reveal two important patterns in how youth view and experience online social platforms. First, they make a clear distinction in their minds: platforms like Instagram are what they consider “social media”—where they spend significant time but rarely form meaningful connections. In contrast, they do not typically label platforms like Discord or Minecraft as “social media,” even though these are where they actually develop deep friendships and find their time meaningful. Second, because youth associate “social media” primarily with platforms like Instagram and its limitations, they become skeptical of social media’s overall potential to support authentic relationships. This narrow definition of social media, shaped by their experiences with Instagram-like platforms, leads them to dismiss the entire category as inherently flawed.
Participants identified several key elements missing from archetypal social media platforms: (1) emotional connection through presence and immersion, (2) diverse forms of individual expression that feel more natural, (3) intuitive social navigation—ranging from privacy to relationship development—that enables them to leverage norms from the physical world, and (4) playful, low-stakes opportunities to develop friendships. Although participants expressed deep dissatisfaction with the time they currently spend on archetypal social media platforms, they felt that spending the same amount of time on platforms designed to prioritize meaningful connections would feel valuable.
We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media. Additionally, we propose the concept of spatial integrity for social interactions in online spaces, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. These dimensions include spatial presence (the ability to understand the space and identify who is present), spatial composition (the meaning and significance of the space), spatial configuration (the organizational structure of the space), and spatial depth (the extent of information that can be perceived within the space). Furthermore, we demonstrate that the FI method, in the context of social media design, not only prompted creative idea generation and enabled participants to transcend current social media design patterns but also empowered them to envision the positive potential of social media, leaving them feeling more hopeful about its future possibilities.
2. Related Work
2.1. Defining Social Media and its Role in Young People’s Lives
Social media does not yet have a widely agreed-upon definition. There is a range of interpretations, from broad (e.g., “any interactive communication medium that enables two-way interaction and feedback” (Kent, 2010)) to specific (e.g., “a different kind of Internet-based applications which build the ideological and technological foundations of Web 2.0, and allow users to create content and exchange it with other people through the Internet” (Kaplan and Haenlein, 2010)). While some platforms, such as Instagram (All, [n. d.]) or Facebook (fac, [n. d.]), are indisputably considered social media, others, such as YouTube (you, [n. d.]) and Discord (discord page, [n. d.]), are less archetypal. There are also different types of social media, and even such classifications have multiple interpretations (Kaplan and Haenlein, 2010; Sharma and Verma, 2018; Scott, 2015). Some major categories include social networking sites (i.e., “web-based services that allow individuals to (1) construct a public or semi-public profile within a bounded system, (2) articulate a list of other users with whom they share a connection, and (3) view and traverse their list of connections and those made by others within the system” (Boyd and Ellison, 2007)) and blogs (i.e., “a way to share that love with the world and encourage an active community of readers who comment on the posts of the author” (Scott, 2015)).
The range of definitions and classifications of social media reflects the wide array of user needs that social media addresses. From a uses and gratifications perspective (Katz et al., 1973), social media supports a variety of utilities, ranging from social interactions to information seeking and entertainment (Whiting and Williams, 2013). Even within such broad categories of uses and gratifications, there are diverse ways these needs are pursued. For example, within social interactions, there may be a focus on bridging social capital and weaker ties versus bonding social capital and stronger ties (Phua et al., 2017). Further, even on a single platform—Instagram, for instance—users engage in a range of activities, from photo sharing for self-promotion and disclosure (Menon, 2022) to more passive consumption of media for entertainment. In other words, Instagram may serve as a personal branding tool for some while functioning as a way to pass the time for others.
Despite the prevalent narrative that social media is harmful, the experiences of youth with these platforms are more nuanced. They face challenges such as social comparison (Yau and Reich, 2019), significant privacy concerns (Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014), pressures to curate their online presence (Yau and Reich, 2019), and feelings of meaningless time spent (Lukoff et al., 2018), along with more subtle discomforts (Landesman et al., 2024). However, these platforms also offer meaningful benefits. For instance, marginalized populations often derive substantial support from these spaces (Li et al., 2023; Bates et al., 2020), and social media serves as an important medium for identity development and self-expression (Davis, 2012; Lee et al., 2022). Given this complex relationship between youth and social media, researchers advocate for strategies that focus on building resilience and equipping teens to navigate these experiences effectively rather than solely aiming to restrict or mitigate harm (Wisniewski et al., 2012b; Wisniewski, 2018a; Agha et al., 2023).
2.2. Social Technology Design Explorations in HCI
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
At the feature level, efforts to support meaningful communication include designs that encourage effortful communication (Liu et al., 2021; Zhang et al., 2022) and those that foster feelings of genuine connections (Stepanova et al., 2022). While these features are important, social media experiences are often holistic, involving a complex interplay of multiple design aspects and user behaviors. As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
Beyond traditional social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the traditional sense.
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear—especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
2.3. Designing Beyond Doom Narratives
Social media and design fixation. Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
Social media design is no exception. Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps. However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
The Fictional Inquiry (FI) method. FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
According to Dindler and Iversen (Dindler and Iversen, 2007), implementing this method involves three key steps: first, clearly defining the inquiry’s purpose, whether it’s for staging design situations, exploring future ideas, or driving organizational change; second, developing an appropriate narrative that participants are comfortable with while being distinct enough from current practices to encourage new thinking; and third, creating a compelling plot that introduces tension or conflict within the narrative to motivate specific workshop activities. Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
3. Method
We conducted co-design workshops with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the workshop, we used the FI method to scaffold participants in generating design ideas.
3.1. Material: The Narrative Framework
Figure 1. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
[…]
This is a visual chart designed to organize relationships and connections in the Harry Potter universe. There are three columns of labeled sections. In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue). Each house box has its crest. To the left, there are two smaller boxes. The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname. Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name. To the right, there are two more boxes. The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family. Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid. At the bottom, there are three large sections. The leftmost section is labeled “Muggles” with a blank space. The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo. The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes. Each section contains an empty space for names or details to be added.
For our study, we selected the Harry Potter world as the narrative framework for our FI method. The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI workshops. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions—which parallel real-world moves between schools or locations—allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI workshop method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
Following this introduction, we presented the FI narrative: participants were asked to envision ways to support social connection within the Hogwarts world as a non-Wizard (i.e., “Muggle” (mug, [n. d.])) trying to stay connected with everyone across Hogwarts and the Muggle world. We then shared a Miro (mir, [n. d.]) board (Figure 1) where participants could sort their friends and family into categories inspired by the Harry Potter series, such as fictional houses (e.g., Gryffindor, Hufflepuff), teams (e.g., Quiddich Team members), and old friends (e.g., muggles), reflecting fictional houses, teams, and communities from the story. This activity encouraged them to think of their relationships within the Harry Potter context and was intended to further enhance immersion in the fictional design space. Next, we asked general questions about their experiences with social media, focusing on what they found meaningful or meaningless and their vision for an ideal social media platform.
During the design phase, participants were prompted to describe the capabilities their ideal magical social media would have and what aspects of their lives they would choose to share. Throughout this phase, we sought to keep participants immersed in the narrative, referring to existing platforms as “Muggle social media” and maintaining consistent engagement with the Hogwarts setting, as outlined by the fictional inquiry guidelines. Finally, we concluded by asking participants to reflect on how the social media they envisioned for Hogwarts compared to Muggle social media. The semi-structured interview protocol is available in Appendix A.
3.3. Participants and Recruitment
We used a combination of purposive and convenience sampling to recruit participants for our study. The first and second authors shared the recruitment materials (that included a link to a screener survey) on their social media accounts, while the first author also distributed flyers across our institution’s campus.
Participants were required to be between 15 and 24 years old, in accordance with UN (United Nations, [n. d.]) and WHO (who, [n. d.]) definitions for youth, and to consent to be video or audio recorded during a 90-minute interview. Video recording was requested following the FI method guidelines, enabling us to analyze gestures if necessary. We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
Participants were also required to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). This last criterion was based on the assumption that individuals familiar with 3D games would likely have experience with 2D platforms as well, equipping them with a broader set of cognitive tools for creative idea generation. By contrast, extensive experience with 2D platforms alone does not necessarily ensure the ability to conceptualize or engage with spatial technologies effectively. While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
Table 1. Demographics Summary. Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”). Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
A total of 84 individuals completed the screener survey. We reached out to the 48 respondents who met the recruitment criteria, and 23 participants provided consent and completed the interview procedure. The mean age of the participant group was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) identifying as women, 3 (13.0%) as men, 2 (8.7%) as non-binary, and 3 (13.0%) that were gender fluid. The full list of participant information is available in Table 1.
3.4. Data Analysis
We analyzed the interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides a systematic yet adaptable framework for interpreting qualitative data. The method’s independence from pre-existing theoretical frameworks made it particularly suitable for our study, given our study’s exploratory focus on youth-envisioned social media design through the FI method. It enabled us to effectively capture and analyze the diverse perspectives shared by our participants.
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts.
3.5. Ethical Considerations
Given J.K. Rowling’s controversial statements regarding transgender individuals (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in our recruitment material and screener survey and reiterated our position during the interview procedure.
4. Results
Participants articulated frustrations with current social media and envisioned Social Media at Hogwarts (SMH) as a transformative 3D platform prioritizing authentic connections with close friends. They proposed customizable personal spaces, communal areas fostering meaningful interactions, and immersive shared experiences like gaming or walking together. They suggested features that allow users to share expressively, such as memory sharing and real-time updates via avatars or aesthetics alongside low-intensity interactions that offer presence and nuanced privacy. Participants prioritized user control over feed content, fostering closer relationships over unsolicited interactions. Table 2 summarizes this section, while Figure 2 visualizes the collective vision, and Appendix  3 outlines key features per participant.
Figure 2. An AI-generated illustration envisioning a platform whose design reflects a synthesis of participants’ ideas. It is a 3D, open-world social media world with personal spaces—where users mirror their identity and interests—and community areas that support shared activities and memories. These homes exist in a neighborhood-based spatial design, where proximity mirrors social closeness, enabling intentional navigation and serendipitous interactions rather than passive scrolling. The neighborhood also includes themed social rooms and community spaces for public gatherings and interest-based group interactions. Users can visit friends’ homes and explore shared memories and content, similar to a Pensieve, with selective access control over different areas, rooms, and objects, ensuring privacy and customizable levels of sharing. An intelligent pet guide assists in exploration, connection, and engagement, fostering personal expression and meaningful interactions beyond traditional social media.
[…]
A whimsical digital illustration depicting a fantasy-style virtual neighborhood at night, with warm glowing houses, cobblestone pathways, and floating orbs of light. The left side shows a village scene, where small, cozy houses are arranged in a grid, surrounded by gardens, fences, and street lamps. A blue glowing node and arrow highlight one particular house, linking it to the right-side image. The right side provides a closer view inside a warmly lit home, where a young girl in a sweater and dress interacts with a cake on a wooden table. A small holographic memory scene emerges from the cake with a floating, glowing pool of light, depicting a group of people joyfully gathered around the cake. A curious orange dog watches nearby. Shelves filled with books, a spiral staircase, and a glowing lantern enhance the cozy, magical atmosphere. The image suggests an interactive, immersive social media experience, where users can explore virtual homes and relive shared memories.
Table 2. Key Features for Envisioned Social Media from Participant Interviews
Desired Feature Key Idea Example Quote
Immersive Experiences Through Spatial Elements (Section 4.1) “In 2D, you’re just observing… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” (P01)
[Section 4.1.1]
Bringing Realism and Presence Through 3D
Emphasized need for physical presence and embodied interactions beyond flat screens “You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen.” (P06)
[Section 4.1.2]
Personal Spaces and Community-Centered Landscapes
Desired personal homes and neighborhoods that mirror real-world social structures “You each get your own house… anything you want.” (P03)
[Section 4.1.3]
Holographic Display and Virtual Reality
Sought more natural and immersive interfaces beyond traditional devices “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” (P01)
Organic and Intentional Social Interactions (Section 4.2) “The antonym to pressure, stress-free, community-oriented, wholesome, accessible, and accepting.” (P09)
[Section 4.2.1]
Flexible and Safe Spaces for Meaningful Conversations
Valued direct conversations in protected spaces “Actual conversations for real connections, individualized chatting… to get to know people… rather than just posting on a thread.” (P22)
[Section 4.2.2]
Community Gathering Places
Desired spaces for shared activities and casual encounters “a little card game area, shared experiences, something fun” (P06)
[Section 4.2.2] (1)
Activity Spaces for Shared Experiences
Valued collaborative activities and shared projects “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” (P06)
[Section 4.2.2] (2)
(Public) Third Places for Casual Encounters
Desired spaces for serendipitous interactions “You strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” (P03)
[Section 4.2.3]
Ambient Co-Presence
Valued low-intensity connections through shared virtual spaces “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.” (P19)
[Section 4.2.4]
Effortful and Intentional Navigation
Saw physical movement as a meaningful way to engage “Ideally, SMH would have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you.” (P05)
Expressive and Lower-Stakes Sharing (Section 4.3) “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” (P01)
[Section 4.3.1]
Expression Through Personal Area and Avatar Customization
Wanted customizable avatars and environments “decorat[ing] a house or customizing an avatar allows others to get to know more about my personality… through my style.” (P16)
[Section 4.3.2]
Real-Time Sharing: Capturing States and Changes
Desired ways to share real-time states “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that” (P04)
[Section 4.3.3]
Sharing Memories: Multisensory and Emotional Experiences
Wanted to share memories in an immersive and clearer way “jumping into a pool or taking the first bite of an ice cream sundae… from your brain, pull it out, and just be able to share that” (P04)
Granular, Intuitive, and Fun Privacy Mechanisms (Section 4.4)
“If people are posting more like their genuine real life and their experiences, privacy would be a little bit more complicated.”
[Section 4.4.1]
Space-Based Privacy
Wanted privacy boundaries based on spatial metaphors “A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.” (P13)
[Section 4.4.2]
Playful Privacy
Desired engaging privacy mechanisms with elements of fantasy “Secret knock or an object that reveals a secret passage to… secret rooms.” (P11)
[Section 4.4.3]
Contextual Privacy
Wanted context-aware privacy settings “Each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” (P14)
[Section 4.4.4]
Invisibility
Sought robust blocking and visibility control “I don’t wanna be able to tell that the person I blocked is even around me.” (P22)
[Section 4.4.5]
Age-Appropriate Spaces
Emphasized need for age separation “People who are 50 should not be engaging with people who are 13.” (P21)
Refocusing Interaction Priorities for Meaningful Engagement (Section 4.5) “We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page.” (P19)
[Section 4.5.1]
Intentional Content Consumption
Desired more control over content exposure “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor.” (P09)
[Section 4.5.2]
Reduced Celebrity and Commercial Influence
Wanted focus on real friends over celebrities “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” (P07)
[Section 4.5.3]
Spatial Representation of Relationships
Sought physical distance to represent emotional closeness “The top 7 people you’re closest to would be the 7 houses in your cul-de-sac area.” (P15)
[Section 4.5.4]
Intentional and Meaningful Use of Time
Valued active engagement over passive consumption “Anytime not scrolling and instead talking to people feels worthwhile, even if it’s just light conversation.” (P06)
4.1. Immersive Experiences Through Spatial Elements
Participants saw 3D and VR platforms as key to emotionally engaging, socially fulfilling interactions. P16 noted current social media (which we will refer to as “Muggle Social Media” or MSM for short) “does a good job at spreading information [but not] emotions,” arguing that “tight interpersonal interactions” require “talk[ing] to you and see[ing] your face,” which 3D and VR enhance. They envisioned SMH as a 3D open world with realistic, highly flexible social interactions, lifelike avatars, distinct public and personal spaces, and holographic interfaces that convey presence.
4.1.1. Bringing Realism and Presence Through 3D
Many found MSM “flat” or limited by “just a screen,” (P10) yearning for greater realism and freedom. P19 compared SMH to “social media like Minecraft” but broader in scope, supporting a “physical world through 3D graphics” (P19) with “more of a connection” (P06) and “an emphasis on communications.” (P06) P06 referred to the 3D massively multiplayer online game (MMO) Sky (sky, [n. d.]), where players can “talk to other people” or “sit down by a candle and chat” but wanted “more of a connection” than “just sitting there, talking to one random person.”
Navigating shared virtual spaces resonated with P03, who noted that actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” P06 similarly highlighted co-presence in 3D:
“it does feel more real… You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen… you’re sharing the chat room. But you’re not really sharing the space [like when] I’m playing Roblox. [In SMH] I’m sharing that space physically, even though it’s not really physical.” (P06)
4.1.2. Personal Spaces and Community-Centered Landscapes
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world settings like homes, neighborhoods, and shared areas. P19 imagined “separate houses” with “a main house,” where users could “recreat[e] and decorat[e] your house.” P03 echoed this: “You each get your own house… anything you want” (P03). P21 expanded on the idea, seeing personal spaces as “subspaces” within a larger ecosystem.
Having a realistic landscape was essential for fostering connection and community for some. P06 described it as having “neighbors” or an “island next door” in a “town area.” This idea of community and proximity reflected participants’ desire to recreate the social structures of the real world in a digital setting. By mirroring the physical world, the landscape was also presumed to create a sense of place, grounding social interactions in a familiar and relatable context where users feel comfort and ownership. For example, P09 noted traversal in a corporal form would feel “fulfilling,” “calming,” and “reassuring.”
4.1.3. Holographic Display and Virtual Reality
A recurring theme was moving beyond traditional devices like phones and computers to embrace holographic and VR-based interfaces. Many participants felt mobile devices would seem cumbersome in a magical, immersive world. P01 noted, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting holographic screens, like “Jarvis in Iron Man,” or using spell-activated “panels” and “big screen[s].” P15 preferred VR glasses, calling them “more natural,” “easier,” and “more fun.” Other concepts included a “makeup mirror” as a portal, a “watch” that immerses users, or a “mirror glass” for video calls or summoning friends via the charm (i.e., accio).
The immersive nature of VR was seen as fostering deeper connections. P01 noted MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” They concluded 3D VR platforms are “more immersive and more awesome” (P03) than MSM, which is inherently limited “because it’s 2D.”
4.2. Organic and Intentional Social Interactions
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.”
In contrast, SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
4.2.1. Flexible and Safe Spaces for Meaningful Conversations
Many participants valued direct, synchronous conversations as the best way to build closer relationships. P22 wanted, “actual conversations” for “real connections,” preferring “individualized chatting…to get to know people…rather than just posting on a thread.” They envisioned “chat rooms you can walk into,” (P02) each dedicated to specific topics, such as a room for “high school stuff,” decorated with artifacts from shared experiences to spark conversations. These were seen as “protected space…where you can talk…without feeling unsafe or judged” contrasting the perceived lack of safety on MSM, where users “don’t know everything around it.”
P09 imagined a flexible, “Room of Requirement” (roo, [n. d.])-like system for creating customized rooms “on a moment’s notice,” reflecting the “purpose” of conversations. P07 also imagined that personalized chatroom aesthetics could make conversations feel more intimate, such as “photos of you together, inside jokes, or fun memories.” Interest-based spaces, like Discord (discord page, [n. d.]), were also valued for providing a “common point” (P13) for conversation and “a good talking point for introductions.” (P07)
4.2.2. Community Gathering Places
SMH’s envisioned community spaces blended intimate, private interactions with dynamic public areas, fostering shared experiences and casual encounters.
Activity Spaces for Shared Experiences. Participants saw shared activities as central to friendship-building. P06 envisioned “a little card game area” like Roblox, where “shared experiences” of “something fun” create “real memories” to be reminisced, “do you remember that time when…?” They added, “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” P12 valued collaborative “projects online,” such as “creat[ing] a house we love,” offering “a sense of self-achievement” and “teamwork” “instead of just like doom scrolling.” Sharing that experience together “on the same topic, on the same project” makes it “a better way to connect, instead of just… posting.”Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
(Public) Third Places for Casual Encounters. Although most of the interactions participants envisioned centered around friends, participants also emphasized public spaces for serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
4.2.3. Ambient Co-Presence
Participants valued the idea of low-intensity connections through ambient co-presence. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.”
Music would facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They reflected on their experience:
“At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music. And it was one of the most awesome nights because we all knew in that moment that collectively we all enjoyed what we were hearing and it was just good. I don’t know how else to describe it. It was just nice.” (P19)
They added that music could also serve as a “conversation starter” while also providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.2.4. Effortful and Intentional Navigation
Physical travel was seen as a key way to make interactions feel more meaningful and intentional. While teleportation options like “spells” (P06) or the “Marauder’s Map” (P21) were seen as the convenient default, several wanted the option of being able to travel around the virtual world, as if they were taking a “nature walk” (P14). As P21 remarked, “Many times… the best memories happen during a walk.”
For some, the value of their redesigned interactions would come from the intentionality they would bring. P02 described engaging in SMH with deliberate purpose rather than habitual use, preferring a slower pace: “Teleporting would be nice, but not in this situation for me… I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” Similarly, P05 envisioned navigation in SMH as requiring focused effort:
“[Ideally, SMH would] have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you… you would have to walk through a town to get there. It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” (P05)
P14 highlighted how such intentional interactions on SMH would feel more “personal” compared to seeing “a random 2D Insta post that they posted 3 months ago.” Participants also noted that SMH’s intentional nature would amplify SMH’s focus on meaningful connections. P05 speculated, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.”
4.3. Expressive and Lower-Stakes Sharing
Participants emphasized that SMH should enable rich personal expression, allowing them to authentically and casually reflect their personalities and interests through memory-sharing and customization, from decorating dorm rooms to avatars.
4.3.1. Expression Through Personal Area and Avatar Customization
Participants wanted to create and personalize virtual environments and avatars that reflect their identities, preferences, and aspirations. Unlike MSM that only offers “two modes” (P10)— “light and dark mode[s]” (P10)—P21 envisioned a highly customizable world where users could create their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” They emphasized how this space could be shared with friends, “showing a little part of yourself.” P03 imagined a “dark academia” Hogwarts-inspired room filled with “so many bookshelves and… Victorian architecture,” to reflect their interests. P01 expanded on these ideas, envisioning a house with rooms “dedicated to each of your interests.” They added how that would enable them to showcase the different facets of their identities: “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.”
“3D avatar[s]” (P11) and profiles were another key way participants envisioned expressing themselves. P22 highlighted how platforms like Myspace (mys, [n. d.]) allowed for a level of customization that reflected individuality, describing how they enjoyed coding profiles to change “the font and really make it reflect me.” P01 also shared that there is an inherent “joy” in “just that process of creating it, and finding out who you are through that process.” Even just choosing a username was seen as a way to “convey to the world and your friends your own identity… which is a positive feeling usually.” Reflecting on Tumblr, P01 explained how such personalization is “mainly for you,” not just for others to see. P17 felt similarly, stating, “I don’t think it necessarily conveys personal aspects of yourself, but I think it will without you meaning it to.” As they put it, “your brain conjured that up, right? So I think that’s part of you.”
Participants highlighted how expression through choice feels authentic. For example, P01 felt that such “creative expression” could feel “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 described this kind of indirect sharing as “mak[ing] social media an extension of everyday interaction,” allowing people to “gain the basic foremost information” about someone without requiring direct conversations. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
P10 explained that being in a 3D, VR environment amplifies this expressive nature, stating that 2D worlds feel more “compressed” and “able to see as much stuff as you can see in 3D.” P16 noted that in VR, others are “basically right in front of you when they’re talking,” making their “mannerisms… clearly easily observable.” They elaborated that one could learn a lot about someone based on “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.” This experience feels more authentic, as though meeting with “more real people… compared to the profile pictures.”
In addition to expressiveness, participants appreciated how these features support subtle and low-stakes sharing. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, valued the ability to “share things more subtly.” For instance, they explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style.” They found this form of self-expression to be “more justified” than posting, which felt “embarrassing” and “just doing this for other people.”
4.3.2. Real-Time Sharing: Capturing States and Changes
Another recurring theme was the ability to reflect real-time states through external artifacts in virtual spaces. P04 shared how they might opt into sharing updates on “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants envisioned that the blending of physical and virtual selves would allow for a more personalized medium of sharing. P01 imagined their avatar reflecting what outfit they’re wearing in the moment or what song they’re listening to, adding that this could help friends intuitively know “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 described how frequent updates to virtual rooms could reveal changes in personal interests, providing friends with “an insight into… what I’m thinking.”
Participants also discussed the ability to adapt avatars to reflect real-time emotions or actions. P21 described avatars that “read your emotions… showing [them] on your faces,” while P19 envisioned avatars mimicking real-life gestures. Music was frequently mentioned as a subtle and meaningful way to share current moods.
4.3.3. Sharing Memories: Multisensory and Emotional Experiences
A major theme was the ability to share memories directly with others, like the “Pensieve” (pen, [n. d.]). P04 imagined sharing experiences like “jumping into a pool or taking the first bite of an ice cream sundae” directly, creating more of an emotional connection. Sharing memories could also help reduce misunderstandings by enabling more accurate and nuanced expression, unlike MSM, where communication is mostly “just words.” (P03)
P07 analogically described memory-sharing as “screen sharing, but with your eyes.” P08 highlighted the practicality of memory-sharing, explaining how it could capture transient moments—like crossing a state line—without needing to pull out a camera. P03 highlighted how this would enable them to “just give” memories to their mom in real-time, instead of doing so “after the fact,” where they often found themselves missing the right opportunities.
P04 added that memory-sharing would foster more deliberate consumption of content:
“[SMH] feels more personal and interactive… if I could take more time in a less easy-to-consume format with what others choose to share.… On Instagram, when you click on someone’s page, you see all the posts at once. You don’t go one by one, taking time to read through every comment.… Here, you’d have to take more time to hone in on a memory…… It’s the satisfaction of consciously consuming content from someone you’re close to, being deliberate in viewing what they share, versus just quickly looking and being done.” (P04)
Some participants imagined that these memories would not just be sensory but should also include emotions and perspectives. P02 explained how sharing memories could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02) For some, memories weren’t just visual; P01 expressed a desire to “transmit smells through social media,” adding another sensory dimension to shared experiences. Given ethical concerns, many felt that only minimal editing, such as “cropping certain parts you don’t want to share” (P01) or versions “sped-up or slowed-down,” (P07) should be allowed.
4.4. Granular, Intuitive, and Fun Privacy Mechanisms
Participants felt that current social media platforms offer limited privacy options, often lacking flexibility and nuance. P16 noted, “It’s kind of a pain to organize, and… right now in Instagram you can put close friends… that’s about it.” They expressed a need for more granular controls, particularly in a platform like SMH, where users might share more authentic aspects of their lives. As P11 explained, “If people are posting more like their genuine real life and their experiences, [privacy] would be a little bit more complicated,” emphasizing the importance of having “flexibility” to control access.
4.4.1. Space-Based Privacy
Participants envisioned using spatial or visual cues to establish privacy boundaries, drawing on intuitions from physical spaces. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home? If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
P13 elaborated on how privacy norms in physical world contexts transfer to the SMH world, stating:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes—or 30 seconds—the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
While managing visibility across multiple rooms might initially seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could simplify privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.4.2. Playful Privacy
Participants envisioned engaging privacy mechanisms imbued with a sense of playfulness and fantasy. P11 proposed a system where houses or specific objects within personal rooms could incorporate elements like a “secret knock” or an object that reveals a “secret passage to… secret rooms” when interacted with in a specific way known only to an exclusive list of friends. Others suggested knowledge-gated access, such as a “secret handshake” (P21) with their friends. P02 drew inspiration from the Marauder’s Map, envisioning a system where users could only see accessible spaces unlocked by summoning secret “spells.” P05 likened playful privacy mechanisms to “a little Easter egg”, stating “There’s that little kid part of everyone” that makes the features enjoyable. They noted that these elements would feel “fun” and “novel” as opposed to MSM’s privacy settings, which often come across as uninspired or confusing, “like a wall of settings where sometimes the title of what the setting is doesn’t even make sense.”
4.4.3. Contextual Privacy
Participants reflected on the challenges of navigating context collapse on MSM (Marwick and Boyd, 2011), and hoped SMH would make contextual disclosure easier:
“On Muggle social media, you’ve got to create a whole new account. But I’m thinking for this app, it just telepathically knows what’s in Muggle area and what’s in Hogwarts area. So you don’t really have to go through the effort of keeping two lives. You can just walk through a portal. And you’re like a new person.” (P02)
P03 proposed a framework where every user has a “base” identity but could customize additional aspects of their representation. P22 imagined a contextual profile where “any group you’re in…you could change your avatar for that,” drawing inspirations Discord where “you can make different profiles for different servers.” P14 highlighted the importance of individual control, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.”
Participants also emphasized that contextual privacy goes beyond person-based access control. For example, they suggested varying access based on temporal factors. P14 proposed that “people who don’t follow you… can’t see past versions… I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” The idea of gradual openness was another suggestion:
“if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.” (P21)
4.4.4. Invisibility
Participants were critical of existing platforms for their lack of a robust blocking system. P22 noted that on Instagram, even with contact syncing off, “they might still suggest accounts based on the follow relationships,” exposing a Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they would rather avoid. They also emphasized the need to block someone thoroughly, including their “alternate accounts,” that “I don’t wanna be able to tell that the person I blocked is even around me.” P14 added that on Instagram, even private accounts allow “a random person” to see personal content such as Story highlights (Bojkov, 2019).
Many users expressed a desire for greater control over their visibility on social media, emphasizing features that would allow them to be effectively “invisible” unless they chose otherwise. P03 envisioned users having “an invisibility cloak” for situations where “you don’t want anyone to see you.” Similarly, P07 proposed a “super private” profile option where others “can’t even search… like you’re not even on the app.”
The ability to prevent unsolicited contact was another recurring theme. P22 suggested a model inspired by Discord, where “unprompted contact” happens only with people “you’ve already interacted with,” and others “can’t even be discovered unless you have something in common.” This approach resonated with P13, who felt strongly that “people should only be able to add you if you meet in person and exchange a QR code.” They reflected on an experience where they were able to find personal information about a waitress online based just on their name, noting “it’s too easy to find people on social media.”
4.4.5. Age-Appropriate Spaces
Participants emphasized the importance of creating age-appropriate spaces for minors. P21, for example, argued “people who are 50” should not be engaging with “people who are 13.” Age verification was widely seen as a critical first step. P16 suggested incorporating “biometrics” into the verification process, while P19 argued that users should “legally prove” their age. Once reliable age verification is in place, participants advocated for creating separate spaces for minors to which older users do not have access. Conversely, P22 proposed that minors should have restrictions, such as being unable to join general spaces like a “town square.”
4.5. Refocusing Interaction Priorities for Meaningful Engagement
Participants expressed a strong desire to refocus their social media experiences on meaningful connections and intentional interactions. This section explores four key themes that emerged from their responses: the need to prioritize genuine connections over algorithmic content recommendations, reducing the influence of celebrities and commercial content, using spatial metaphors to represent relationship strengths, and fostering more intentional and meaningful use of time. These themes reflect participants’ vision for social media platforms that better align with their desires for authentic social connection and purposeful engagement.
4.5.1. Intentional Content Consumption
Participants emphasized the importance of refocusing social media on genuine, personal connections rather than extraneous content. P19 criticized how platforms like Instagram have shifted away from its initial focus:
“We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page. We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
This sentiment was echoed by P22, who lamented that on Instagram, “half of the timeline is accounts I don’t even follow that were just randomly recommended to me.” They described this experience as “less keeping up with friends and more just getting random stuff shoved in my face.”
On a related note, a major source of frustration was the infinite scrolling design of current platforms. P11 explained, “There’s always more content on social media… this fuels the mindset that… I need to make sure I’m not missing anything.” P14 also criticized the endless content stream on TikTok, which exposes users to “clickbaity” and often harmful content that “doesn’t stimulate [users] at all” and “messes with [users’] attention span.” They expressed a desire to “be more proactive” and “genuinely use [their] brain” instead of “just scrolling mindlessly for an hour.” P09 envisioned a system like an elevator: “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor… Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” This sentiment was echoed by several participants who envisioned brain-controlled interfaces that would “automatically show [users] things [they] would really like” (P07) based on their immediate needs or interests.
4.5.2. Reduced Celebrity and Commercial Influence
Several participants advocated minimizing celebrity influence. P02 stated, “I don’t want celebrity stuff… I don’t want any celebrity drama,” while P01 suggested removing celebrities altogether if the platform prioritized personal enjoyment. If included, they proposed isolating them to a “celebrity zone.” Prioritizing posts from close friends was a key preference. P07 explained, “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH offering greater control over interactions and content, noting it would be “more designed for people you actually know… like your friends, classmates, or Quidditch team,” rather than celebrities or acquaintances. P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
Other participants expressed a need for control over how non-personal content is organized and placed on the platform. P04 suggested being able to “figure out how I want to arrange strangers’ spaces in my… platform.” They also wanted the ability to “prioritize things and feel more in control of how much information is fed to them,” imagining a system where apps and notifications are ranked by personal relevance. P18 saw ads as inevitable but wished they could appear in a separate “billboard” that users could skip.
4.5.3. Spatial Representation of Relationships
Youth envisioned SMH as a platform that reflects the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Participants appreciated this design as it prioritized interactions with closer friends, creating a sense of community and authenticity. P02 highlighted, “The fact that only my close friends have access to it… is the defining factor,” enabling users to focus on “better conversations with other people [that] already know who I am,” without the pressure to “advertise who I am” as they might feel on MSM.
Some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. P03 imagined, “The people around you… would be people from the same house as you,” cultivating a shared sense of “togetherness” and “community building” through common spaces and interactions. Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
This design approach—aligning physical landscapes with emotional closeness—was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.” Unlike MSM feeds, which participants felt prioritized engagement over relationship strength, SMH’s spatial arrangement allowed users to engage naturally with those closest to them. P09 further elaborated on how this physical representation could enhance navigation: “have your friends’ rooms, or like your kind of close friends list, and maybe you’re kind of living in a pod with them, so you can move between the rooms a lot quicker.”
4.5.4. Intentional and Meaningful Use of Time
Participants expressed a strong preference for digital experiences that feel meaningful and intentional, contrasting these with the negative feelings often associated with aimless social media use. P10 described how social media can feel “meaningless” when they “didn’t learn something,” while P18 reflected on the regret they feel after “mindlessly scrolling.” Similarly, P14 shared that scrolling through Instagram Reels while “literally just laying in [their] bed” feels like “wasting [their] time.”
In contrast, participants found interactions involving direct communication and engagement more fulfilling. P16 shared that “anytime not scrolling” and instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this sentiment, noting that social media use feels more meaningful when they are “actually talking to someone.” They explained that such interactions create a sense of “connectivity,” particularly when keeping in touch with people “not in close proximity.”
The distinction between passive and active engagement was another recurring theme. P10 explained that while doomscrolling does not make them feel “fulfilled,” playing games like Minecraft is different because it involves “progress” they can “save and come back to.” Similarly, P14 shared that the lack of connectivity on social media drives their “doomscrolling,” but they believe platforms that encourage interaction could reduce screen time by making time spent online more meaningful. P06 also note that they noted that they “don’t feel bad… most of the time [they] use Roblox,” as it involves engaging activities such as “playing games with friends” and “customizing spaces.”
Given such differences in how time spent is perceived, participants expected time spent on SMH to feel more meaningful than that on MSM. P04 noted that while MSM allows for quicker interpersonal interactions, it often leads to “doomscrolling.” In contrast, SMH would, for example, encourage “thoughtfully engaging with someone’s memories,” leaving them feeling “much more fulfilled.” P05, similarly, explained that they would not feel bad about spending more time on a platform like SMH because it involves “more effort and more time… a more intentional way of interacting.” P07 affirmed this, stating they would be willing to spend more time on a platform like SMH given interactions where they would find interactions “more enjoyable.”
4.6. Reimagining Social Media’s Potential
Given the nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what social media could be. Therefore, while they acknowledged potential challenges—such as the platform being overly immersive, addictive, or demanding—they largely viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants saw SMH as an inspiring alternative to existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They also appreciated the FI workshops for empowering them to reimagine social media and to be hopeful about its possibilities in ways they had previously thought impossible.
4.6.1. Fostering Hope for the Positive Potentials of Social Media
Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become. For example, P06 shared how SMH combines several desirable aspects of social media in a cohesive way, creating an environment focused on meaningful interactions: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there. I kind of like all of those here.”
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “fun” and “novel,” like a “breath of fresh air.”
Beyond features, participants expressed how SMH inspired hope for the future of social media. P10 pointed out that overcoming the design fixation around social media is crucial:
“The way that we can make social media better is if we expand the realm of what the definition of social media is. Instead of it being where someone just shares stuff, it can also be places where people connect, and people do this and that.” (P10)
Several participants noted that the FI workshop played a pivotal role in helping them gain a broader vision of social technologies and envision new possibilities for social media. P15 reflected on how their perspective evolved throughout the process:
“When we started, I literally was giving all these suggestions, which was completely the same thing which existed… I was repeating their concepts. So I think throughout this entire interview, I evolved my idea… actually thinking about what is new… So I think this has been a super hit. I like the entire conversation.” (P15)
They appreciated that SMH was not merely a “clone of the same thing… without actually adding anything else” but rather a concept that was “completely different… not even the same.” Similarly, P14 shared how the workshop shifted their initial skepticism:
“I don’t know what else you could do, because it’s online… so I really don’t know what else there could be to make it more meaningful.” (P14)
After engaging with the workshop, they felt empowered to imagine broader possibilities:
“I’m not really a tech person. So I don’t really think about [how to make social media better] that much. But I feel like this fake scenario… [helped] remind myself, ‘Girl, you have full control, this is no restrictions.’ After I got into that I… have a way bigger vision of what social media can become. (P14)
4.6.2. Celebrating Youthful Creativity and Play
SMH allows a playful and engaging social media experience, allowing users to immerse themselves in actively creative activities and self-expression. P06 appreciated this ability to personalize their experience on SMH, stating, “I think… everyone likes sharing parts of themselves.” P21 likened SMH to a “game” due to its creative freedom. They elaborated, “The creative freedom you have over the platform makes it feel more fictional than it is real.”
Participants found the overall experience of SMH to be youthful, fun, and deeply immersive, offering a space that feels both magical and meaningful. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
“Yeah, [I would prefer the 3D version] actually 100%. I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult. But who cares? I feel like something that’s more enjoyable would be [the SMH].” (P14)
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness. P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
4.6.3. Enabling Otherwise Inaccessible Experiences
SMH creates opportunities for users to engage in experiences and interactions that may otherwise be out of reach. P09 highlighted how individuals “experiencing stress” or those who may not have access to “really, really beautiful landscape and environments” could walk around and “admir[e]” the “calming” landscape of SMH. P16 envisioned SMH as a space where users could engage in activities like climbing a mountain in Nepal, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
P02 saw SMH as a platform that would “seem more like real life, but better.” They imagined being able to engage with friends in idyllic settings, such as “on the beach all the time.” Furthermore, they believed SMH could strengthen friendships through magical features. In their words, “Since people can’t read others’ emotions in real life, SMH would… be a way for you to connect with your friends on a deeper level.” P18 also appreciated how SMH’s 3D interface facilitated them to interact with “family across the world” in an immersive way. SMH would offer a sense of presence that goes beyond conventional video calls, enabling users to engage with loved ones “not just through a screen.”
5. Discussion
Figure 3. (a) Current youth perceptions of social media platforms and where they experience meaningful connections, highlighting the disconnect between platforms commonly recognized as “social media” versus where youth actually experience meaningful social connection. The SMH—youth-envisioned ideal social media—is represented by the overlap, combining both aspects. (b) A more nuanced viewpoint clarifies “social media” as an umbrella term for a range of platforms with different primary purposes—friendship-building, entertainment, status, information-seeking, tie maintenance, and more—rather than a few archetypal platforms.
[…]
This diagram is divided into two parts, offering insights into how youth perceive social media and how they categorize its purposes. On the left, there are two overlapping circles. The circle on the left represents platforms that youth identify as “social media,” such as Instagram, Facebook, and TikTok. The circle on the right illustrates platforms where youth feel meaningfully connected with friends, including FaceTime, Discord, Minecraft, and Roblox. In the overlapping area falls Social Media at Hogwarts (SMH) youth-envisioned ideal social media that combines the qualities of both categories. On the right, a Venn diagram shows potential categories of social media platforms by their primary purposes, such as “Entertainment-Centered,” “Status-Centered,” and “Information-Seeking-Centered.” “Social Media at Hogwarts” is grouped under “Strong Tie-Centered.”
5.1. “Social Media is 2D”: Youth Perspectives on What Constitutes Social Media
Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction. This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways in which they happen to be designed today (Kim et al., 2024b).
Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections. As illustrated in Figure 3(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections with friends (such as Discord or Minecraft). When envisioning their ideal social platform, participants described something that would bridge this divide—a space combining the deliberate pursuit of social connection they associate with traditional social media with the organic, engaged interactions they naturally experience in gaming and chat platforms.
This misalignment between expectations and experiences around social media is problematic because negative perceptions of social media use can diminish psychological well-being (Lee and Hancock, 2024; Kim et al., 2024a). Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce. As shown in Figure 3(b), social media serves diverse uses and gratifications—from information seeking and entertainment to status enhancement and network maintenance. Rather than maintaining an entirely negative view of social media confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations.
For instance, critiquing Instagram for not facilitating authentic, deep friendships misses the point; while such connections might occasionally form there, it is not the platform’s central form of engagement. However, it is important to note that platforms aren’t static—their designs evolve, and user engagement patterns shift over time. Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts. This fluidity in platform use underscores the importance of teaching (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025) youth to be more discerning about the connections they seek and what’s possible on each platform rather than enforcing anxiety-inducing narratives about social media’s harmful effects.
5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings confirm prior studies that while the youth are largely dissatisfied with social media and how they spend their time on archetypal social media platforms, they do not necessarily want to reduce their social media usage—rather, they seek platforms that facilitate deeper, more meaningful connections. Our participants identified four key elements that could transform social media from a space of passive consumption to one of genuine connection:
  1. (1)
    Presence and Immersion. Youth envision platforms that foster genuine emotional connection through real-time presence. While current platforms feel “flat” and emotionally disconnected, participants valued synchronous interactions that create a sense of co-existence. Even simple activities like sharing quick updates about eating ice cream or walking outside, or simply existing in a shared virtual space were seen as meaningful when done with immersive co-presence.
  2. (2)
    Supporting Individual Preferences. Youth desire more natural and diverse ways to express themselves. Participants envisioned platforms where they could show their personality through both active sharing and subtle signals, similar to how personality naturally emerges in physical spaces. From memory-sharing to customizing their virtual spaces and avatars, they want to express themselves spontaneously and casually rather than feeling constrained to carefully curated posts.
  3. (3)
    Intuitive Social Navigation. Meaningful social media experiences require design that mirrors how relationships naturally unfold. Youth value platforms that enable organic relationship growth through proximity and repeated low-intensity interactions, similar to how sitting in the same classroom gradually builds connections. They also seek designs that replicate the subtle, embodied norms of physical spaces—where spatial arrangements naturally signal degrees of privacy and social contexts, much like how a bedroom implies different privacy expectations than a living room.
  4. (4)
    Playful Engagement. Youth emphasize that meaningful connection often grows through joy and shared experiences, yet they find current platforms overly serious and high-stakes. They envision platforms that encourage whimsical exploration and casual interaction—where they can discover hidden objects, explore virtual landscapes, or simply hang out in a “third place.” This playfulness creates low-stakes opportunities for friendship development beyond the performative nature of mainstream social media.
These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams. Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement. While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
Our participants envision environments where relationships deepen naturally through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement—challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
In the design of virtual environments that foster meaningful social connections, spatial cognition—the way we perceive and interpret space—remains critically relevant. Even in virtual spaces, human interactions are shaped by spatial affordances, mirroring many aspects of physical-world interactions. While virtual spaces need not replicate physical environments exactly—indeed, they can transcend physical limitations in some ways while falling short in others—certain spatial principles remain fundamental to how humans build and maintain relationships. This relevance is reflected in our pervasive use of spatial metaphors when describing virtual interactions, such as “chat rooms,” “cyberspace,” or “folders.” While some platforms like Discord and social VR applications successfully implement spatial design elements, mainstream social media platforms often lack these affordances. Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces.
  1. (1)
    Spatial Presence: Understanding the Space and Who’s There. Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others). While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
  2. (2)
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offers only transient content streams. Our participants emphasized the importance of dedicated areas—like communal lounges or study spaces—where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
  3. (3)
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity—a user’s feed” is meant to be personal space yet lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
  4. (4)
    Spatial Depth: How Much Information Can Be Perceived. Spatial depth determines how much information users can perceive in a space. 3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001). Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression. Incorporating spatial depth could enable richer self-expression and more natural interactions.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—are critical for designing virtual environments that support meaningful social connections. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram that we explored in Section 5.2. As spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).)
The concept of spatial integrity offers a framework for understanding why many social media platforms feel disconnected and unsatisfying. By thoughtfully incorporating these principles, we can design online environments that better align with human social-spatial cognition, addressing the limitations our participants identified and better supporting meaningful social interactions and the organic development of relationships.
5.4. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be. Youth have come to equate social media with certain limiting characteristics: endless content streams, superficial interactions, and 2D interfaces focused on broadcasting posts. Interestingly, while they report meaningful social connections and positive experiences in other online social spaces like Minecraft or Discord, they view mainstream social media as inherently tied to emotionally disconnected experiences and superficial interactions. This cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The Hogwarts-inspired context provided fresh metaphors for exploring social spaces, privacy, and connection. Unlike traditional design workshops where youth typically propose incremental improvements to existing platforms, these sessions enabled participants to imagine entirely new possibilities for meaningful connection. More importantly, the workshops helped participants expand their understanding of what social platforms could be, empowering them to articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This study demonstrates how design methods can offer a generative path toward hope in domains where narratives of doom are prevalent.
5.5. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, potentially biasing them toward favorable attitudes toward incorporating spatial elements or online environments that center play. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research could empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him—including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
6. Conclusion
Through co-design workshops employing the Fictional Inquiry (FI) method and leveraging a Hogwarts-inspired narrative, we explored the youth’s vision for an ideal social platform. The interviews revealed a disconnect between mainstream social media and the kinds of meaningful online connections young people seek. While platforms like Instagram have become synonymous with “social media,” youth find their most authentic social interactions happening elsewhere—in gaming environments, chat platforms, and other spaces that they don’t typically seek as destinations for social connection.
The participants collectively emphasized four key elements missing from current social platforms: emotional connection through real-time presence and immersion, natural forms of self-expression beyond carefully curated posts, intuitive social navigation that mirrors physical-world relationship development, and playful low-stakes opportunities for friendship growth. These findings challenge the widespread assumption that engaging youth requires addictive features or endless content streams. Rather, young people express a clear desire for time that feels meaningfully spent, imagining that social media platforms enabling meaningful connection through shared experiences and deliberate engagement would provide exactly that.
Importantly, this research demonstrates how moving beyond standard design patterns through methods like Fictional Inquiry can help youth articulate possibilities they might have previously dismissed as impossible within conventional social media paradigms. Rather than accepting current platforms’ limitations as inevitable, participants were able to envision radically different approaches to online social connection and feel empowered and hopeful about the positive potential of social media that they initially felt pessimistic about.
Further, we introduce spatial integrity as a framework for facilitating meaningful social connections in online environments, helping to understand why many social platforms feel disconnected and unsatisfying while pointing toward design principles that could better support meaningful online relationships. From 2D platforms to spatial technologies, our study highlights an opportunity and hopeful future for creating social experiences that work with, rather than against, our natural needs and intuitions for social connection.
Appendix A FI Workshop Protocol
Introduction
Welcome, and thank you for participating in our design workshop. Before we begin, I want to ensure that you’ve read the consent form I sent earlier. Have you had a chance to review it?
[Response]
Great! Just to reiterate, this session will be recorded for research purposes. Do I have your consent to proceed with the recording?
[Response]
Thank you. If at any point you feel uncomfortable or wish to stop the interview, please let me know.
General
Let’s start with some general questions about your experiences with 3D gaming and Harry Potter.
  • Which 3D games do you play and how long have you been playing them?
  • (Which is your favorite?) Can you describe the 3D elements or the spatial elements in that game a little bit, please?
  • Have you ever played in VR mode if that is available?
  • Any AR games you have played? Like Pokémon Go?
  • What Harry Potter books or movies have you watched?
  • Which is your favorite?
  • Which Hogwarts house would you put yourself in?
Narrative/Plot
As mentioned in the consent form, our goal today is to explore innovative ways to improve social media. We particularly invited 3D game players with a love for Harry Potter because we believe this will help us think creatively and “outside the box” about what an ideal, magical social media could look like, beyond the confines of a small 2D smartphone screen.
Don’t worry about being too creative—we’re here to guide you through the design process, not to evaluate you. We want to explore different design directions with your help, and we’ll be actively involved while also giving you the space to share your ideas.
To give our brainstorming some structure, we’ll start with a specific scenario to solve during this session. I’m going to ask you some guiding questions to help you address the scenario:
Imagine you are a student at Hogwarts. You have friends who are Muggle-born, friends from magical families, and even friends at other wizarding schools around the world. You’re also getting to know me, as I’m a new student you’ve just met at Hogwarts. Keeping in touch with everyone is tough because traditional Muggle social media like Instagram doesn’t really portray your wizard self very well. You don’t have a smartphone in the first place because Hogwarts people don’t need electronic devices, so you can’t really charge your smartphone.
One day, a brilliant Muggle-born student had an idea. They decided to use magic to create a magical social media where everyone—Muggles, Hogwarts students, professors, and friends from other magical schools—could all communicate and share their lives in the most ideal way. Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?
Character Map
Before we dive into the designing part, let’s first map out your relationships in your new life as a Hogwarts student. Use the Miro board link (Figure 1) to write down the names or nicknames of people that you would like to put in each section. For example, you might want to put “younger sister” or “Jane” under Family, some of your coworkers under Quidditch Team members, yourself under the house that you think you belong to, your older middle school friends as Muggles, etc.
Please note the two boundaries in the four rectangles for each house. You can place your closer friends in the inner box and not-so-close ones in the outer box. And you can put the same name in multiple boxes (or parallelograms) as you wish.
  • How do you currently communicate with or stay connected with each of these people via Muggle social media?
  • What are your biggest pet peeves/issues with Muggle social media that you’d like to solve with magical powers?
  • What moments do you feel bad/guilty about using social media, if at all?
  • When does time spent on social media feel meaningless/meaningful/fulfilling, if at all?
  • Pet peeves around privacy settings?
  • When do you feel most connected to your friends when interacting with them on social media?
  • What would a better social media be if you were to use adjectives to describe it? What does “better” mean to you?
Design
Now, let’s think about what this magical world social media might look like. Remember, there are no right or wrong answers here. We’re looking for your imagination and creativity, so feel free to think outside the box and have fun with it!
  • What key capabilities would this thing have for communicating with [GROUP]*?
  • If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)
  • What would you want to share about yourself on this ideal magical “thing” with [GROUP]?
  • If you had [GROUP’] join this “thing,” how might its form adapt to include this new group?
*: Iterate for 1) Romantic Relationships and Closest Friends, 2) Closer Hogwarts Friends, 3) Muggle Friends, 4) Friends at Other Magical Schools, 5) Quidditch Team Members, 6) House Elf
Connecting Back
  • What do you like about this magical “thing” compared to Muggle social media?
  • What do you dislike about it?
  • Imagine using the magical “thing” we discussed earlier. How do you think it could resolve some of the issues people might face with Muggle social media?
  • How might this new magical “thing” better support your needs for self-presentation and identity management compared to traditional social media?
  • What challenges might arise in using such a “thing” at Hogwarts?
Thank you so much for your time and insights. Your contributions are incredibly valuable to our research. Do you have any final thoughts or questions before we conclude?
Appendix B Participant Interview Summary
Table 3. Key Features for Envisioned Social Media from Participant Interviews
PID Envisioned Social Media Features
P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.

### v2

@@
-Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections
+Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry
 Abstract.
-We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision an ideal remote social connection.
-Using the Fictional Inquiry (FI) method within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled as “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord).
-Participants envisioned an immersive 3D platform that would bridge this gap by prioritizing meaningful social connections, enabling time that feels well spent through presence and immersion, natural individual expression, intuitive social navigation that leverages physical-world norms, and playful, low-stakes opportunities for gradual friendship development.
-We introduce the design framework of spatial integrity, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online.
-The FI method proved effective in generating innovative ideas while empowering youth by fostering a sense of hope and agency over the future of social media through their own design contributions.
+We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments.
+Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces.
+Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics.
+We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth.
+Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction.
+Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
 social media; fictional inquiry; youth; design
@@
 1. Introduction
-Most youth today use social media daily and spend significant time on platforms such as Instagram and TikTok (Anderson, 2023).
-However, these same youth often express deep dissatisfaction with how they spend this time—not because they want less social connection, but because they desire more meaningful online interactions (Kim et al., 2024c).
-They often feel constrained by the design patterns of archetypal social media platforms, which prioritize uninteresting and attention-grabbing content consumption over genuine connection (Landesman et al., 2024).
-This disconnect has contributed to broader societal concerns about social media’s impact on youth well-being (Oth, 2023).
-Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023).
-This narrative has prompted restrictive policy responses worldwide: 41 U.S.
-states suing Meta (Lima-Strong and Nix, 2023), proposals for “warning labels” on social media use (Murthy, 2024), restrictive legislation like the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act” (Oth, [n.
-d.]), and outright bans on social media use for those under 16 in countries like Australia (Kim, 2024).
-The EU’s GDPR (gdp, [n.
-d.]) similarly limits teen access.
-However, such alarmist narratives oversimplify a complex reality: the effectiveness of these narratives and policies (as well as the causal role of social media in youth mental health) remains unclear.
-Despite recognizing the issues raised in these narratives, the youth report more positive than negative personal experiences (Vogels, 2023), even as they acknowledge potential risks to others (Vogels, 2023).
-Rather than perpetuating fear-based narratives or attempting to completely rebuild existing platforms, researchers and teens themselves emphasize the importance of learning to navigate social media effectively (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025).
-Furthermore, such alarmist responses and outright restrictions may be misguided—they impede the development of resilience (Wisniewski, 2018b) and induce extreme and unhelpful anxiety and mistrust without fostering constructive actions (Kim et al., 2024a), ultimately leading to lower psychological well-being associated with social media use (Lee and Hancock, 2024).
-This misalignment suggests that the solution lies not in the wholesale rejection of social media or attempts to limit youth access, but in expanding our understanding of what social platforms can be.
-Rather than trying to “fix” existing platforms or promote fear-based narratives, we need to broaden our vision of what social media could be.
-To explore alternative visions for social technologies that prioritize meaningful connection, we employed the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) in co-design workshops with 23 participants (ages 15-24) experienced in 3D gaming environments (e.g., Minecraft (min, [n.
-d.]), Roblox (Roblox Corporation, [n.
-d.])).
-Participants were asked to envision a social media platform for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n.
+Meaningful social connections play a critical role in young people’s development and well-being (Hawkley and Cacioppo, 2010), and many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n.
 d.]).
-This magical context freed participants from current platform constraints while maintaining familiar social dynamics.
-Our findings reveal two important patterns in how youth view and experience online social platforms.
-First, they make a clear distinction in their minds: platforms like Instagram are what they consider “social media”—where they spend significant time but rarely form meaningful connections.
-In contrast, they do not typically label platforms like Discord or Minecraft as “social media,” even though these are where they actually develop deep friendships and find their time meaningful.
-Second, because youth associate “social media” primarily with platforms like Instagram and its limitations, they become skeptical of social media’s overall potential to support authentic relationships.
-This narrow definition of social media, shaped by their experiences with Instagram-like platforms, leads them to dismiss the entire category as inherently flawed.
-Participants identified several key elements missing from archetypal social media platforms: (1) emotional connection through presence and immersion, (2) diverse forms of individual expression that feel more natural, (3) intuitive social navigation—ranging from privacy to relationship development—that enables them to leverage norms from the physical world, and (4) playful, low-stakes opportunities to develop friendships.
-Although participants expressed deep dissatisfaction with the time they currently spend on archetypal social media platforms, they felt that spending the same amount of time on platforms designed to prioritize meaningful connections would feel valuable.
+However, recent trends indicate that social media platforms are increasingly becoming less focused on facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025).
+Initially designed around sharing personal updates with friends, these platforms now often emphasize entertainment, news, and engaging larger, less intimate audiences.
+Meta’s own data illustrates this shift: between 2023 and 2025, the share of time users spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025).
+Mark Zuckerberg described this change as aligned with the company’s focus on supporting users “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
+This shift away from personal interactions raises concerns about the future of meaningful online relationships, particularly for younger users who have grown up with social media as a central part of their social lives.
+Most youth today engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023).
+Despite this high engagement, many young users report dissatisfaction with the quality of their social interactions, not because they wish to reduce their usage, but because they expect these platforms to better support their social connections (Kim et al., 2024b).
+Current platform designs frequently prioritize passive consumption of content over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
+Given the central importance of social connections, particularly for youth, and recognizing social media’s vital role in facilitating these interactions, it is increasingly important to explore new design possibilities.
+Although young people rely heavily on social media, they frequently express dissatisfaction with their current experiences, indicating a significant gap between their social needs and the interactions enabled by existing platforms.
+Current mainstream social media platforms employ highly standardized design patterns, which limit our collective imagination and contribute to feelings of pessimism among both younger and older users.
+Consequently, there is a compelling need for innovative designs that complement existing platforms by placing a stronger emphasis on meaningful interpersonal connections.
+This prompts the critical question of understanding what types of social interactions current designs might be overlooking or inadequately supporting.
+To address this question, we conducted a series of co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007).
+We involved 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n.
+d.]) and Roblox (Roblox Corporation, [n.
+d.]).
+Participants were asked to collaboratively imagine a social media platform specifically designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n.
+d.]).
+This imaginative context allowed participants to step away from the constraints of existing platforms and envision more varied and meaningful social interactions.
+Additionally, because 3D gaming environments are widely recognized as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated that participants could leverage insights from their gaming experiences to propose novel social media interactions.
+Specifically, our research aims to answer the following question: What types of interactions do youth familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
+Our findings reveal that youth with 3D gaming experiences envision ideal social media platforms not as immersive, navigable environments designed to feel more like neighborhoods, community centers, and shared social spaces.
+Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming.
+Together, these visions highlight the youth’s desire for social media that feels more relational, meaningful, and personally rewarding.
 We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media.
@@
 2. Related Work
-2.1. Defining Social Media and its Role in Young People’s Lives
-Social media does not yet have a widely agreed-upon definition.
-There is a range of interpretations, from broad (e.g., “any interactive communication medium that enables two-way interaction and feedback” (Kent, 2010)) to specific (e.g., “a different kind of Internet-based applications which build the ideological and technological foundations of Web 2.0, and allow users to create content and exchange it with other people through the Internet” (Kaplan and Haenlein, 2010)).
-While some platforms, such as Instagram (All, [n.
-d.]) or Facebook (fac, [n.
-d.]), are indisputably considered social media, others, such as YouTube (you, [n.
-d.]) and Discord (discord page, [n.
-d.]), are less archetypal.
-There are also different types of social media, and even such classifications have multiple interpretations (Kaplan and Haenlein, 2010; Sharma and Verma, 2018; Scott, 2015).
-Some major categories include social networking sites (i.e., “web-based services that allow individuals to (1) construct a public or semi-public profile within a bounded system, (2) articulate a list of other users with whom they share a connection, and (3) view and traverse their list of connections and those made by others within the system” (Boyd and Ellison, 2007)) and blogs (i.e., “a way to share that love with the world and encourage an active community of readers who comment on the posts of the author” (Scott, 2015)).
-The range of definitions and classifications of social media reflects the wide array of user needs that social media addresses.
-From a uses and gratifications perspective (Katz et al., 1973), social media supports a variety of utilities, ranging from social interactions to information seeking and entertainment (Whiting and Williams, 2013).
-Even within such broad categories of uses and gratifications, there are diverse ways these needs are pursued.
-For example, within social interactions, there may be a focus on bridging social capital and weaker ties versus bonding social capital and stronger ties (Phua et al., 2017).
-Further, even on a single platform—Instagram, for instance—users engage in a range of activities, from photo sharing for self-promotion and disclosure (Menon, 2022) to more passive consumption of media for entertainment.
-In other words, Instagram may serve as a personal branding tool for some while functioning as a way to pass the time for others.
-Despite the prevalent narrative that social media is harmful, the experiences of youth with these platforms are more nuanced.
-They face challenges such as social comparison (Yau and Reich, 2019), significant privacy concerns (Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014), pressures to curate their online presence (Yau and Reich, 2019), and feelings of meaningless time spent (Lukoff et al., 2018), along with more subtle discomforts (Landesman et al., 2024).
-However, these platforms also offer meaningful benefits.
-For instance, marginalized populations often derive substantial support from these spaces (Li et al., 2023; Bates et al., 2020), and social media serves as an important medium for identity development and self-expression (Davis, 2012; Lee et al., 2022).
-Given this complex relationship between youth and social media, researchers advocate for strategies that focus on building resilience and equipping teens to navigate these experiences effectively rather than solely aiming to restrict or mitigate harm (Wisniewski et al., 2012b; Wisniewski, 2018a; Agha et al., 2023).
+2.1. Meaningful Social Connection in Online Environments
+2.1.1. Defining Meaningful Social Interaction
+Litt et al.
+(Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” While these interactions are described as “deeply subjective,” the primary distinction from non-meaningful ones lies in the presence of impact.
+This impact extends “beyond the immediate interaction itself,” typically involving positive emotions, the exchange of information, or concrete benefits.
+The study identified three key types of impact: “emotional impact” (involving feelings, mood changes, or relationship strengthening), “informational impact” (relating to the exchange of knowledge, advice, or understanding), and “tangible impact” (providing concrete benefits such as goods or services).
+In contrast, non-meaningful interactions were often explicitly described as lacking any such impact.
+Factors like the presence of strong ties (e.g., friends and family) and community ties, as well as participation in shared activities, were found to be associated with meaningfulness.
+However, these appeared to act more as facilitators that enable impact rather than as the primary source of meaning.
+Other attributes quantitatively linked to greater meaningfulness included whether interactions were planned in advance, memorialized (for example, through photos or videos), and synchronous.
+Interestingly, the study found that the communication medium itself, whether online or offline, was less important in predicting meaningfulness after accounting for these other characteristics.
+While in this paper we treat meaningful social interaction as any interaction that participants say they would actively seek out in an ideal scenario.
+reflecting their subjective or perceived meaningfulness, we expected these themes to echo in how our participants talked about their experiences.
+2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
+Social media platforms are often seen as a primary means of connecting with others, especially for young people (Boyd, 2014a).
+However, despite the widespread use of these platforms, many users report feeling lonely or disconnected (Twenge et al., 2019; Primack et al., 2017).
+This disconnect highlights a significant gap between the intended purpose of social media and the actual experiences of users.
+Social connection is just one of the many reasons people turn to social media, but it remains one of the most important.
+From staying in touch with close friends to meeting new people or maintaining community ties, users often come to these platforms seeking connection.
+Mainstream platforms like Instagram (All, [n.
+d.]), YouTube (you, [n.
+d.]), and Discord (discord page, [n.
+d.]) support some of these needs, but they are designed to address a wide range of user motivations, including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022).
+From a uses and gratifications perspective (Katz et al., 1973), even within a single platform, users pursue diverse goals, ranging from building social capital (Phua et al., 2017) to engaging in general media consumption (Menon, 2022).
+As such, although people may approach these platforms with the intention of connecting socially, they may encounter experiences that prioritize other forms of engagement, potentially leaving their social needs unmet.
+This gap highlights the importance of developing a clearer understanding of what it means to support meaningful social connection in online environments.
+While social media offers both risks, such as social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b), and benefits, such as support for marginalized groups (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and opportunities for identity development and self-expression (Davis, 2012; Lee et al., 2022), social connection remains a central motivation for many users.
+Particularly for young people, who increasingly rely on digital platforms for social interaction, it is critical to move beyond treating connection as one of many competing platform goals and toward a more focused understanding of how platforms can meaningfully support users’ social relationships.
 2.2. Social Technology Design Explorations in HCI
+2.2.1. Designing for Meaningful Social Connections
 HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features.
-Social media platforms such as BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression.
+Social media platforms such as BeReal (Kim et al., 2024b), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression.
 This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction.
@@
 As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
-Beyond traditional social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions.
+Beyond archetypal social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions.
 Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments.
 Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions.
-However, these platforms often are not generally recognized as social media in the traditional sense.
-What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear—especially when considering the possibilities offered by emerging technologies such as spatial computing platforms.
+However, these platforms often are not generally recognized as social media in the generic sense.
+What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear, especially when considering the possibilities offered by emerging technologies such as spatial computing platforms.
 Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
-2.3. Designing Beyond Doom Narratives
-Social media and design fixation.
+2.2.2. Video Games as a Source of Meaningful Social Connection
+Video games are increasingly recognized as valuable environments for fostering meaningful social connection, particularly among youth.
+Large-scale survey research shows that video games serve as a major venue for the development and maintenance of friendships, especially for boys (Lenhart, 2015c).
+Teens not only play games with friends they know in person but also frequently make new friends through online gameplay (Lenhart, 2015a), with more than half of teens reporting they have met at least one friend online (Lenhart, 2015a, b).
+These interactions extend beyond gameplay itself, supporting ongoing social relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
+Research also highlights the social and emotional benefits of gaming, including opportunities for cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014).
+Certain game designs promote prosocial outcomes by encouraging players to help and support others, both in-game and beyond (Gentile et al., 2009).
+Moreover, games are not limited to surface-level entertainment; they can also offer deeper, eudaimonic experiences that involve meaning-making, personal growth, and feelings of connection with others (Daneels et al., 2023).
+These findings suggest that video games can be powerful tools for fostering meaningful social connections, particularly among youth, and highlight the potential for game-like features to enhance social media experiences.
+2.3. Recognizing and Overcoming Social Media Design Fixation
 Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design.
@@
 This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
-Social media design is no exception.
-Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps.
-However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)).
+Social media design is no exception (Sundaram, [n.
+d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020).
+Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps (Oth, [n.
+d.]).
+However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024b) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)).
 These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n.
 d.]; Pardes, 2020).
-The Fictional Inquiry (FI) method.
+2.3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
 FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007).
@@
 Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
+Further, one of the most powerful ways to overcome design fixation is analogical transfer (Novick, 1988; Buchanan, 1992).
+We see video games as a rich source of analogies for social media design, offering unique affordances and interaction patterns that can inspire new approaches to fostering meaningful social connections.
+By examining how video games facilitate social interactions, we can identify design patterns that can be adapted to social media contexts, helping to break free from design fixation and explore novel alternatives.
 3. Method
-We conducted co-design workshops with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms.
+We conducted co-design interviews with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms.
 All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n.
 d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers.
-During the workshop, we used the FI method to scaffold participants in generating design ideas.
+During the interviews, we used the FI method to scaffold participants in generating design ideas.
 3.1. Material: The Narrative Framework
@@
 The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
-Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI workshops.
+Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI interviews.
 The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement.
@@
 Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building.
-The characters’ experiences or life transitions—which parallel real-world moves between schools or locations—allow for exploring how social technologies might support relationship maintenance across changing contexts.
+The characters’ experiences or life transitions, which parallel real-world moves between schools or locations, allow for exploring how social technologies might support relationship maintenance across changing contexts.
 The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
 3.2. Procedure
-The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI workshop method.
+The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI method.
 The interviews averaged around 90 minutes.
@@
 We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
-Participants were also required to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox).
-This last criterion was based on the assumption that individuals familiar with 3D games would likely have experience with 2D platforms as well, equipping them with a broader set of cognitive tools for creative idea generation.
-By contrast, extensive experience with 2D platforms alone does not necessarily ensure the ability to conceptualize or engage with spatial technologies effectively.
+They also had to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox).
 While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
@@
 Using this final set of codes, the first author proceeded to code all the transcripts.
+All versions of the codes and the final codebook are available in the supplementary materials.
 3.5. Ethical Considerations
@@
 4. Results
-Participants articulated frustrations with current social media and envisioned Social Media at Hogwarts (SMH) as a transformative 3D platform prioritizing authentic connections with close friends.
-They proposed customizable personal spaces, communal areas fostering meaningful interactions, and immersive shared experiences like gaming or walking together.
-They suggested features that allow users to share expressively, such as memory sharing and real-time updates via avatars or aesthetics alongside low-intensity interactions that offer presence and nuanced privacy.
-Participants prioritized user control over feed content, fostering closer relationships over unsolicited interactions.
-Table 2 summarizes this section, while Figure 2 visualizes the collective vision, and Appendix  3 outlines key features per participant.
-Figure 2.
-An AI-generated illustration envisioning a platform whose design reflects a synthesis of participants’ ideas.
-It is a 3D, open-world social media world with personal spaces—where users mirror their identity and interests—and community areas that support shared activities and memories.
-These homes exist in a neighborhood-based spatial design, where proximity mirrors social closeness, enabling intentional navigation and serendipitous interactions rather than passive scrolling.
-The neighborhood also includes themed social rooms and community spaces for public gatherings and interest-based group interactions.
-Users can visit friends’ homes and explore shared memories and content, similar to a Pensieve, with selective access control over different areas, rooms, and objects, ensuring privacy and customizable levels of sharing.
-An intelligent pet guide assists in exploration, connection, and engagement, fostering personal expression and meaningful interactions beyond traditional social media.
-[…]
-A whimsical digital illustration depicting a fantasy-style virtual neighborhood at night, with warm glowing houses, cobblestone pathways, and floating orbs of light.
-The left side shows a village scene, where small, cozy houses are arranged in a grid, surrounded by gardens, fences, and street lamps.
-A blue glowing node and arrow highlight one particular house, linking it to the right-side image.
-The right side provides a closer view inside a warmly lit home, where a young girl in a sweater and dress interacts with a cake on a wooden table.
-A small holographic memory scene emerges from the cake with a floating, glowing pool of light, depicting a group of people joyfully gathered around the cake.
-A curious orange dog watches nearby.
-Shelves filled with books, a spiral staircase, and a glowing lantern enhance the cozy, magical atmosphere.
-The image suggests an interactive, immersive social media experience, where users can explore virtual homes and relive shared memories.
-Table 2. Key Features for Envisioned Social Media from Participant Interviews
-Desired Feature Key Idea Example Quote
-Immersive Experiences Through Spatial Elements (Section 4.1) “In 2D, you’re just observing… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” (P01)
-[Section 4.1.1]
-Bringing Realism and Presence Through 3D
-Emphasized need for physical presence and embodied interactions beyond flat screens “You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen.” (P06)
-[Section 4.1.2]
-Personal Spaces and Community-Centered Landscapes
-Desired personal homes and neighborhoods that mirror real-world social structures “You each get your own house… anything you want.” (P03)
-[Section 4.1.3]
-Holographic Display and Virtual Reality
-Sought more natural and immersive interfaces beyond traditional devices “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” (P01)
-Organic and Intentional Social Interactions (Section 4.2) “The antonym to pressure, stress-free, community-oriented, wholesome, accessible, and accepting.” (P09)
-[Section 4.2.1]
-Flexible and Safe Spaces for Meaningful Conversations
-Valued direct conversations in protected spaces “Actual conversations for real connections, individualized chatting… to get to know people… rather than just posting on a thread.” (P22)
-[Section 4.2.2]
-Community Gathering Places
-Desired spaces for shared activities and casual encounters “a little card game area, shared experiences, something fun” (P06)
-[Section 4.2.2] (1)
-Activity Spaces for Shared Experiences
-Valued collaborative activities and shared projects “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” (P06)
-[Section 4.2.2] (2)
-(Public) Third Places for Casual Encounters
-Desired spaces for serendipitous interactions “You strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” (P03)
-[Section 4.2.3]
-Ambient Co-Presence
-Valued low-intensity connections through shared virtual spaces “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.” (P19)
-[Section 4.2.4]
-Effortful and Intentional Navigation
-Saw physical movement as a meaningful way to engage “Ideally, SMH would have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you.” (P05)
-Expressive and Lower-Stakes Sharing (Section 4.3) “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” (P01)
-[Section 4.3.1]
-Expression Through Personal Area and Avatar Customization
-Wanted customizable avatars and environments “decorat[ing] a house or customizing an avatar allows others to get to know more about my personality… through my style.” (P16)
-[Section 4.3.2]
-Real-Time Sharing: Capturing States and Changes
-Desired ways to share real-time states “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that” (P04)
-[Section 4.3.3]
-Sharing Memories: Multisensory and Emotional Experiences
-Wanted to share memories in an immersive and clearer way “jumping into a pool or taking the first bite of an ice cream sundae… from your brain, pull it out, and just be able to share that” (P04)
-Granular, Intuitive, and Fun Privacy Mechanisms (Section 4.4)
-“If people are posting more like their genuine real life and their experiences, privacy would be a little bit more complicated.”
-[Section 4.4.1]
-Space-Based Privacy
-Wanted privacy boundaries based on spatial metaphors “A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.” (P13)
-[Section 4.4.2]
-Playful Privacy
-Desired engaging privacy mechanisms with elements of fantasy “Secret knock or an object that reveals a secret passage to… secret rooms.” (P11)
-[Section 4.4.3]
-Contextual Privacy
-Wanted context-aware privacy settings “Each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” (P14)
-[Section 4.4.4]
-Invisibility
-Sought robust blocking and visibility control “I don’t wanna be able to tell that the person I blocked is even around me.” (P22)
-[Section 4.4.5]
-Age-Appropriate Spaces
-Emphasized need for age separation “People who are 50 should not be engaging with people who are 13.” (P21)
-Refocusing Interaction Priorities for Meaningful Engagement (Section 4.5) “We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else.
-We didn’t need the explore page.” (P19)
-[Section 4.5.1]
-Intentional Content Consumption
-Desired more control over content exposure “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor.” (P09)
-[Section 4.5.2]
-Reduced Celebrity and Commercial Influence
-Wanted focus on real friends over celebrities “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” (P07)
-[Section 4.5.3]
-Spatial Representation of Relationships
-Sought physical distance to represent emotional closeness “The top 7 people you’re closest to would be the 7 houses in your cul-de-sac area.” (P15)
-[Section 4.5.4]
-Intentional and Meaningful Use of Time
-Valued active engagement over passive consumption “Anytime not scrolling and instead talking to people feels worthwhile, even if it’s just light conversation.” (P06)
-4.1. Immersive Experiences Through Spatial Elements
-Participants saw 3D and VR platforms as key to emotionally engaging, socially fulfilling interactions.
-P16 noted current social media (which we will refer to as “Muggle Social Media” or MSM for short) “does a good job at spreading information [but not] emotions,” arguing that “tight interpersonal interactions” require “talk[ing] to you and see[ing] your face,” which 3D and VR enhance.
-They envisioned SMH as a 3D open world with realistic, highly flexible social interactions, lifelike avatars, distinct public and personal spaces, and holographic interfaces that convey presence.
-4.1.1. Bringing Realism and Presence Through 3D
-Many found MSM “flat” or limited by “just a screen,” (P10) yearning for greater realism and freedom.
-P19 compared SMH to “social media like Minecraft” but broader in scope, supporting a “physical world through 3D graphics” (P19) with “more of a connection” (P06) and “an emphasis on communications.” (P06) P06 referred to the 3D massively multiplayer online game (MMO) Sky (sky, [n.
-d.]), where players can “talk to other people” or “sit down by a candle and chat” but wanted “more of a connection” than “just sitting there, talking to one random person.”
-Navigating shared virtual spaces resonated with P03, who noted that actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” P06 similarly highlighted co-presence in 3D:
-“it does feel more real… You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen… you’re sharing the chat room.
-But you’re not really sharing the space [like when] I’m playing Roblox.
-[In SMH] I’m sharing that space physically, even though it’s not really physical.” (P06)
-4.1.2. Personal Spaces and Community-Centered Landscapes
-Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world settings like homes, neighborhoods, and shared areas.
-P19 imagined “separate houses” with “a main house,” where users could “recreat[e] and decorat[e] your house.” P03 echoed this: “You each get your own house… anything you want” (P03).
-P21 expanded on the idea, seeing personal spaces as “subspaces” within a larger ecosystem.
-Having a realistic landscape was essential for fostering connection and community for some.
-P06 described it as having “neighbors” or an “island next door” in a “town area.” This idea of community and proximity reflected participants’ desire to recreate the social structures of the real world in a digital setting.
-By mirroring the physical world, the landscape was also presumed to create a sense of place, grounding social interactions in a familiar and relatable context where users feel comfort and ownership.
-For example, P09 noted traversal in a corporal form would feel “fulfilling,” “calming,” and “reassuring.”
-4.1.3. Holographic Display and Virtual Reality
-A recurring theme was moving beyond traditional devices like phones and computers to embrace holographic and VR-based interfaces.
-Many participants felt mobile devices would seem cumbersome in a magical, immersive world.
-P01 noted, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting holographic screens, like “Jarvis in Iron Man,” or using spell-activated “panels” and “big screen[s].” P15 preferred VR glasses, calling them “more natural,” “easier,” and “more fun.” Other concepts included a “makeup mirror” as a portal, a “watch” that immerses users, or a “mirror glass” for video calls or summoning friends via the charm (i.e., accio).
-The immersive nature of VR was seen as fostering deeper connections.
-P01 noted MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” They concluded 3D VR platforms are “more immersive and more awesome” (P03) than MSM, which is inherently limited “because it’s 2D.”
-4.2. Organic and Intentional Social Interactions
-Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14).
-This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23).
-P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.”
-In contrast, SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison.
-P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
-4.2.1. Flexible and Safe Spaces for Meaningful Conversations
-Many participants valued direct, synchronous conversations as the best way to build closer relationships.
-P22 wanted, “actual conversations” for “real connections,” preferring “individualized chatting…to get to know people…rather than just posting on a thread.” They envisioned “chat rooms you can walk into,” (P02) each dedicated to specific topics, such as a room for “high school stuff,” decorated with artifacts from shared experiences to spark conversations.
-These were seen as “protected space…where you can talk…without feeling unsafe or judged” contrasting the perceived lack of safety on MSM, where users “don’t know everything around it.”
-P09 imagined a flexible, “Room of Requirement” (roo, [n.
-d.])-like system for creating customized rooms “on a moment’s notice,” reflecting the “purpose” of conversations.
-P07 also imagined that personalized chatroom aesthetics could make conversations feel more intimate, such as “photos of you together, inside jokes, or fun memories.” Interest-based spaces, like Discord (discord page, [n.
-d.]), were also valued for providing a “common point” (P13) for conversation and “a good talking point for introductions.” (P07)
-4.2.2. Community Gathering Places
-SMH’s envisioned community spaces blended intimate, private interactions with dynamic public areas, fostering shared experiences and casual encounters.
-Activity Spaces for Shared Experiences.
-Participants saw shared activities as central to friendship-building.
-P06 envisioned “a little card game area” like Roblox, where “shared experiences” of “something fun” create “real memories” to be reminisced, “do you remember that time when…?” They added, “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” P12 valued collaborative “projects online,” such as “creat[ing] a house we love,” offering “a sense of self-achievement” and “teamwork” “instead of just like doom scrolling.” Sharing that experience together “on the same topic, on the same project” makes it “a better way to connect, instead of just… posting.”Even “mindless scrolling” became meaningful when shared.
-P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
-(Public) Third Places for Casual Encounters.
-Although most of the interactions participants envisioned centered around friends, participants also emphasized public spaces for serendipitous interactions.
-P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look!
-This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
-4.2.3. Ambient Co-Presence
-Participants valued the idea of low-intensity connections through ambient co-presence.
+Informed by their experiences with 3D games and familiar media worlds, our participants imagined forms of social interaction that extend beyond what is typically supported on mainstream platforms like Instagram.
+Through the speculative frame of Social Media at Hogwarts (SMH), they explored how digital environments might enable more embodied, relational, and emotionally meaningful forms of connection.
+Their visions converged around six interrelated themes, which, collectively, contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs, ultimately expanding their perception of the horizons for social connection on such platforms.
+4.1. Intuitive Social Navigation
+Youth described seeking interactions that feel natural, embodied, and socially meaningful, reflecting the way people navigate relationships and spaces in everyday life.
+They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape how they connect with others, express themselves, and manage privacy in intuitive and socially familiar ways.
+4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
+Participants described how embodied, immersive qualities such as realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space that fosters presence, emotional connection, and socially meaningful interaction.
+In contrast, current social media (which we will refer to as “Muggle Social Media” or MSM for short), was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 similarly noted that MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” P03 described SMH as “more immersive and more awesome” than MSM, which they saw as inherently limited “because it’s 2D.” P16 emphasized that meaningful social connection requires “talk[ing] to you and see[ing] your face,” something they believed SMH could make possible.
+For others, the fulfillment came from the embodied experience of physical actions.
+P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They also added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
+P03 similarly reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” As P06 put it, SMH “feels more real” because you “walk around a physical world” and “feel like you’re actually in the area of these people.”
+This sense of presence was felt even in seemingly passive actions. As P01 explained,
+“I don’t know, there’s just a feeling about it that makes it different.
+I can’t really explain.
+Depending on 2D and 3D, but I feel like in 2D, you can’t really change anything about their homepage.
+You just observe it.
+The most you can do is add a like or something, and that would change their like count.
+But then they would know who you are, so that takes away the anonymity.
+But if you’re just looking at it, then you’re not really affecting anything.
+That’s just for your knowledge.
+But in a 3D setting, even if you don’t leave any physical footprints, in a way, they’re still there because you opened the door, you went in, you sat down.
+Maybe just that mark in a way that you’ve left on the place takes away from the anonymity.
+I don’t know if that makes sense but.” (P01)
+For P01, 3D interaction not only supported more embodied engagement but also introduced a felt presence that lingered, even when no direct interaction took place.
+4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
+Participants proposed that using spatial and visual cues modeled after real-world environments could make managing access and privacy feel more intuitive, socially natural, and contextually appropriate.
+P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
+P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
+“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home?
+If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
+While managing visibility across multiple rooms might seem more complex, participants believed this approach would feel intuitive and natural.
+P11 suggested that spatial contexts could lower the efforts required for privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
+4.2. Shared Experiences
+Participants emphasized that shared experiences, whether through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
+4.2.1. Fostering Shared Memories Through Collaborative Activities
+Participants described shared activities as a meaningful foundation for building friendships.
+P06 contrasted casual solo games with more immersive, co-experienced ones.
+For example, while playing a simple emoji game was described as “just something fun we can both do,” it lacked a true shared experience since “we don’t actually see each other’s screens” and would just “screenshot our scores and send them to each other.” By contrast, they described Roblox as a space that fosters shared moments that become lasting memories, like “do you remember that time when that kid in the voice chat said this,” or “when we got scared by that jump scare.” They explained, “it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
+Similarly, P12 described the value of collaborative projects, such as “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting it with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than simply “posting” content individually.
+The value of shared experiences and memories was not limited to completing tasks or reaching goals.
+Even simple activities like walking together were seen as meaningful opportunities for connection.
+P03 imagined moments where you could just “walk with your friend,” as if taking a “nature walk” (P14).
+P21 echoed this, describing how walking itself could spark memorable conversations and shared moments: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
+4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
+Participants also saw low-pressure, ambient forms of social interaction, such as co-existing in shared spaces or listening to music together, as natural and organic ways to build connections and shared memories.
 P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background.
 They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations.
-P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.”
-Music would facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They reflected on their experience:
+P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.” Even “mindless scrolling” became meaningful when shared.
+P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
+Music would often facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They came up with the idea reflecting on their own experience:
 “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.
@@
 You can just listen to the music.”
-4.2.4. Effortful and Intentional Navigation
-Physical travel was seen as a key way to make interactions feel more meaningful and intentional.
-While teleportation options like “spells” (P06) or the “Marauder’s Map” (P21) were seen as the convenient default, several wanted the option of being able to travel around the virtual world, as if they were taking a “nature walk” (P14).
-As P21 remarked, “Many times… the best memories happen during a walk.”
-For some, the value of their redesigned interactions would come from the intentionality they would bring.
-P02 described engaging in SMH with deliberate purpose rather than habitual use, preferring a slower pace: “Teleporting would be nice, but not in this situation for me… I wouldn’t really want anything to be super fast decisions.
-I want everything to be very thought through.” Similarly, P05 envisioned navigation in SMH as requiring focused effort:
-“[Ideally, SMH would] have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you… you would have to walk through a town to get there.
-It’s not challenging to navigate, but you have to intentionally go there to see it.
-It isn’t something you can do mindlessly.” (P05)
-P14 highlighted how such intentional interactions on SMH would feel more “personal” compared to seeing “a random 2D Insta post that they posted 3 months ago.” Participants also noted that SMH’s intentional nature would amplify SMH’s focus on meaningful connections.
-P05 speculated, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
-For others, the fulfillment came from the embodied experience of physical actions.
-P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.”
-4.3. Expressive and Lower-Stakes Sharing
-Participants emphasized that SMH should enable rich personal expression, allowing them to authentically and casually reflect their personalities and interests through memory-sharing and customization, from decorating dorm rooms to avatars.
-4.3.1. Expression Through Personal Area and Avatar Customization
-Participants wanted to create and personalize virtual environments and avatars that reflect their identities, preferences, and aspirations.
-Unlike MSM that only offers “two modes” (P10)— “light and dark mode[s]” (P10)—P21 envisioned a highly customizable world where users could create their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” They emphasized how this space could be shared with friends, “showing a little part of yourself.” P03 imagined a “dark academia” Hogwarts-inspired room filled with “so many bookshelves and… Victorian architecture,” to reflect their interests.
-P01 expanded on these ideas, envisioning a house with rooms “dedicated to each of your interests.” They added how that would enable them to showcase the different facets of their identities: “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.”
-“3D avatar[s]” (P11) and profiles were another key way participants envisioned expressing themselves.
-P22 highlighted how platforms like Myspace (mys, [n.
-d.]) allowed for a level of customization that reflected individuality, describing how they enjoyed coding profiles to change “the font and really make it reflect me.” P01 also shared that there is an inherent “joy” in “just that process of creating it, and finding out who you are through that process.” Even just choosing a username was seen as a way to “convey to the world and your friends your own identity… which is a positive feeling usually.” Reflecting on Tumblr, P01 explained how such personalization is “mainly for you,” not just for others to see.
-P17 felt similarly, stating, “I don’t think it necessarily conveys personal aspects of yourself, but I think it will without you meaning it to.” As they put it, “your brain conjured that up, right?
-So I think that’s part of you.”
-Participants highlighted how expression through choice feels authentic.
-For example, P01 felt that such “creative expression” could feel “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 described this kind of indirect sharing as “mak[ing] social media an extension of everyday interaction,” allowing people to “gain the basic foremost information” about someone without requiring direct conversations.
-P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
-P10 explained that being in a 3D, VR environment amplifies this expressive nature, stating that 2D worlds feel more “compressed” and “able to see as much stuff as you can see in 3D.” P16 noted that in VR, others are “basically right in front of you when they’re talking,” making their “mannerisms… clearly easily observable.” They elaborated that one could learn a lot about someone based on “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.” This experience feels more authentic, as though meeting with “more real people… compared to the profile pictures.”
-In addition to expressiveness, participants appreciated how these features support subtle and low-stakes sharing.
-P16, who described themselves as “fairly private” and not particularly “public” about their preferences, valued the ability to “share things more subtly.” For instance, they explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style.” They found this form of self-expression to be “more justified” than posting, which felt “embarrassing” and “just doing this for other people.”
-4.3.2. Real-Time Sharing: Capturing States and Changes
-Another recurring theme was the ability to reflect real-time states through external artifacts in virtual spaces.
-P04 shared how they might opt into sharing updates on “what I’m doing at the moment.
-For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
-Participants envisioned that the blending of physical and virtual selves would allow for a more personalized medium of sharing.
-P01 imagined their avatar reflecting what outfit they’re wearing in the moment or what song they’re listening to, adding that this could help friends intuitively know “‘Oh.
-that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 described how frequent updates to virtual rooms could reveal changes in personal interests, providing friends with “an insight into… what I’m thinking.”
-Participants also discussed the ability to adapt avatars to reflect real-time emotions or actions.
-P21 described avatars that “read your emotions… showing [them] on your faces,” while P19 envisioned avatars mimicking real-life gestures.
-Music was frequently mentioned as a subtle and meaningful way to share current moods.
-4.3.3. Sharing Memories: Multisensory and Emotional Experiences
-A major theme was the ability to share memories directly with others, like the “Pensieve” (pen, [n.
-d.]).
-P04 imagined sharing experiences like “jumping into a pool or taking the first bite of an ice cream sundae” directly, creating more of an emotional connection.
-Sharing memories could also help reduce misunderstandings by enabling more accurate and nuanced expression, unlike MSM, where communication is mostly “just words.” (P03)
-P07 analogically described memory-sharing as “screen sharing, but with your eyes.” P08 highlighted the practicality of memory-sharing, explaining how it could capture transient moments—like crossing a state line—without needing to pull out a camera.
-P03 highlighted how this would enable them to “just give” memories to their mom in real-time, instead of doing so “after the fact,” where they often found themselves missing the right opportunities.
-P04 added that memory-sharing would foster more deliberate consumption of content:
-“[SMH] feels more personal and interactive… if I could take more time in a less easy-to-consume format with what others choose to share.… On Instagram, when you click on someone’s page, you see all the posts at once.
-You don’t go one by one, taking time to read through every comment.… Here, you’d have to take more time to hone in on a memory…… It’s the satisfaction of consciously consuming content from someone you’re close to, being deliberate in viewing what they share, versus just quickly looking and being done.” (P04)
-Some participants imagined that these memories would not just be sensory but should also include emotions and perspectives.
-P02 explained how sharing memories could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02) For some, memories weren’t just visual; P01 expressed a desire to “transmit smells through social media,” adding another sensory dimension to shared experiences.
-Given ethical concerns, many felt that only minimal editing, such as “cropping certain parts you don’t want to share” (P01) or versions “sped-up or slowed-down,” (P07) should be allowed.
-4.4. Granular, Intuitive, and Fun Privacy Mechanisms
-Participants felt that current social media platforms offer limited privacy options, often lacking flexibility and nuance.
-P16 noted, “It’s kind of a pain to organize, and… right now in Instagram you can put close friends… that’s about it.” They expressed a need for more granular controls, particularly in a platform like SMH, where users might share more authentic aspects of their lives.
-As P11 explained, “If people are posting more like their genuine real life and their experiences, [privacy] would be a little bit more complicated,” emphasizing the importance of having “flexibility” to control access.
-4.4.1. Space-Based Privacy
-Participants envisioned using spatial or visual cues to establish privacy boundaries, drawing on intuitions from physical spaces.
-P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
-P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
-“I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home?
-If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
-P13 elaborated on how privacy norms in physical world contexts transfer to the SMH world, stating:
-“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.
-It’s where you spend most of your time, and it shows your personality the most.
-It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes—or 30 seconds—the way you can on Instagram.
-If you wanted to add someone there, all you’d have to do is remove a couple of things.
-It’s easier than physically changing a space.” (P13)
-While managing visibility across multiple rooms might initially seem more complex, participants believed this approach would feel intuitive and natural.
-P11 suggested that spatial contexts could simplify privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
-4.4.2. Playful Privacy
-Participants envisioned engaging privacy mechanisms imbued with a sense of playfulness and fantasy.
-P11 proposed a system where houses or specific objects within personal rooms could incorporate elements like a “secret knock” or an object that reveals a “secret passage to… secret rooms” when interacted with in a specific way known only to an exclusive list of friends.
-Others suggested knowledge-gated access, such as a “secret handshake” (P21) with their friends.
-P02 drew inspiration from the Marauder’s Map, envisioning a system where users could only see accessible spaces unlocked by summoning secret “spells.” P05 likened playful privacy mechanisms to “a little Easter egg”, stating “There’s that little kid part of everyone” that makes the features enjoyable.
-They noted that these elements would feel “fun” and “novel” as opposed to MSM’s privacy settings, which often come across as uninspired or confusing, “like a wall of settings where sometimes the title of what the setting is doesn’t even make sense.”
-4.4.3. Contextual Privacy
-Participants reflected on the challenges of navigating context collapse on MSM (Marwick and Boyd, 2011), and hoped SMH would make contextual disclosure easier:
-“On Muggle social media, you’ve got to create a whole new account.
-But I’m thinking for this app, it just telepathically knows what’s in Muggle area and what’s in Hogwarts area.
-So you don’t really have to go through the effort of keeping two lives.
-You can just walk through a portal.
-And you’re like a new person.” (P02)
-P03 proposed a framework where every user has a “base” identity but could customize additional aspects of their representation.
-P22 imagined a contextual profile where “any group you’re in…you could change your avatar for that,” drawing inspirations Discord where “you can make different profiles for different servers.” P14 highlighted the importance of individual control, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.”
-Participants also emphasized that contextual privacy goes beyond person-based access control.
-For example, they suggested varying access based on temporal factors.
-P14 proposed that “people who don’t follow you… can’t see past versions… I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” The idea of gradual openness was another suggestion:
-“if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on.
-Or turn it off.” (P21)
-4.4.4. Invisibility
-Participants were critical of existing platforms for their lack of a robust blocking system.
-P22 noted that on Instagram, even with contact syncing off, “they might still suggest accounts based on the follow relationships,” exposing a Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they would rather avoid.
-They also emphasized the need to block someone thoroughly, including their “alternate accounts,” that “I don’t wanna be able to tell that the person I blocked is even around me.” P14 added that on Instagram, even private accounts allow “a random person” to see personal content such as Story highlights (Bojkov, 2019).
-Many users expressed a desire for greater control over their visibility on social media, emphasizing features that would allow them to be effectively “invisible” unless they chose otherwise.
-P03 envisioned users having “an invisibility cloak” for situations where “you don’t want anyone to see you.” Similarly, P07 proposed a “super private” profile option where others “can’t even search… like you’re not even on the app.”
-The ability to prevent unsolicited contact was another recurring theme.
-P22 suggested a model inspired by Discord, where “unprompted contact” happens only with people “you’ve already interacted with,” and others “can’t even be discovered unless you have something in common.” This approach resonated with P13, who felt strongly that “people should only be able to add you if you meet in person and exchange a QR code.” They reflected on an experience where they were able to find personal information about a waitress online based just on their name, noting “it’s too easy to find people on social media.”
-4.4.5. Age-Appropriate Spaces
-Participants emphasized the importance of creating age-appropriate spaces for minors.
-P21, for example, argued “people who are 50” should not be engaging with “people who are 13.” Age verification was widely seen as a critical first step.
-P16 suggested incorporating “biometrics” into the verification process, while P19 argued that users should “legally prove” their age.
-Once reliable age verification is in place, participants advocated for creating separate spaces for minors to which older users do not have access.
-Conversely, P22 proposed that minors should have restrictions, such as being unable to join general spaces like a “town square.”
-4.5. Refocusing Interaction Priorities for Meaningful Engagement
-Participants expressed a strong desire to refocus their social media experiences on meaningful connections and intentional interactions.
-This section explores four key themes that emerged from their responses: the need to prioritize genuine connections over algorithmic content recommendations, reducing the influence of celebrities and commercial content, using spatial metaphors to represent relationship strengths, and fostering more intentional and meaningful use of time.
-These themes reflect participants’ vision for social media platforms that better align with their desires for authentic social connection and purposeful engagement.
-4.5.1. Intentional Content Consumption
-Participants emphasized the importance of refocusing social media on genuine, personal connections rather than extraneous content.
-P19 criticized how platforms like Instagram have shifted away from its initial focus:
+4.3. Communal Ambience
+SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison.
+P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
+Participants described several design features, such as neighborhood structures, shared gathering spaces, and stronger boundary controls, that could help cultivate this kind of communal, low-pressure environment.
+4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
+Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to naturally foster social interaction and relationship-building.
+P22 described in detail:
+“So, in a virtual world, you might have your house, and then you could go to a community board or town square with postings for different groups … From there, you could travel to those group spaces.
+I don’t know if you’ve ever played or seen Club Penguin, but in that world, you had your igloo to decorate and could visit the town … The town had different buildings for different activities.
+I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
+Youth envisioned SMH as also reflecting the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties.
+Specifically, some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction.
+One participant elaborated:
+“And in this magical world, I imagine the people around you in your neighborhood would be from the same house as you, kind of like how people live together in dorms on a college campus.
+The nearby stores would be the ones your group usually goes to.
+So you’d share that sense of togetherness because you all go to the same stores or belong to the same house.
+I think that would help build a real sense of community.” (P03)
+Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
+This design approach of aligning physical landscapes with emotional closeness was seen as a way to streamline interactions and focus on meaningful relationships.
+P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
+4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
+Participants imagined that in addition to private homes and personal spaces, SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating opportunities for connection.
+P05 described these areas as “neutral spaces”, not tied to any specific activity or relationship status, where groups could gather for a range of purposes.
+They imagined “some places to sit, and some kind of flat table-like surface” that could support casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces, they explained, would function as “a third, neutral territory” where people could interact more organically, without the friction of managing invitations or switching platforms.
+Participants described shared social spaces within neighborhoods, such as common rooms, courtyards, or clubhouses, where residents could casually interact outside their personal spaces.
+P01 described leaving one’s home to gather in a “courtyard or game room” at the center of the neighborhood.
+Drawing on examples from Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms that cater to different interests, such as a Quidditch field or library, creating opportunities for people to meet others through shared activities.
+P09 similarly imagined a “virtual dining room” that would allow people to casually talk while doing something familiar, like eating together.
+Unlike the more segmented house-based accommodation at Hogwarts, this dining space was described as open and inclusive, allowing people to connect across social groups.
+Others described third places located in public areas that supported casual, serendipitous interactions.
+P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look!
+This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
+4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
+To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized the importance of filtering out irrelevant content, such as celebrity drama or political commentary, and prioritizing updates from close friends and trusted communities.
+P19 criticized how platforms like Instagram have shifted away from their initial focus:
 “We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else.
@@
 We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
-This sentiment was echoed by P22, who lamented that on Instagram, “half of the timeline is accounts I don’t even follow that were just randomly recommended to me.” They described this experience as “less keeping up with friends and more just getting random stuff shoved in my face.”
-On a related note, a major source of frustration was the infinite scrolling design of current platforms.
-P11 explained, “There’s always more content on social media… this fuels the mindset that… I need to make sure I’m not missing anything.” P14 also criticized the endless content stream on TikTok, which exposes users to “clickbaity” and often harmful content that “doesn’t stimulate [users] at all” and “messes with [users’] attention span.” They expressed a desire to “be more proactive” and “genuinely use [their] brain” instead of “just scrolling mindlessly for an hour.” P09 envisioned a system like an elevator: “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor… Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” This sentiment was echoed by several participants who envisioned brain-controlled interfaces that would “automatically show [users] things [they] would really like” (P07) based on their immediate needs or interests.
-4.5.2. Reduced Celebrity and Commercial Influence
 Several participants advocated minimizing celebrity influence.
@@
 P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
-Other participants expressed a need for control over how non-personal content is organized and placed on the platform.
-P04 suggested being able to “figure out how I want to arrange strangers’ spaces in my… platform.” They also wanted the ability to “prioritize things and feel more in control of how much information is fed to them,” imagining a system where apps and notifications are ranked by personal relevance.
-P18 saw ads as inevitable but wished they could appear in a separate “billboard” that users could skip.
-4.5.3. Spatial Representation of Relationships
-Youth envisioned SMH as a platform that reflects the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties.
-Participants appreciated this design as it prioritized interactions with closer friends, creating a sense of community and authenticity.
-P02 highlighted, “The fact that only my close friends have access to it… is the defining factor,” enabling users to focus on “better conversations with other people [that] already know who I am,” without the pressure to “advertise who I am” as they might feel on MSM.
-Some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction.
-P03 imagined, “The people around you… would be people from the same house as you,” cultivating a shared sense of “togetherness” and “community building” through common spaces and interactions.
-Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
-This design approach—aligning physical landscapes with emotional closeness—was seen as a way to streamline interactions and focus on meaningful relationships.
-P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.” Unlike MSM feeds, which participants felt prioritized engagement over relationship strength, SMH’s spatial arrangement allowed users to engage naturally with those closest to them.
-P09 further elaborated on how this physical representation could enhance navigation: “have your friends’ rooms, or like your kind of close friends list, and maybe you’re kind of living in a pod with them, so you can move between the rooms a lot quicker.”
-4.5.4. Intentional and Meaningful Use of Time
-Participants expressed a strong preference for digital experiences that feel meaningful and intentional, contrasting these with the negative feelings often associated with aimless social media use.
-P10 described how social media can feel “meaningless” when they “didn’t learn something,” while P18 reflected on the regret they feel after “mindlessly scrolling.” Similarly, P14 shared that scrolling through Instagram Reels while “literally just laying in [their] bed” feels like “wasting [their] time.”
-In contrast, participants found interactions involving direct communication and engagement more fulfilling.
-P16 shared that “anytime not scrolling” and instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this sentiment, noting that social media use feels more meaningful when they are “actually talking to someone.” They explained that such interactions create a sense of “connectivity,” particularly when keeping in touch with people “not in close proximity.”
-The distinction between passive and active engagement was another recurring theme.
-P10 explained that while doomscrolling does not make them feel “fulfilled,” playing games like Minecraft is different because it involves “progress” they can “save and come back to.” Similarly, P14 shared that the lack of connectivity on social media drives their “doomscrolling,” but they believe platforms that encourage interaction could reduce screen time by making time spent online more meaningful.
-P06 also note that they noted that they “don’t feel bad… most of the time [they] use Roblox,” as it involves engaging activities such as “playing games with friends” and “customizing spaces.”
-Given such differences in how time spent is perceived, participants expected time spent on SMH to feel more meaningful than that on MSM.
-P04 noted that while MSM allows for quicker interpersonal interactions, it often leads to “doomscrolling.” In contrast, SMH would, for example, encourage “thoughtfully engaging with someone’s memories,” leaving them feeling “much more fulfilled.” P05, similarly, explained that they would not feel bad about spending more time on a platform like SMH because it involves “more effort and more time… a more intentional way of interacting.” P07 affirmed this, stating they would be willing to spend more time on a platform like SMH given interactions where they would find interactions “more enjoyable.”
-4.6. Reimagining Social Media’s Potential
-Given the nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what social media could be.
-Therefore, while they acknowledged potential challenges—such as the platform being overly immersive, addictive, or demanding—they largely viewed these downsides as worthwhile trade-offs for its many benefits.
-Overall, participants saw SMH as an inspiring alternative to existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional.
-They also appreciated the FI workshops for empowering them to reimagine social media and to be hopeful about its possibilities in ways they had previously thought impossible.
-4.6.1. Fostering Hope for the Positive Potentials of Social Media
+4.3.4. Developing Socially Grounded and Themed Conversational Spaces
+Participants described how themed, socially grounded conversational spaces designed for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
+“Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread.
+Like how I talked last time about Reddit, you know, there’s a thread, and then you make a comment on it, but there are a hundred other comments, so no one’s going to respond to you.
+I want to focus more on having actual conversations with people, making real connections, and more individualized chatting.
+I find it’s way easier to get to know people a lot better.” (P22)
+Participants described how the design and feel of conversational spaces could make social interactions on SMH feel safer, more personal, and more inviting than on existing platforms.
+Many valued the idea of walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to a broad, undefined audience.
+P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations.
+They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
+Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected.
+P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate.
+P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n.
+d.]) where users could create custom rooms on demand to match the “purpose” of their conversation.
+Participants also valued interest-based spaces, similar to Discord (discord page, [n.
+d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
+4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
+To ensure SMH feels like a safe, personal environment, participants stressed the need for more comprehensive boundary controls beyond privacy tools found on existing platforms that they perceived as limited and inconsistent.
+P22, for example, explained that even with contact syncing turned off on Instagram, the platform still suggested their Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they were trying to avoid, based on mutual follow relationships.
+They stressed the need for stronger, more thorough blocking controls, including the ability to block “alternate accounts” and fully remove any trace of a blocked person’s presence, explaining, “I don’t wanna be able to tell that the person I blocked is even around me.” P14 similarly pointed out that even private accounts on Instagram still expose personal content, like Story Highlights, to “a random person”, revealing perceived gaps in platform-level privacy protections (Bojkov, 2019).
+Participants expressed a desire for stronger control over their own visibility, imagining features that would let them opt out of being seen or discovered entirely.
+P03 suggested a feature like an “invisibility cloak” for moments when users “don’t want anyone to see you.” P07 proposed a “super private” profile mode where users “can’t even be searched… like you’re not even on the app” unless they choose to make themselves visible.
+Preventing unsolicited contact from strangers was another recurring priority.
+P22 proposed a model inspired by Discord, where only people you’ve already interacted with can contact you, and others “can’t even be discovered unless you have something in common.” This idea resonated with P13, who argued that new connections should require mutual consent, suggesting that people “should only be able to add you if you meet in person and exchange a QR code.” They shared a personal experience of being able to find a stranger online based on just their first name, reflecting that “it’s too easy to find people on social media.”
+4.4. Nuanced Self-Presentation
+Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14).
+This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23).
+P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting the full range of who they are, without the pressures of curation or performance.
+4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
+Participants consistently described customizable avatars and personal spaces as meaningful tools for self-expression, offering more creativity, personality, and authenticity than the static profiles or limited settings found on MSM.
+Unlike MSM, which typically offers only “two modes” (P10), i.e., “light and dark” (P10), participants imagined a much richer and more immersive experience.
+For example, P21 envisioned a fully customizable world where users could build their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” These spaces could be shared with friends as a way of “showing a little part of yourself.” Similarly, P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture,” to reflect their personal interests.
+Customizable “3D avatar[s]” (P11) and profiles were also seen as key ways of showing individuality.
+P22 reflected on the early days of Myspace (mys, [n.
+d.]), describing how they enjoyed coding their profile layouts and changing “the font and really make it reflect me.” Even small details, like choosing a username, were described as ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted that such forms of expression are often “mainly for you,” not just for public display.
+P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right?
+So I think that’s part of you.”
+Beyond visual customization, participants imagined SMH supporting more ambient, indirect forms of expression, offering subtle cues about what someone is doing, feeling, or thinking without needing to make a formal post.
+For instance, P01 described wanting their avatar to reflect real-time states, like the outfit they are wearing or the song they are listening to, so that friends could intuitively notice things like, “‘Oh.
+that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ Similarly, P10 imagined that friends could pick up on changes in their interests simply by observing how they update their virtual space over time, providing “an insight into… what I’m thinking.”
+Several participants highlighted the contrast between the flat, compressed feel of 2D platforms and the immersive qualities of 3D environments.
+P10 noted that in 3D, others can see “as much stuff as you can see in 3D,” making it feel richer and more complete.
+P16 added that VR creates a sense of presence that feels more like meeting “real people… compared to the profile pictures,” because you can pick up on their “mannerisms” and notice details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
+Participants also described how these kinds of choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests.
+As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation.
+P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
+Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
+“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.
+It’s where you spend most of your time, and it shows your personality the most.
+It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes or 30 seconds the way you can on Instagram.
+If you wanted to add someone there, all you’d have to do is remove a couple of things.
+It’s easier than physically changing a space.” (P13)
+Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes.
+P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention.
+They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.
+4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
+Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides of themselves across various spaces and relationships, and reducing the pressure to maintain a single, polished identity.
+Rather than being limited to a single, polished identity, as is often the case on MSM, participants wanted ways to represent the full range of their interests, identities, and relationships.
+P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things.
+You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are.
+P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
+4.4.3. Sharing Everyday Activities and Lived Experiences to Deepend Understanding and Connection
+Participants valued the ability to share casual, everyday moments and rich, embodied memories, filling the gaps between the highlights usually shared on MSM into a more natural, meaningful extension of daily life and personal connection.
+P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment.
+For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
+Participants also imagined SMH supporting the sharing of more personal, embodied memories as a way to deepen understanding and reduce misinterpretation.
+P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n.
+d.]) as a way to create stronger emotional connections.
+Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully.
+P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
+Some participants imagined that these shared memories could include sensory or emotional details, not just visuals.
+P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication.
+Others compared it to live streaming personal experiences.
+P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record.
+P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
+4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
+Participants emphasized that flexible, contextual privacy is essential for supporting authentic self-presentation, allowing people to manage who sees what, when, and how, without feeling overexposed or misrepresented.
+They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life.
+P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
+Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces.
+P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts.
+Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation.
+P03 proposed a model where everyone maintains a “base” identity but customizes additional layers.
+P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
+Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors.
+P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time.
+As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on.
+Or turn it off.”
+Participants also highlighted the importance of individual control over these transitions and boundaries.
+P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
+4.5. Intentional Engagement
+Participants described their growing frustration with how current social media platforms prioritize passive consumption over meaningful connection.
+In response, they envisioned SMH as a space that promotes intentional, socially rewarding interactions.
+This section explores four key themes that reflect their desire to refocus social media on authentic connection, purposeful engagement, and meaningful use of time.
+4.5.1. Encouraging Deliberate Navigation and Content Discovery
+Participants imagined SMH as a space that encourages deliberate, effortful discovery, replacing the passive, algorithm-driven feeds of MSM with experiences that require users to actively choose what and who they engage with.
+P22 described their frustration with Instagram, explaining that “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage more intentional, purposeful engagement, both when browsing content and when navigating the virtual world.
+P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful.
+Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
+Participants also imagined features that would prompt users to actively choose what they wanted to engage with before entering a space.
+P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
+Physical movement through SMH’s world was also seen as key to this intentionality.
+Participants valued having to walk, explore, and navigate to find people, spaces, or activities, turning even small actions into social moments of presence and care.
+P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
+This desire for intentional interaction extended to how users move through the platform itself.
+Participants imagined physical navigation, such as walking or traveling through virtual space, as a way to make engagement feel more meaningful.
+P02 explained that they would prefer slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions.
+I want everything to be very thought through.” While features like “spells” (P06) or the “Marauder’s Map” (P21) were imagined as convenient options, participants valued the ability to walk and explore, turning simple actions like visiting a friend or browsing a shop into shared social moments.
+P03 explained that this kind of embodied interaction feels “more connected than regular social media” (P03) because it requires active presence rather than passive scrolling.
+P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them.
+They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it.
+It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
+Participants believed that this kind of effortful navigation would naturally discourage casual lurking or influencer-driven engagement, shifting the platform’s focus toward genuine connection.
+As P05 put it, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
+4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
+Beyond how users navigate content, participants also emphasized the importance of shifting the platform’s core purpose from endless consumption toward meaningful social connection.
+Participants expressed frustration with how current platforms promote mindless, passive consumption, and described their hopes for SMH to instead foster more meaningful, intentional connections through active social engagement.
+They described frustration with how easily time slips away on platforms like Instagram and TikTok without leaving them feeling unfulfilled.
+P10 noted that social media often feels “meaningless” when they “didn’t learn something,” and P18 reflected on the regret of “mindlessly scrolling.” P14 similarly described spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
+A major contributor to this feeling was the infinite scrolling design of MSM.
+P11 explained that “there’s always more content on social media,” which fuels the fear of missing out and the urge to keep scrolling.
+P14 critiqued TikTok for exposing users to “clickbaity” and often low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
+In contrast, participants described active, social engagement, even if casual, as far more rewarding.
+P16 shared that “anytime not scrolling” but instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this, noting that meaningful use comes when they are “actually talking to someone,” especially to stay connected with people “not in close proximity.” They reflected that while users might spend more time on a platform like SMH, that time would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to passively watching “50-second TikTok videos of somebody making slime for like an hour.”
+Participants also contrasted passive consumption with active participation in interactive environments.
+P10 explained that while doomscrolling feels unproductive, activities like playing Minecraft feel different because they offer “progress” that can be “saved and come back to.” P06 echoed this sentiment, noting that they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
+Reflecting on SMH’s potential to create this kind of active, engaging experience, P09 acknowledged that “time spent might not be mitigated,” but emphasized that “the way you’re spending it” would make the difference.
+They explained that SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting that meaningful engagement, rather than passive consumption, was what ultimately mattered for them.
+4.6. Playfulness and Whimsy
+Participants described SMH as a platform that feels youthful, playful, and deeply immersive, offering not just entertainment but opportunities for creative self-expression, social bonding, and emotional well-being.
+By embracing whimsy and delight, SMH was seen as making social media feel more personal, less pressured, and more meaningful.
+4.6.1. Embracing Creative Joy and Playful Self-Expression
+Participants imagined SMH as a platform that transforms social media into a space for creative joy and playful self-expression, offering users the freedom to experiment with identity, explore personal interests, and connect with others through creativity.
+Several participants described how SMH offered far more than static profiles or simple feeds.
+P06 reflected that “everyone likes sharing parts of themselves,” appreciating how the platform allows users to shape not just what they say, but how they appear and interact with others.
+P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
+For some, this act of personal customization felt like a meaningful process of identity exploration.
+P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression.
+P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know.
+That’s where the fun aspect comes in, really, for me.”
+This spirit of creativity extends beyond the individual to the wider environment and community.
+Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa!
+They built that, that’s so cool” (P06).
+SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here.
+No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06).
+The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms.
+Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
+4.6.2. Designing Engaging and Playful Privacy Mechanisms
+Participants imagined privacy not as a tedious and confusing task but as part of the playful experience, envisioning interactive, game-like mechanisms that turn boundary-setting into something fun, discoverable, and socially meaningful.
+P11 envisioned systems where users could control access to spaces through gestures like a “secret knock” or hidden objects that revealed “secret passages” accessible only to trusted friends.
+Others suggested knowledge-based gates, such as “secret handshakes” (P21) shared among friends.
+Drawing inspiration from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces only when correctly summoned.
+These features were described not just as functional but as adding delight and novelty to social interaction.
+P05 compared them to “little Easter eggs,” remarking that “there’s that little kid part of everyone” that makes such features enjoyable.
+They contrasted these with conventional privacy settings on mainstream social media, which they described as “a wall of settings” that often feel confusing or uninviting.
+4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
+Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities.
+P01 questioned the need for phones at all, asking, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting “holographic screens, like Jarvis in Iron Man,” or activating “panels” and “big screens” through spells.
+Others suggested alternative magical devices.
+P15 preferred VR glasses, describing them as “more natural,” “easier,” and “more fun.” Participants also imagined objects like a “makeup mirror” that could serve as a portal, a “watch” that immerses users into SMH, or a “mirror glass” for video calls or summoning friends with spells like accio.
+These magical affordances were not only imagined as novel interfaces but also as ways to make extraordinary experiences more accessible.
+P09 highlighted that SMH could give people “experiencing stress” or lacking access to “really, really beautiful landscape and environments” the chance to walk around and “admir[e]” calming, scenic spaces.
+P16 imagined that users could take part in activities they might never otherwise experience, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
+4.7. Expanding Horizons
+Given the speculative nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what interactions on social media platforms could embody.
+Therefore, while they acknowledged potential challenges, such as the platform being over-immersive or distracting, they overwhelmingly viewed these downsides as worthwhile trade-offs for its many benefits.
+Overall, participants described SMH as an inspiring alternative or expansion of existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional.
+They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
+4.7.1. Breaking Free from Platform Monotony with Novel Interactions
 Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become.
@@
 Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram.
-P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “fun” and “novel,” like a “breath of fresh air.”
-Beyond features, participants expressed how SMH inspired hope for the future of social media. P10 pointed out that overcoming the design fixation around social media is crucial:
+P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “novel,” like a “breath of fresh air.”
+For some participants, SMH’s novelty lay in giving themselves permission to prioritize enjoyment over seriousness.
+P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun.
+For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
+“Yeah, [I would prefer the 3D version] actually 100%.
+I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult.
+But who cares?
+I feel like something that’s more enjoyable would be [the SMH].” (P14)
+P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms.
+P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness.
+P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
+4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
+Through the FI process, participants realized they were not limited by the structures of existing platforms.
+Instead, they felt empowered to imagine social technologies that center joy, creativity, and authentic human connection, qualities they had once believed were beyond the reach of digital spaces.
+P10 pointed out that overcoming the design fixation around social media is crucial:
 “The way that we can make social media better is if we expand the realm of what the definition of social media is.
@@
 (P14)
-4.6.2. Celebrating Youthful Creativity and Play
-SMH allows a playful and engaging social media experience, allowing users to immerse themselves in actively creative activities and self-expression.
-P06 appreciated this ability to personalize their experience on SMH, stating, “I think… everyone likes sharing parts of themselves.” P21 likened SMH to a “game” due to its creative freedom.
-They elaborated, “The creative freedom you have over the platform makes it feel more fictional than it is real.”
-Participants found the overall experience of SMH to be youthful, fun, and deeply immersive, offering a space that feels both magical and meaningful.
-For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
-“Yeah, [I would prefer the 3D version] actually 100%.
-I feel like,… the only reason I said 2D is because I felt like the need to be like a mature adult.
-But who cares?
-I feel like something that’s more enjoyable would be [the SMH].” (P14)
-P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms.
-P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness.
-P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
-4.6.3. Enabling Otherwise Inaccessible Experiences
-SMH creates opportunities for users to engage in experiences and interactions that may otherwise be out of reach.
-P09 highlighted how individuals “experiencing stress” or those who may not have access to “really, really beautiful landscape and environments” could walk around and “admir[e]” the “calming” landscape of SMH.
-P16 envisioned SMH as a space where users could engage in activities like climbing a mountain in Nepal, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
-P02 saw SMH as a platform that would “seem more like real life, but better.” They imagined being able to engage with friends in idyllic settings, such as “on the beach all the time.” Furthermore, they believed SMH could strengthen friendships through magical features.
-In their words, “Since people can’t read others’ emotions in real life, SMH would… be a way for you to connect with your friends on a deeper level.” P18 also appreciated how SMH’s 3D interface facilitated them to interact with “family across the world” in an immersive way.
-SMH would offer a sense of presence that goes beyond conventional video calls, enabling users to engage with loved ones “not just through a screen.”
 5. Discussion
-Figure 3.
-(a) Current youth perceptions of social media platforms and where they experience meaningful connections, highlighting the disconnect between platforms commonly recognized as “social media” versus where youth actually experience meaningful social connection.
-The SMH—youth-envisioned ideal social media—is represented by the overlap, combining both aspects.
-(b) A more nuanced viewpoint clarifies “social media” as an umbrella term for a range of platforms with different primary purposes—friendship-building, entertainment, status, information-seeking, tie maintenance, and more—rather than a few archetypal platforms.
+Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
 […]
-This diagram is divided into two parts, offering insights into how youth perceive social media and how they categorize its purposes.
-On the left, there are two overlapping circles.
-The circle on the left represents platforms that youth identify as “social media,” such as Instagram, Facebook, and TikTok.
-The circle on the right illustrates platforms where youth feel meaningfully connected with friends, including FaceTime, Discord, Minecraft, and Roblox.
-In the overlapping area falls Social Media at Hogwarts (SMH) youth-envisioned ideal social media that combines the qualities of both categories.
-On the right, a Venn diagram shows potential categories of social media platforms by their primary purposes, such as “Entertainment-Centered,” “Status-Centered,” and “Information-Seeking-Centered.” “Social Media at Hogwarts” is grouped under “Strong Tie-Centered.”
-5.1. “Social Media is 2D”: Youth Perspectives on What Constitutes Social Media
-Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction.
-This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways in which they happen to be designed today (Kim et al., 2024b).
-Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections.
-As illustrated in Figure 3(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections with friends (such as Discord or Minecraft).
-When envisioning their ideal social platform, participants described something that would bridge this divide—a space combining the deliberate pursuit of social connection they associate with traditional social media with the organic, engaged interactions they naturally experience in gaming and chat platforms.
-This misalignment between expectations and experiences around social media is problematic because negative perceptions of social media use can diminish psychological well-being (Lee and Hancock, 2024; Kim et al., 2024a).
-Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce.
-As shown in Figure 3(b), social media serves diverse uses and gratifications—from information seeking and entertainment to status enhancement and network maintenance.
-Rather than maintaining an entirely negative view of social media confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations.
-For instance, critiquing Instagram for not facilitating authentic, deep friendships misses the point; while such connections might occasionally form there, it is not the platform’s central form of engagement.
-However, it is important to note that platforms aren’t static—their designs evolve, and user engagement patterns shift over time.
-Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts.
-This fluidity in platform use underscores the importance of teaching (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025) youth to be more discerning about the connections they seek and what’s possible on each platform rather than enforcing anxiety-inducing narratives about social media’s harmful effects.
+A Venn diagram comparing two categories of online platforms.
+The left circle represents platforms focused on social networking and broadcasting, with icons for Instagram, Facebook, and TikTok.
+The right circle represents platforms where youth engage in meaningful connections, with icons for Zoom, Minecraft, Discord, and Roblox.
+In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
+Figure 2 captures how participants described their vision for ideal social media as something that pulls benefits from both types of platforms they already use.
+On one side are social networking and broadcasting platforms like Instagram, Facebook, and TikTok, which make it easy to stay socially visible but often feel shallow or performative.
+On the other side are platforms like FaceTime, Minecraft, Discord, and Roblox, where youth experience more organic and engaging connections, but where social interaction is often secondary to other goals like gaming or productivity.
+Our participants imagined “Social Media at Hogwarts (SMH)” as a platform that blends these strengths, creating a space where meaningful social connection effectively surfaces not just a peripheral effect, but a key focus of the platform.
 5.1. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
-Our findings confirm prior studies that while the youth are largely dissatisfied with social media and how they spend their time on archetypal social media platforms, they do not necessarily want to reduce their social media usage—rather, they seek platforms that facilitate deeper, more meaningful connections.
-Our participants identified four key elements that could transform social media from a space of passive consumption to one of genuine connection:
+Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online.
+Instead, they seek spaces that support deeper, more intentional social interactions.
+Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative.
+The changes in social media platform design that they sought were as follows:
   1. (1)
-    Presence and Immersion.
-Youth envision platforms that foster genuine emotional connection through real-time presence.
-While current platforms feel “flat” and emotionally disconnected, participants valued synchronous interactions that create a sense of co-existence.
-Even simple activities like sharing quick updates about eating ice cream or walking outside, or simply existing in a shared virtual space were seen as meaningful when done with immersive co-presence.
+    Immersive Experiences Through Spatial Elements.
+Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences.
+They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed.
+They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied.
+They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures.
+Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
   2. (2)
-    Supporting Individual Preferences.
-Youth desire more natural and diverse ways to express themselves.
-Participants envisioned platforms where they could show their personality through both active sharing and subtle signals, similar to how personality naturally emerges in physical spaces.
-From memory-sharing to customizing their virtual spaces and avatars, they want to express themselves spontaneously and casually rather than feeling constrained to carefully curated posts.
+    Organic and Intentional Social Interactions.
+Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions.
+They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally.
+They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out.
+Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people.
+Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
   3. (3)
-    Intuitive Social Navigation.
-Meaningful social media experiences require design that mirrors how relationships naturally unfold.
-Youth value platforms that enable organic relationship growth through proximity and repeated low-intensity interactions, similar to how sitting in the same classroom gradually builds connections.
-They also seek designs that replicate the subtle, embodied norms of physical spaces—where spatial arrangements naturally signal degrees of privacy and social contexts, much like how a bedroom implies different privacy expectations than a living room.
+    Expressive and Lower-Stakes Sharing.
+Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts.
+They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars.
+They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection.
+Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
   4. (4)
-    Playful Engagement.
-Youth emphasize that meaningful connection often grows through joy and shared experiences, yet they find current platforms overly serious and high-stakes.
-They envision platforms that encourage whimsical exploration and casual interaction—where they can discover hidden objects, explore virtual landscapes, or simply hang out in a “third place.” This playfulness creates low-stakes opportunities for friendship development beyond the performative nature of mainstream social media.
+    Granular, Intuitive, and Playful Privacy Mechanisms.
+Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid.
+They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries.
+They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable.
+Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments.
+They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
+  5. (5)
+    Refocusing Interaction Priorities for Meaningful Engagement.
+Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about.
+They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams.
+They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections.
+Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape.
+Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
 These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams.
@@
 While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
-Our participants envision environments where relationships deepen naturally through shared experiences and deliberate engagement rather than through metrics-driven interactions.
-This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement—challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
+Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions.
+This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
 5.2. Spatial Integrity for Meaningful Social Connections in Online Environments
-In the design of virtual environments that foster meaningful social connections, spatial cognition—the way we perceive and interpret space—remains critically relevant.
-Even in virtual spaces, human interactions are shaped by spatial affordances, mirroring many aspects of physical-world interactions.
-While virtual spaces need not replicate physical environments exactly—indeed, they can transcend physical limitations in some ways while falling short in others—certain spatial principles remain fundamental to how humans build and maintain relationships.
-This relevance is reflected in our pervasive use of spatial metaphors when describing virtual interactions, such as “chat rooms,” “cyberspace,” or “folders.” While some platforms like Discord and social VR applications successfully implement spatial design elements, mainstream social media platforms often lack these affordances.
-Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces.
+Concept Description Supported Interactions
+Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence.
+Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation.
+Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
+Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction.
+Platforms like Discord create stable communal spaces with clear channels and streams.
+Users value this for fostering deeper relationships.
+Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
+Spatial Composition The segmentation of a social network into defined areas and pathways.
+Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces.
+Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
+Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting.
+3D social media limits expression to explicit actions like posting and commenting.
+Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
+Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
+Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
   1. (1)
@@
 Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019).
-Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offers only transient content streams.
-Our participants emphasized the importance of dedicated areas—like communal lounges or study spaces—where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
+Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams.
+Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
   3. (3)
@@
 Clear boundaries and pathways help people recognize territories and manage social norms.
-Current platforms lack such clarity—a user’s feed” is meant to be personal space yet lacks clear boundaries, while profiles can be accessed without users’ awareness.
+Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness.
 This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
@@
 Incorporating spatial depth could enable richer self-expression and more natural interactions.
-These four dimensions of spatial integrity—presence, composition, configuration, and depth—are critical for designing virtual environments that support meaningful social connections.
-While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram that we explored in Section 5.2.
-As spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms.
-(It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).)
-The concept of spatial integrity offers a framework for understanding why many social media platforms feel disconnected and unsatisfying.
-By thoughtfully incorporating these principles, we can design online environments that better align with human social-spatial cognition, addressing the limitations our participants identified and better supporting meaningful social interactions and the organic development of relationships.
+According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire.
+The relationship between these dimensions and the interactions they support is laid out in Table 2.
+These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections.
+By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish.
+While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram.
+Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms.
+(It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
 5.3. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
+There are societal concerns about social media’s impact on youth well-being (Oth, 2023).
+Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023).
+This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
+However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality.
+Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023).
+At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024).
+This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connection that youth actually seek.
+Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
 Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be.
@@
 Our study has several key limitations to consider.
-First, all participants had prior experience with 3D gaming platforms, potentially biasing them toward favorable attitudes toward incorporating spatial elements or online environments that center play.
+First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently generalize to youth who are not familiar with such environments.
 Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations.
@@
 This suggests opportunities to refine the narrative approach.
-Future research could empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces.
+Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces.
 This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
 Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant.
-They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him—including personal details they would not share with friends.
+They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends.
 Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share.
@@
 6. Conclusion
-Through co-design workshops employing the Fictional Inquiry (FI) method and leveraging a Hogwarts-inspired narrative, we explored the youth’s vision for an ideal social platform.
-The interviews revealed a disconnect between mainstream social media and the kinds of meaningful online connections young people seek.
-While platforms like Instagram have become synonymous with “social media,” youth find their most authentic social interactions happening elsewhere—in gaming environments, chat platforms, and other spaces that they don’t typically seek as destinations for social connection.
-The participants collectively emphasized four key elements missing from current social platforms: emotional connection through real-time presence and immersion, natural forms of self-expression beyond carefully curated posts, intuitive social navigation that mirrors physical-world relationship development, and playful low-stakes opportunities for friendship growth.
-These findings challenge the widespread assumption that engaging youth requires addictive features or endless content streams.
-Rather, young people express a clear desire for time that feels meaningfully spent, imagining that social media platforms enabling meaningful connection through shared experiences and deliberate engagement would provide exactly that.
-Importantly, this research demonstrates how moving beyond standard design patterns through methods like Fictional Inquiry can help youth articulate possibilities they might have previously dismissed as impossible within conventional social media paradigms.
-Rather than accepting current platforms’ limitations as inevitable, participants were able to envision radically different approaches to online social connection and feel empowered and hopeful about the positive potential of social media that they initially felt pessimistic about.
-Further, we introduce spatial integrity as a framework for facilitating meaningful social connections in online environments, helping to understand why many social platforms feel disconnected and unsatisfying while pointing toward design principles that could better support meaningful online relationships.
-From 2D platforms to spatial technologies, our study highlights an opportunity and hopeful future for creating social experiences that work with, rather than against, our natural needs and intuitions for social connection.
+This study explored how 3D gamer youth reimagine social media when given the freedom to think beyond the limits of today’s platforms.
+Through the fictional world of Hogwarts, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption that defines mainstream social media.
+Their visions called for platforms that feel more like 3D games than archetypal ones like Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to the complexity of real relationships and incorporate elements of playfulness and whimsy.
+Building on these insights, we introduced the concept of spatial integrity, a set of foundational design principles that make such visions possible.
+While spatial integrity alone does not guarantee meaningful connection, it provides the structural conditions that support the kinds of relational, embodied, and playful interactions participants described.
+The study underscores that youth are not necessarily seeking less time on social media, but rather more fulfilling and emotionally present interactions within these digital spaces.
+Their visions point towards a future where social media feels more like a dynamic, shared social world like those found in 3D gaming environments.
+While our findings are rooted in the experiences of youth already familiar with 3D gaming, they offer a hopeful design direction for spaces that support meaningful social connections that youth desire.
 Appendix A FI Workshop Protocol

### v3

@@
 1. Introduction
-Meaningful social connections play a critical role in young people’s development and well-being (Hawkley and Cacioppo, 2010), and many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n.
+Meaningful social connections are critical for young people’s development and well-being (Hawkley and Cacioppo, 2010), and social media has great potential to support these.
+At the individual level, richer and healthier interactions support youth well-being and identity development (Kim et al., 2024c).
+At the societal level, platforms act as critical infrastructures for sustaining friendships and community ties, making their design choices consequential for social capital (Phua et al., 2017a; Ellison et al., 2007).
+In fact, many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n.
 d.]).
-However, recent trends indicate that social media platforms are increasingly becoming less focused on facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025).
-Initially designed around sharing personal updates with friends, these platforms now often emphasize entertainment, news, and engaging larger, less intimate audiences.
-Meta’s own data illustrates this shift: between 2023 and 2025, the share of time users spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025).
-Mark Zuckerberg described this change as aligned with the company’s focus on supporting users “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
-This shift away from personal interactions raises concerns about the future of meaningful online relationships, particularly for younger users who have grown up with social media as a central part of their social lives.
-Most youth today engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023).
-Despite this high engagement, many young users report dissatisfaction with the quality of their social interactions, not because they wish to reduce their usage, but because they expect these platforms to better support their social connections (Kim et al., 2024b).
-Current platform designs frequently prioritize passive consumption of content over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
-Given the central importance of social connections, particularly for youth, and recognizing social media’s vital role in facilitating these interactions, it is increasingly important to explore new design possibilities.
-Although young people rely heavily on social media, they frequently express dissatisfaction with their current experiences, indicating a significant gap between their social needs and the interactions enabled by existing platforms.
-Current mainstream social media platforms employ highly standardized design patterns, which limit our collective imagination and contribute to feelings of pessimism among both younger and older users.
-Consequently, there is a compelling need for innovative designs that complement existing platforms by placing a stronger emphasis on meaningful interpersonal connections.
-This prompts the critical question of understanding what types of social interactions current designs might be overlooking or inadequately supporting.
-To address this question, we conducted a series of co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007).
-We involved 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n.
+However, social media platforms increasingly shift away from facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025).
+While designers initially created these platforms around sharing personal updates with friends, they now emphasize entertainment, news, and engaging larger, less intimate audiences.
+Meta’s data illustrates this shift: between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025).
+Mark Zuckerberg described this change as aligned with the company’s focus on “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
+This shift raises concerns about meaningful online relationships, particularly for young people who have grown up with social media as central to their social lives.
+Most young people engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023).
+Despite high engagement, many report dissatisfaction with their social interactions, not because they want reduced usage, but because they expect platforms to better support their connections (Kim et al., 2024c).
+Current designs prioritize passive content consumption over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
+Given social connections’ importance for young people and social media’s vital role in facilitating them, we must explore new design possibilities.
+Young people frequently express dissatisfaction with current experiences, revealing a significant gap between their social needs and existing platform interactions.
+Mainstream platforms employ standardized design patterns that constrain creative possibilities and foster pessimism among users about social media’s potential.
+This creates a compelling need for innovative designs that complement existing platforms by emphasizing meaningful interpersonal connections.
+Therefore, we ask the following research question: What types of social interactions might current social media designs be overlooking or inadequately supporting?
+To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) with 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n.
 d.]) and Roblox (Roblox Corporation, [n.
 d.]).
-Participants were asked to collaboratively imagine a social media platform specifically designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n.
+Participants collaboratively imagined a social media platform designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n.
 d.]).
-This imaginative context allowed participants to step away from the constraints of existing platforms and envision more varied and meaningful social interactions.
-Additionally, because 3D gaming environments are widely recognized as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated that participants could leverage insights from their gaming experiences to propose novel social media interactions.
-Specifically, our research aims to answer the following question: What types of interactions do youth familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
-Our findings reveal that youth with 3D gaming experiences envision ideal social media platforms not as immersive, navigable environments designed to feel more like neighborhoods, community centers, and shared social spaces.
+This imaginative context allowed participants to step away from existing platform constraints and envision more varied and meaningful social interactions.
+Since researchers recognize 3D gaming environments as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated participants could leverage gaming insights to propose novel social media interactions.
+Specifically, our research addresses: What types of interactions do young people familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
+Our findings reveal that young people with 3D gaming experience envision ideal social media platforms as immersive, navigable environments that designers should create to feel like neighborhoods, community centers, and shared social spaces.
 Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming.
-Together, these visions highlight the youth’s desire for social media that feels more relational, meaningful, and personally rewarding.
-We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media.
-Additionally, we propose the concept of spatial integrity for social interactions in online spaces, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online.
-These dimensions include spatial presence (the ability to understand the space and identify who is present), spatial composition (the meaning and significance of the space), spatial configuration (the organizational structure of the space), and spatial depth (the extent of information that can be perceived within the space).
-Furthermore, we demonstrate that the FI method, in the context of social media design, not only prompted creative idea generation and enabled participants to transcend current social media design patterns but also empowered them to envision the positive potential of social media, leaving them feeling more hopeful about its future possibilities.
+These visions highlight young people’s desire for social media that feels more relational, meaningful, and personally rewarding.
+We contribute empirical findings on young people’s perceptions of current social media platforms’ shortcomings and their desired experiences.
+We propose the concept of spatial integrity for social interactions in online spaces, encompassing four dimensions of spatial affordances that facilitate meaningful social connections: spatial presence (understanding the space and identifying who is present), spatial composition (the space’s meaning and significance), spatial configuration (the space’s organizational structure), and spatial depth (the extent of perceivable information within the space).
+We demonstrate that the FI method in social media design contexts prompts creative idea generation, enables participants to transcend current design patterns, and empowers them to envision social media’s positive potential, leaving them more hopeful about future possibilities.
 2. Related Work
@@
 Litt et al.
-(Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” While these interactions are described as “deeply subjective,” the primary distinction from non-meaningful ones lies in the presence of impact.
-This impact extends “beyond the immediate interaction itself,” typically involving positive emotions, the exchange of information, or concrete benefits.
-The study identified three key types of impact: “emotional impact” (involving feelings, mood changes, or relationship strengthening), “informational impact” (relating to the exchange of knowledge, advice, or understanding), and “tangible impact” (providing concrete benefits such as goods or services).
-In contrast, non-meaningful interactions were often explicitly described as lacking any such impact.
-Factors like the presence of strong ties (e.g., friends and family) and community ties, as well as participation in shared activities, were found to be associated with meaningfulness.
-However, these appeared to act more as facilitators that enable impact rather than as the primary source of meaning.
-Other attributes quantitatively linked to greater meaningfulness included whether interactions were planned in advance, memorialized (for example, through photos or videos), and synchronous.
-Interestingly, the study found that the communication medium itself, whether online or offline, was less important in predicting meaningfulness after accounting for these other characteristics.
-While in this paper we treat meaningful social interaction as any interaction that participants say they would actively seek out in an ideal scenario.
-reflecting their subjective or perceived meaningfulness, we expected these themes to echo in how our participants talked about their experiences.
+(Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself,” involving positive emotions, information exchange, or concrete benefits.
+The study identified three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services).
+Non-meaningful interactions lacked any such impact.
+Strong ties (friends and family), community ties, and shared activities were associated with meaningfulness, but acted as facilitators enabling impact rather than primary sources of meaning.
+Other attributes linked to meaningfulness included advance planning, memorialization (photos or videos), and synchronicity.
+The communication medium itself—online or offline—was less important than these other characteristics.
+We define meaningful social interaction as any interaction that participants would actively seek in an ideal scenario, reflecting their subjective meaningfulness.
+We expected these themes to emerge in participants’ descriptions of their experiences.
 2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
-Social media platforms are often seen as a primary means of connecting with others, especially for young people (Boyd, 2014a).
-However, despite the widespread use of these platforms, many users report feeling lonely or disconnected (Twenge et al., 2019; Primack et al., 2017).
-This disconnect highlights a significant gap between the intended purpose of social media and the actual experiences of users.
-Social connection is just one of the many reasons people turn to social media, but it remains one of the most important.
-From staying in touch with close friends to meeting new people or maintaining community ties, users often come to these platforms seeking connection.
-Mainstream platforms like Instagram (All, [n.
+Though social media platforms are primary connection tools for young people (Boyd, 2014a), many users report feeling lonely or disconnected despite widespread use (Twenge et al., 2019; Primack et al., 2017), highlighting a gap between intended purpose and actual experience.
+While social connection remains a primary motivation for social media use, mainstream platforms like Instagram (All, [n.
 d.]), YouTube (you, [n.
 d.]), and Discord (discord page, [n.
-d.]) support some of these needs, but they are designed to address a wide range of user motivations, including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022).
-From a uses and gratifications perspective (Katz et al., 1973), even within a single platform, users pursue diverse goals, ranging from building social capital (Phua et al., 2017) to engaging in general media consumption (Menon, 2022).
-As such, although people may approach these platforms with the intention of connecting socially, they may encounter experiences that prioritize other forms of engagement, potentially leaving their social needs unmet.
-This gap highlights the importance of developing a clearer understanding of what it means to support meaningful social connection in online environments.
-While social media offers both risks, such as social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b), and benefits, such as support for marginalized groups (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and opportunities for identity development and self-expression (Davis, 2012; Lee et al., 2022), social connection remains a central motivation for many users.
-Particularly for young people, who increasingly rely on digital platforms for social interaction, it is critical to move beyond treating connection as one of many competing platform goals and toward a more focused understanding of how platforms can meaningfully support users’ social relationships.
+d.]) are designed for diverse user motivations including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022).
+From a uses and gratifications perspective (Katz et al., 1973), users pursue varied goals from building social capital (Phua et al., 2017b) to media consumption (Menon, 2022).
+Users seeking social connection may encounter experiences prioritizing other engagement forms, leaving social needs unmet.
+This gap underscores the need for a clearer understanding of meaningful social connections in online environments.
+While social media presents risks like social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b) alongside benefits like marginalized group support (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and identity development opportunities (Davis, 2012; Lee et al., 2022), social connection remains central for many users.
+For young people increasingly relying on digital platforms, we must move beyond treating connection as one competing goal toward a focused understanding of how platforms can meaningfully support social relationships.
 2.2. Social Technology Design Explorations in HCI
 2.2.1. Designing for Meaningful Social Connections
-HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features.
-Social media platforms such as BeReal (Kim et al., 2024b), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression.
-This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction.
-However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
-At the feature level, efforts to support meaningful communication include designs that encourage effortful communication (Liu et al., 2021; Zhang et al., 2022) and those that foster feelings of genuine connections (Stepanova et al., 2022).
-While these features are important, social media experiences are often holistic, involving a complex interplay of multiple design aspects and user behaviors.
-As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
-Beyond archetypal social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions.
-Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments.
-Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions.
-However, these platforms often are not generally recognized as social media in the generic sense.
-What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear, especially when considering the possibilities offered by emerging technologies such as spatial computing platforms.
-Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
+HCI research has extensively explored designs fostering meaningful social interactions online at both platform and feature levels.
+Platforms like BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) introduced unique affordances emphasizing spontaneity, ephemerality, and creative self-expression.
+This research identified ways to build relationships through features prioritizing authenticity and playful interaction.
+However, users report increasingly toxic interactions as platforms gain popularity, with design choices encouraging network expansion and compulsive behaviors like tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), detracting from meaningful interaction goals.
+Feature-level efforts include designs encouraging effortful communication (Liu et al., 2021; Zhang et al., 2022) and genuine connections (Stepanova et al., 2022).
+However, social media experiences are holistic, involving a complex interplay of multiple design aspects and user behaviors.
+Individual features alone cannot fundamentally alter overarching social media narratives.
+Beyond archetypal social media, recent work investigates spatial social technologies emphasizing immersive, embodied interactions.
+Social VR platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide virtual spaces and embodiment for co-presence and shared experiences, facilitating more personal and tangible interactions than 2D environments.
+Platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend gaming and social networking, enabling virtual environment navigation, avatar customization, and dynamic group interactions.
+However, these platforms are not generally recognized as conventional social media.
+What youth envision as ideal social media at the landscape and feature levels remains unclear, especially considering emerging technologies like spatial computing platforms.
+Understanding these perspectives is crucial for reimagining social media’s future to meet users’ evolving needs.
 2.2.2. Video Games as a Source of Meaningful Social Connection
-Video games are increasingly recognized as valuable environments for fostering meaningful social connection, particularly among youth.
-Large-scale survey research shows that video games serve as a major venue for the development and maintenance of friendships, especially for boys (Lenhart, 2015c).
-Teens not only play games with friends they know in person but also frequently make new friends through online gameplay (Lenhart, 2015a), with more than half of teens reporting they have met at least one friend online (Lenhart, 2015a, b).
-These interactions extend beyond gameplay itself, supporting ongoing social relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
-Research also highlights the social and emotional benefits of gaming, including opportunities for cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014).
-Certain game designs promote prosocial outcomes by encouraging players to help and support others, both in-game and beyond (Gentile et al., 2009).
-Moreover, games are not limited to surface-level entertainment; they can also offer deeper, eudaimonic experiences that involve meaning-making, personal growth, and feelings of connection with others (Daneels et al., 2023).
-These findings suggest that video games can be powerful tools for fostering meaningful social connections, particularly among youth, and highlight the potential for game-like features to enhance social media experiences.
+Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth.
+Large-scale surveys show games serve as major venues for friendship development and maintenance, especially for boys (Lenhart, 2015c).
+Teens play with known friends and make new ones through online gameplay (Lenhart, 2015a), with over half reporting meeting at least one friend online (Lenhart, 2015a, b).
+These interactions extend beyond gameplay, supporting ongoing relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
+Research highlights gaming’s social and emotional benefits, including cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014).
+Certain designs promote prosocial outcomes by encouraging players to help others in-game and beyond (Gentile et al., 2009).
+Games offer deeper, eudaimonic experiences involving meaning-making, personal growth, and connection (Daneels et al., 2023).
+These findings suggest video games can foster meaningful social connections among youth and highlight the potential for game-like features in social media.
 2.3. Recognizing and Overcoming Social Media Design Fixation
-Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design.
-While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking.
-This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
-Social media design is no exception (Sundaram, [n.
+Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented design challenge.
+Though generating divergent ideas is critical for innovation, we often impose unnecessary restrictions.
+This self-limiting process confines design to localized problems rather than exploring broader, optimal solutions.
+Social media design faces similar fixation (Sundaram, [n.
 d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020).
-Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps (Oth, [n.
+Ad-based revenue models incentivize increased time on apps (Oth, [n.
 d.]).
-However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024b) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)).
-These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n.
+Platforms use compulsive strategies like friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)), creating distrustful user communities (Kim et al., 2025).
+Rather than proactively exploring new paradigms of designs and interactions, much recent work on social media design has disproportionately focused on reactively addressing existing harm (Kim et al., 2024b).
+The domination of mainstream platforms along with this reactive approach to design leaves users dissatisfied and reinforces perceptions that all social media experiences are fundamentally the same (Sundaram, [n.
 d.]; Pardes, 2020).
-2.3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
-FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007).
-At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios.
-These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings.
-The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints.
-While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
-According to Dindler and Iversen (Dindler and Iversen, 2007), implementing this method involves three key steps: first, clearly defining the inquiry’s purpose, whether it’s for staging design situations, exploring future ideas, or driving organizational change; second, developing an appropriate narrative that participants are comfortable with while being distinct enough from current practices to encourage new thinking; and third, creating a compelling plot that introduces tension or conflict within the narrative to motivate specific workshop activities.
-Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
-Further, one of the most powerful ways to overcome design fixation is analogical transfer (Novick, 1988; Buchanan, 1992).
-We see video games as a rich source of analogies for social media design, offering unique affordances and interaction patterns that can inspire new approaches to fostering meaningful social connections.
-By examining how video games facilitate social interactions, we can identify design patterns that can be adapted to social media contexts, helping to break free from design fixation and explore novel alternatives.
+Analogical transfer offers one powerful method for overcoming design fixation (Novick, 1988; Buchanan, 1992).
+Therefore, in this paper, we attempt to break out of the design and cognitive fixation that assumes the future of social media is bleak by freeing participants from real-world constraints and beliefs and leveraging elements from fictional settings to serve as sources for analogical transfer.
 3. Method
-We conducted co-design interviews with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms.
-All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n.
-d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers.
-During the interviews, we used the FI method to scaffold participants in generating design ideas.
+We conducted co-design interviews with 23 participants aged 15-24 to examine how we might envision ideal social media platforms free from real-world constraints and design fixation on archetypal platforms.
+All participants had 3D gaming experience and Harry Potter (Rowling, 2015; har, [n.
+d.]) familiarity, a popular young adult fantasy series where some possess magical powers.
+We used the FI method to scaffold participants’ design idea generation.
+All procedures were approved by the university’s Institutional Review Board (IRB).
+3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
+FI is an exploratory co-design method enabling participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007).
+FI creates environments where people imagine beyond present limitations through carefully crafted scenarios incorporating immersive elements like physical artifacts or role-playing activities.
+Participants take generative roles envisioning futuristic or fictional settings.
+The method effectively overcomes design fixation by encouraging analogies from fictional scenarios and free imagination outside real-world constraints.
+Unlike other ideation techniques, FI emphasizes developing futuristic design materials through collaborative designer-user dialogue (Dindler and Iversen, 2007).
+Dindler and Iversen (Dindler and Iversen, 2007) outline three key implementation steps: defining the inquiry’s purpose (staging design situations, exploring future ideas, or driving organizational change); developing an appropriate narrative that participants find comfortable yet distinct enough from current practices to encourage new thinking; and creating compelling plots introducing tension or conflict to motivate workshop activities.
+Unlike traditional role-playing, FI encourages participants to maintain their identities and expertise while operating in fictional contexts, exploring new possibilities through personal motivations and capabilities while circumventing societal constraints.
+We view video games as rich analogy sources for social media design, offering unique affordances and interaction patterns inspiring new approaches to meaningful social connections.
+By examining how games facilitate social interactions, we can identify adaptable design patterns for social media contexts, breaking free from fixation to explore novel alternatives.
 3.2. Material: The Narrative Framework
@@
 Each section contains an empty space for names or details to be added.
-For our study, we selected the Harry Potter world as the narrative framework for our FI method.
-The Harry Potter series follows the journey of an orphaned boy who discovers he is a wizard and, as a result, is invited to attend “Hogwarts School of Witchcraft and Wizardry,” where he forms deep friendships and gradually uncovers his destiny as he clashes with the story’s antagonist.
-Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI interviews.
-The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement.
-The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S.
+We selected the Harry Potter world as our FI narrative framework.
+The series follows an orphaned boy who discovers he is a wizard and attends “Hogwarts School of Witchcraft and Wizardry,” forming deep friendships while uncovering his destiny and clashing with the antagonist.
+Among numerous magical worlds, Harry Potter proved particularly suitable for FI interviews.
+Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement.
+By 2023, 98% of US youth aged 18-27 identified as casual or avid fans, with half describing themselves as avid fans (Share of Harry Potter fans U.S.
 2023, [n.
 d.]).
-Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K.
+The book series sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K.
 Rowling’s Harry Potter and the Sorcerer’s Stone, [n.
-d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
-The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells.
-Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies.
-Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building.
-The characters’ experiences or life transitions, which parallel real-world moves between schools or locations, allow for exploring how social technologies might support relationship maintenance across changing contexts.
-The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
+d.]) and was translated into over 80 languages by 2020 (The Rowling Library, 2020).
+Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints.
+Young protagonists and coming-of-age narrative align with our focus on youth identity and relationship building.
+Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts.
+The series’ emphasis on community and belonging suits examining youth identity expression and self-presentation in virtual spaces.
 3.3. Procedure
-The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI method.
-The interviews averaged around 90 minutes.
-We began the interview by asking which 3D games they had played and what they liked about those experiences.
-We also asked questions about Harry Potter, such as their favorite book in the series.
-This served to break the ice and help them start immersing themselves in the Harry Potter world.
-Following this introduction, we presented the FI narrative: participants were asked to envision ways to support social connection within the Hogwarts world as a non-Wizard (i.e., “Muggle” (mug, [n.
-d.])) trying to stay connected with everyone across Hogwarts and the Muggle world.
-We then shared a Miro (mir, [n.
-d.]) board (Figure 1) where participants could sort their friends and family into categories inspired by the Harry Potter series, such as fictional houses (e.g., Gryffindor, Hufflepuff), teams (e.g., Quiddich Team members), and old friends (e.g., muggles), reflecting fictional houses, teams, and communities from the story.
-This activity encouraged them to think of their relationships within the Harry Potter context and was intended to further enhance immersion in the fictional design space.
-Next, we asked general questions about their experiences with social media, focusing on what they found meaningful or meaningless and their vision for an ideal social media platform.
-During the design phase, participants were prompted to describe the capabilities their ideal magical social media would have and what aspects of their lives they would choose to share.
-Throughout this phase, we sought to keep participants immersed in the narrative, referring to existing platforms as “Muggle social media” and maintaining consistent engagement with the Hogwarts setting, as outlined by the fictional inquiry guidelines.
-Finally, we concluded by asking participants to reflect on how the social media they envisioned for Hogwarts compared to Muggle social media.
-The semi-structured interview protocol is available in Appendix A.
+The first author conducted all interviews over Zoom, with the second author joining one session to validate the FI method.
+Interviews averaged 90 minutes.
+We began by asking about 3D gaming experiences and Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
+We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (mug, [n.
+d.])) connecting across Hogwarts and Muggle worlds.
+We shared a Miro (mir, [n.
+d.]) board (Figure 1) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles).
+This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion.
+We then asked about social media experiences, focusing on meaningful versus meaningless aspects and ideal platform visions.
+During the design phase, participants described their ideal magical social media capabilities and what life aspects they would share.
+We maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines.
+On the other hand, we encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform.
+Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take?
+(Is it a magical mushroom?
+A secret room?
+A creature?)” We concluded by asking participants to compare their envisioned Hogwarts social media with Muggle social media.
+The semi-structured protocol is in Appendix A.
 3.4. Participants and Recruitment
-We used a combination of purposive and convenience sampling to recruit participants for our study.
-The first and second authors shared the recruitment materials (that included a link to a screener survey) on their social media accounts, while the first author also distributed flyers across our institution’s campus.
-Participants were required to be between 15 and 24 years old, in accordance with UN (United Nations, [n.
+We used purposive and convenience sampling for recruitment.
+The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
+Participants needed to be 15-24 years old per UN (United Nations, [n.
 d.]) and WHO (who, [n.
-d.]) definitions for youth, and to consent to be video or audio recorded during a 90-minute interview.
-Video recording was requested following the FI method guidelines, enabling us to analyze gestures if necessary.
-We included both teenage and young adult participants (ages 15 to 24) to capture perspectives from distinct life stages, as these groups often navigate significant transitions and share a focus on identity exploration, self-presentation, and relationships.
-They also had to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox).
-While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
+d.]) youth definitions and consent to video/audio recording during 90-minute interviews.
+Video recording followed FI guidelines, enabling gesture analysis if needed.
+We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
+Participants needed sufficient Harry Potter familiarity (agreeing with “I remember some details about the plot, characters, and magical aspects”) and prior 3D gaming experience (e.g., Minecraft, Roblox).
+We did not require extensive knowledge of Harry Potter because we felt that would be an unnecessary bias that is not particularly helpful in envisioning new social media design.
+While this made it somewhat difficult for a majority of participants to remember specific artifacts from the Harry Potter world, they knew enough to recognize the complete and extensive freeing extent of the magical powers in that setting.
+Though we imposed no social media experience constraints, all participants had some social media engagement.
 Table 1.
@@
 P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
-A total of 84 individuals completed the screener survey.
-We reached out to the 48 respondents who met the recruitment criteria, and 23 participants provided consent and completed the interview procedure.
-The mean age of the participant group was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) identifying as women, 3 (13.0%) as men, 2 (8.7%) as non-binary, and 3 (13.0%) that were gender fluid.
-The full list of participant information is available in Table 1.
+Eighty-four individuals completed the screener survey.
+We contacted 48 who met criteria; 23 provided consent and completed interviews.
+Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants.
+While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel, 2019), we note that our participant demographics skew toward women.
+Full demographic information appears in Table 1.
+(Lenhart, 2015b, c)
 3.5. Data Analysis
-We analyzed the interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides a systematic yet adaptable framework for interpreting qualitative data.
-The method’s independence from pre-existing theoretical frameworks made it particularly suitable for our study, given our study’s exploratory focus on youth-envisioned social media design through the FI method.
-It enabled us to effectively capture and analyze the diverse perspectives shared by our participants.
-Our analysis began with the first two authors independently coding the same two transcripts line by line.
-During this phase, the codes were descriptive and closely adhered to the data.
-The authors convened to discuss and resolve any discrepancies in their coding.
-Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024).
-After four coding iterations, we reached saturation in our codes.
-Using this final set of codes, the first author proceeded to code all the transcripts.
-All versions of the codes and the final codebook are available in the supplementary materials.
+We analyzed interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides systematic yet adaptable qualitative data interpretation.
+Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
+Since not all participants agreed to have their videos on during interviews, we did not conduct video analysis of gestures that participants sometimes used when describing their ideas.
+Additionally, drawing was not part of our protocol because we have experienced in prior interviews for other studies that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination.
+All data therefore consisted of transcripts from verbal interviews.
+Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data.
+Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atl, 2024).
+After four iterations, we reached code saturation.
+The first author then coded all transcripts using the final code set.
+All code versions and the final codebook are in supplementary materials.
 3.6. Ethical Considerations
 Given J.K.
-Rowling’s controversial statements regarding transgender individuals (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K.
+Rowling’s controversial transgender statements (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K.
 Rowling has made statements regarding transgender individuals that many find harmful and discriminatory.
@@
 Thank you for your understanding and for helping us advance this important research.
-in our recruitment material and screener survey and reiterated our position during the interview procedure.
+in recruitment materials and surveys, reiterating our position during interviews.
 4. Results
-Informed by their experiences with 3D games and familiar media worlds, our participants imagined forms of social interaction that extend beyond what is typically supported on mainstream platforms like Instagram.
-Through the speculative frame of Social Media at Hogwarts (SMH), they explored how digital environments might enable more embodied, relational, and emotionally meaningful forms of connection.
-Their visions converged around six interrelated themes, which, collectively, contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs, ultimately expanding their perception of the horizons for social connection on such platforms.
+Drawing from their experiences with 3D games and familiar media worlds, participants imagined social interactions extending beyond mainstream platforms like Instagram.
+Through the speculative frame of exploring a “new magical ‘thing’ for social communication” for Hogwarts students, they envisioned how digital environments might enable more embodied, relational, and emotionally meaningful connection.
+Their visions converged around six interrelated themes that collectively contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs.
+We refer to this synthesized concept as “Social Media at Hogwarts,” or SMH, throughout this paper.
 4.1. Intuitive Social Navigation
-Youth described seeking interactions that feel natural, embodied, and socially meaningful, reflecting the way people navigate relationships and spaces in everyday life.
-They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape how they connect with others, express themselves, and manage privacy in intuitive and socially familiar ways.
+Youth sought interactions that feel natural, embodied, and socially meaningful, reflecting how people navigate relationships and spaces in everyday life.
+They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy management in intuitive, socially familiar ways.
 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
-Participants described how embodied, immersive qualities such as realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space that fosters presence, emotional connection, and socially meaningful interaction.
-In contrast, current social media (which we will refer to as “Muggle Social Media” or MSM for short), was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 similarly noted that MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” P03 described SMH as “more immersive and more awesome” than MSM, which they saw as inherently limited “because it’s 2D.” P16 emphasized that meaningful social connection requires “talk[ing] to you and see[ing] your face,” something they believed SMH could make possible.
-For others, the fulfillment came from the embodied experience of physical actions.
-P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They also added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
-P03 similarly reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” As P06 put it, SMH “feels more real” because you “walk around a physical world” and “feel like you’re actually in the area of these people.”
+Participants described how embodied, immersive qualities like realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space fostering presence, emotional connection, and meaningful interaction.
+In contrast, current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation, explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
+For others, fulfillment came from embodied physical actions.
+P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing,” explaining “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
+P03 reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
 This sense of presence was felt even in seemingly passive actions. As P01 explained,
@@
 I don’t know if that makes sense but.” (P01)
-For P01, 3D interaction not only supported more embodied engagement but also introduced a felt presence that lingered, even when no direct interaction took place.
+For P01, 3D interaction supported more embodied engagement and introduced a felt presence that lingered, even without direct interaction.
 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
-Participants proposed that using spatial and visual cues modeled after real-world environments could make managing access and privacy feel more intuitive, socially natural, and contextually appropriate.
+Participants proposed using spatial and visual cues modeled after real-world environments to make managing access and privacy feel more intuitive, socially natural, and contextually appropriate.
 P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
@@
 If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
-While managing visibility across multiple rooms might seem more complex, participants believed this approach would feel intuitive and natural.
-P11 suggested that spatial contexts could lower the efforts required for privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
+While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive and natural.
+P11 suggested spatial contexts could reduce privacy control efforts, stating “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
 4.2. Shared Experiences
-Participants emphasized that shared experiences, whether through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
+Participants emphasized that shared experiences, through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
 4.2.1. Fostering Shared Memories Through Collaborative Activities
-Participants described shared activities as a meaningful foundation for building friendships.
+Participants described shared activities as meaningful foundations for building friendships.
 P06 contrasted casual solo games with more immersive, co-experienced ones.
-For example, while playing a simple emoji game was described as “just something fun we can both do,” it lacked a true shared experience since “we don’t actually see each other’s screens” and would just “screenshot our scores and send them to each other.” By contrast, they described Roblox as a space that fosters shared moments that become lasting memories, like “do you remember that time when that kid in the voice chat said this,” or “when we got scared by that jump scare.” They explained, “it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
-Similarly, P12 described the value of collaborative projects, such as “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting it with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than simply “posting” content individually.
-The value of shared experiences and memories was not limited to completing tasks or reaching goals.
-Even simple activities like walking together were seen as meaningful opportunities for connection.
-P03 imagined moments where you could just “walk with your friend,” as if taking a “nature walk” (P14).
-P21 echoed this, describing how walking itself could spark memorable conversations and shared moments: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
+While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” they described Roblox as fostering lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
+P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
+The value of shared experiences extended beyond completing tasks.
+P21 described how walking could spark memorable conversations: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
-Participants also saw low-pressure, ambient forms of social interaction, such as co-existing in shared spaces or listening to music together, as natural and organic ways to build connections and shared memories.
-P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background.
-They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations.
-P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.” Even “mindless scrolling” became meaningful when shared.
-P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
-Music would often facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They came up with the idea reflecting on their own experience:
+Participants saw low-pressure, ambient social interaction, like co-existing in shared spaces or listening to music together, as natural ways to build connections and shared memories.
+P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual study room with Lo-Fi music.
+These spaces would mirror Discord servers’ 24/7 activity, offering comforting presence even without direct conversation.
+Even “mindless scrolling” became meaningful when shared, as P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
+Music often facilitated these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs,” reflecting on their experience:
 “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.
@@
 It was just nice.” (P19)
-They added that music could also serve as a “conversation starter” while also providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music.
+They added that music could serve as a “conversation starter” while providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music.
 You can just listen to the music.”
 4.3. Communal Ambience
-SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison.
-P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
-Participants described several design features, such as neighborhood structures, shared gathering spaces, and stronger boundary controls, that could help cultivate this kind of communal, low-pressure environment.
+SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison.
+P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features fostering organic, intentional, personalized interactions and closer friendships.
+They described several design features–neighborhood structures, shared gathering spaces, and stronger boundary controls–that could cultivate this communal, low-pressure environment.
 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
-Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to naturally foster social interaction and relationship-building.
+Participants wanted SMH’s 3D world to mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster social interaction and relationship-building.
 P22 described in detail:
@@
 I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
-Youth envisioned SMH as also reflecting the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties.
-Specifically, some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction.
+Youth envisioned SMH reflecting real-world relationship dynamics, using physical distances within the landscape to represent social tie strength.
+Some participants envisioned SMH resembling a college campus or neighborhood, where physical closeness naturally fosters interaction.
 One participant elaborated:
@@
 I think that would help build a real sense of community.” (P03)
-Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
-This design approach of aligning physical landscapes with emotional closeness was seen as a way to streamline interactions and focus on meaningful relationships.
+P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach.
+This approach of aligning physical landscapes with emotional closeness would streamline interactions and focus on meaningful relationships.
 P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
-Participants imagined that in addition to private homes and personal spaces, SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating opportunities for connection.
-P05 described these areas as “neutral spaces”, not tied to any specific activity or relationship status, where groups could gather for a range of purposes.
-They imagined “some places to sit, and some kind of flat table-like surface” that could support casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces, they explained, would function as “a third, neutral territory” where people could interact more organically, without the friction of managing invitations or switching platforms.
-Participants described shared social spaces within neighborhoods, such as common rooms, courtyards, or clubhouses, where residents could casually interact outside their personal spaces.
-P01 described leaving one’s home to gather in a “courtyard or game room” at the center of the neighborhood.
-Drawing on examples from Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms that cater to different interests, such as a Quidditch field or library, creating opportunities for people to meet others through shared activities.
-P09 similarly imagined a “virtual dining room” that would allow people to casually talk while doing something familiar, like eating together.
-Unlike the more segmented house-based accommodation at Hogwarts, this dining space was described as open and inclusive, allowing people to connect across social groups.
-Others described third places located in public areas that supported casual, serendipitous interactions.
-P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look!
-This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
+Participants imagined SMH neighborhoods including shared third places where people could casually gather, lowering social barriers and creating connection opportunities.
+P05 described these as “neutral spaces” not tied to specific activities or relationships, where groups could gather for various purposes.
+They imagined “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
+Participants described shared neighborhood spaces like common rooms, courtyards, or clubhouses where residents could casually interact outside personal spaces.
+P01 described gathering in a “courtyard or game room” at the neighborhood center.
+Drawing on Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities.
+P09 imagined a “virtual dining room” allowing casual conversation while eating together.
+Unlike Hogwarts’ segmented house-based accommodation, this dining space was open and inclusive, enabling cross-social group connections.
+Others described public third places supporting casual, serendipitous interactions.
+P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
-To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized the importance of filtering out irrelevant content, such as celebrity drama or political commentary, and prioritizing updates from close friends and trusted communities.
+To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends and trusted communities.
 P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
 We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
-Several participants advocated minimizing celebrity influence.
-P02 stated, “I don’t want celebrity stuff… I don’t want any celebrity drama,” while P01 suggested removing celebrities altogether if the platform prioritized personal enjoyment.
-If included, they proposed isolating them to a “celebrity zone.” Prioritizing posts from close friends was a key preference.
-P07 explained, “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH offering greater control over interactions and content, noting it would be “more designed for people you actually know… like your friends, classmates, or Quidditch team,” rather than celebrities or acquaintances.
-P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
+Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference for close connections: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH being “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
-Participants described how themed, socially grounded conversational spaces designed for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
+Participants described how themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
 “Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread.
@@
 I find it’s way easier to get to know people a lot better.” (P22)
-Participants described how the design and feel of conversational spaces could make social interactions on SMH feel safer, more personal, and more inviting than on existing platforms.
-Many valued the idea of walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to a broad, undefined audience.
+Participants described how conversational space design could make SMH interactions feel safer, more personal, and more inviting than existing platforms.
+Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences.
 P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations.
@@
 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
-To ensure SMH feels like a safe, personal environment, participants stressed the need for more comprehensive boundary controls beyond privacy tools found on existing platforms that they perceived as limited and inconsistent.
-P22, for example, explained that even with contact syncing turned off on Instagram, the platform still suggested their Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they were trying to avoid, based on mutual follow relationships.
-They stressed the need for stronger, more thorough blocking controls, including the ability to block “alternate accounts” and fully remove any trace of a blocked person’s presence, explaining, “I don’t wanna be able to tell that the person I blocked is even around me.” P14 similarly pointed out that even private accounts on Instagram still expose personal content, like Story Highlights, to “a random person”, revealing perceived gaps in platform-level privacy protections (Bojkov, 2019).
-Participants expressed a desire for stronger control over their own visibility, imagining features that would let them opt out of being seen or discovered entirely.
-P03 suggested a feature like an “invisibility cloak” for moments when users “don’t want anyone to see you.” P07 proposed a “super private” profile mode where users “can’t even be searched… like you’re not even on the app” unless they choose to make themselves visible.
-Preventing unsolicited contact from strangers was another recurring priority.
-P22 proposed a model inspired by Discord, where only people you’ve already interacted with can contact you, and others “can’t even be discovered unless you have something in common.” This idea resonated with P13, who argued that new connections should require mutual consent, suggesting that people “should only be able to add you if you meet in person and exchange a QR code.” They shared a personal experience of being able to find a stranger online based on just their first name, reflecting that “it’s too easy to find people on social media.”
+To ensure SMH feels safe and personal, participants stressed needing comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools.
+P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (Huang and Vitak, 2022)) to people they wanted to avoid, based on mutual follows.
+They stressed needing stronger, more thorough blocking controls, including blocking “alternate accounts” and fully removing blocked person traces, explaining “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (Bojkov, 2019).
+Participants wanted stronger control over their visibility, imagining features letting them opt out of being seen or discovered entirely.
+P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
+Preventing unsolicited stranger contact was another priority.
+P22 proposed a Discord-inspired model where only people you’ve interacted with can contact you, while P13 argued people “should only be able to add you if you meet in person and exchange a QR code,” reflecting that “it’s too easy to find people on social media.”
 4.4. Nuanced Self-Presentation
 Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14).
-This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23).
-P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting the full range of who they are, without the pressures of curation or performance.
+This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23).
+P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting their full range without curation or performance pressures.
 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
-Participants consistently described customizable avatars and personal spaces as meaningful tools for self-expression, offering more creativity, personality, and authenticity than the static profiles or limited settings found on MSM.
-Unlike MSM, which typically offers only “two modes” (P10), i.e., “light and dark” (P10), participants imagined a much richer and more immersive experience.
-For example, P21 envisioned a fully customizable world where users could build their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” These spaces could be shared with friends as a way of “showing a little part of yourself.” Similarly, P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture,” to reflect their personal interests.
-Customizable “3D avatar[s]” (P11) and profiles were also seen as key ways of showing individuality.
-P22 reflected on the early days of Myspace (mys, [n.
-d.]), describing how they enjoyed coding their profile layouts and changing “the font and really make it reflect me.” Even small details, like choosing a username, were described as ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted that such forms of expression are often “mainly for you,” not just for public display.
+Participants consistently described customizable avatars and personal spaces as meaningful self-expression tools, offering far more creativity than MSM’s limited settings.
+P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
+Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality.
+P22 reflected on early Myspace (mys, [n.
+d.]), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames were ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display.
 P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right?
 So I think that’s part of you.”
-Beyond visual customization, participants imagined SMH supporting more ambient, indirect forms of expression, offering subtle cues about what someone is doing, feeling, or thinking without needing to make a formal post.
-For instance, P01 described wanting their avatar to reflect real-time states, like the outfit they are wearing or the song they are listening to, so that friends could intuitively notice things like, “‘Oh.
-that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ Similarly, P10 imagined that friends could pick up on changes in their interests simply by observing how they update their virtual space over time, providing “an insight into… what I’m thinking.”
-Several participants highlighted the contrast between the flat, compressed feel of 2D platforms and the immersive qualities of 3D environments.
-P10 noted that in 3D, others can see “as much stuff as you can see in 3D,” making it feel richer and more complete.
-P16 added that VR creates a sense of presence that feels more like meeting “real people… compared to the profile pictures,” because you can pick up on their “mannerisms” and notice details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
-Participants also described how these kinds of choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests.
+Beyond visual customization, participants imagined SMH supporting ambient, indirect expression, offering subtle cues about what someone is doing, feeling, or thinking without formal posts.
+P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh.
+that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
+Several participants highlighted the contrast between 2D platforms’ flat, compressed feel and 3D environments’ immersive qualities.
+P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
+Participants described how these choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests.
 As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation.
@@
 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
-Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides of themselves across various spaces and relationships, and reducing the pressure to maintain a single, polished identity.
-Rather than being limited to a single, polished identity, as is often the case on MSM, participants wanted ways to represent the full range of their interests, identities, and relationships.
+Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships while reducing pressure to maintain a single, polished identity.
+Rather than being limited to one polished identity, as on MSM, participants wanted ways to represent their full range of interests, identities, and relationships.
 P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things.
@@
 P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
-4.4.3. Sharing Everyday Activities and Lived Experiences to Deepend Understanding and Connection
-Participants valued the ability to share casual, everyday moments and rich, embodied memories, filling the gaps between the highlights usually shared on MSM into a more natural, meaningful extension of daily life and personal connection.
+4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
+Participants valued sharing casual, everyday moments and rich, embodied memories, filling gaps between MSM’s typical highlights to create a more natural, meaningful extension of daily life and personal connection.
 P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment.
 For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
-Participants also imagined SMH supporting the sharing of more personal, embodied memories as a way to deepen understanding and reduce misinterpretation.
+Participants imagined SMH supporting personal, embodied memory-sharing to deepen understanding and reduce misinterpretation.
 P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n.
@@
 P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
-Some participants imagined that these shared memories could include sensory or emotional details, not just visuals.
+Some participants imagined these shared memories including sensory or emotional details beyond visuals.
 P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication.
@@
 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
-Participants emphasized that flexible, contextual privacy is essential for supporting authentic self-presentation, allowing people to manage who sees what, when, and how, without feeling overexposed or misrepresented.
+Participants emphasized that flexible, contextual privacy is essential for authentic self-presentation, allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented.
 They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life.
@@
 4.5. Intentional Engagement
-Participants described their growing frustration with how current social media platforms prioritize passive consumption over meaningful connection.
-In response, they envisioned SMH as a space that promotes intentional, socially rewarding interactions.
-This section explores four key themes that reflect their desire to refocus social media on authentic connection, purposeful engagement, and meaningful use of time.
+Participants described growing frustration with current social media platforms prioritizing passive consumption over meaningful connection.
+In response, they envisioned SMH as a space promoting intentional, socially rewarding interactions.
+This section explores themes reflecting their desire to refocus social media on authentic connection, purposeful engagement, and meaningful time use.
 4.5.1. Encouraging Deliberate Navigation and Content Discovery
-Participants imagined SMH as a space that encourages deliberate, effortful discovery, replacing the passive, algorithm-driven feeds of MSM with experiences that require users to actively choose what and who they engage with.
-P22 described their frustration with Instagram, explaining that “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage more intentional, purposeful engagement, both when browsing content and when navigating the virtual world.
+Participants imagined SMH encouraging deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with.
+P22 described frustration with Instagram, explaining “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage intentional, purposeful engagement when browsing content and navigating the virtual world.
 P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful.
 Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
-Participants also imagined features that would prompt users to actively choose what they wanted to engage with before entering a space.
+Participants imagined features prompting users to actively choose what they wanted to engage with before entering a space.
 P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
-Physical movement through SMH’s world was also seen as key to this intentionality.
-Participants valued having to walk, explore, and navigate to find people, spaces, or activities, turning even small actions into social moments of presence and care.
+Physical movement through SMH’s world was key to this intentionality.
 P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
-This desire for intentional interaction extended to how users move through the platform itself.
-Participants imagined physical navigation, such as walking or traveling through virtual space, as a way to make engagement feel more meaningful.
-P02 explained that they would prefer slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions.
-I want everything to be very thought through.” While features like “spells” (P06) or the “Marauder’s Map” (P21) were imagined as convenient options, participants valued the ability to walk and explore, turning simple actions like visiting a friend or browsing a shop into shared social moments.
-P03 explained that this kind of embodied interaction feels “more connected than regular social media” (P03) because it requires active presence rather than passive scrolling.
+P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions.
+I want everything to be very thought through.” While convenient features existed, participants valued walking and exploring, turning simple actions into shared social moments.
 P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them.
@@
 It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
-Participants believed that this kind of effortful navigation would naturally discourage casual lurking or influencer-driven engagement, shifting the platform’s focus toward genuine connection.
-As P05 put it, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
+P05 believed this effortful navigation would naturally discourage superficial engagement: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
-Beyond how users navigate content, participants also emphasized the importance of shifting the platform’s core purpose from endless consumption toward meaningful social connection.
-Participants expressed frustration with how current platforms promote mindless, passive consumption, and described their hopes for SMH to instead foster more meaningful, intentional connections through active social engagement.
-They described frustration with how easily time slips away on platforms like Instagram and TikTok without leaving them feeling unfulfilled.
-P10 noted that social media often feels “meaningless” when they “didn’t learn something,” and P18 reflected on the regret of “mindlessly scrolling.” P14 similarly described spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
-A major contributor to this feeling was the infinite scrolling design of MSM.
-P11 explained that “there’s always more content on social media,” which fuels the fear of missing out and the urge to keep scrolling.
-P14 critiqued TikTok for exposing users to “clickbaity” and often low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
-In contrast, participants described active, social engagement, even if casual, as far more rewarding.
-P16 shared that “anytime not scrolling” but instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this, noting that meaningful use comes when they are “actually talking to someone,” especially to stay connected with people “not in close proximity.” They reflected that while users might spend more time on a platform like SMH, that time would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to passively watching “50-second TikTok videos of somebody making slime for like an hour.”
-Participants also contrasted passive consumption with active participation in interactive environments.
-P10 explained that while doomscrolling feels unproductive, activities like playing Minecraft feel different because they offer “progress” that can be “saved and come back to.” P06 echoed this sentiment, noting that they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
-Reflecting on SMH’s potential to create this kind of active, engaging experience, P09 acknowledged that “time spent might not be mitigated,” but emphasized that “the way you’re spending it” would make the difference.
-They explained that SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting that meaningful engagement, rather than passive consumption, was what ultimately mattered for them.
+Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection.
+They expressed frustration with current platforms promoting mindless, passive consumption and hoped SMH would foster meaningful, intentional connections through active social engagement.
+They described frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled.
+P14 captured this sentiment, describing spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
+MSM’s infinite scrolling design contributed to this feeling.
+P11 explained “there’s always more content on social media,” fueling fear of missing out and urges to keep scrolling.
+P14 critiqued TikTok for exposing users to “clickbaity” low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
+In contrast, participants described active, social engagement as far more rewarding.
+P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
+Participants contrasted passive consumption with active participation in interactive environments.
+P10 explained that while doomscrolling feels unproductive, activities like Minecraft offer “progress” that can be “saved and come back to.” P06 noted they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
+Reflecting on SMH’s potential, P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference.
+They explained SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting meaningful engagement, rather than passive consumption, ultimately mattered.
 4.6. Playfulness and Whimsy
-Participants described SMH as a platform that feels youthful, playful, and deeply immersive, offering not just entertainment but opportunities for creative self-expression, social bonding, and emotional well-being.
-By embracing whimsy and delight, SMH was seen as making social media feel more personal, less pressured, and more meaningful.
+Participants described SMH as a platform feeling youthful, playful, and deeply immersive, offering entertainment plus opportunities for creative self-expression, social bonding, and emotional well-being.
+By embracing whimsy and delight, SMH would make social media feel more personal, less pressured, and more meaningful.
 4.6.1. Embracing Creative Joy and Playful Self-Expression
-Participants imagined SMH as a platform that transforms social media into a space for creative joy and playful self-expression, offering users the freedom to experiment with identity, explore personal interests, and connect with others through creativity.
-Several participants described how SMH offered far more than static profiles or simple feeds.
-P06 reflected that “everyone likes sharing parts of themselves,” appreciating how the platform allows users to shape not just what they say, but how they appear and interact with others.
+Participants imagined SMH transforming social media into a space for creative joy and playful self-expression, offering freedom to experiment with identity, explore personal interests, and connect through creativity.
 P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
-For some, this act of personal customization felt like a meaningful process of identity exploration.
+For some, personal customization felt like meaningful identity exploration.
 P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression.
@@
 That’s where the fun aspect comes in, really, for me.”
-This spirit of creativity extends beyond the individual to the wider environment and community.
+This creativity extends beyond individuals to the wider environment and community.
 Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa!
@@
 4.6.2. Designing Engaging and Playful Privacy Mechanisms
-Participants imagined privacy not as a tedious and confusing task but as part of the playful experience, envisioning interactive, game-like mechanisms that turn boundary-setting into something fun, discoverable, and socially meaningful.
-P11 envisioned systems where users could control access to spaces through gestures like a “secret knock” or hidden objects that revealed “secret passages” accessible only to trusted friends.
-Others suggested knowledge-based gates, such as “secret handshakes” (P21) shared among friends.
-Drawing inspiration from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces only when correctly summoned.
-These features were described not just as functional but as adding delight and novelty to social interaction.
-P05 compared them to “little Easter eggs,” remarking that “there’s that little kid part of everyone” that makes such features enjoyable.
-They contrasted these with conventional privacy settings on mainstream social media, which they described as “a wall of settings” that often feel confusing or uninviting.
+Participants imagined privacy not as tedious and confusing but as part of playful experience, envisioning interactive, game-like mechanisms turning boundary-setting into something fun, discoverable, and socially meaningful.
+P11 envisioned controlling space access through gestures like “secret knock” or hidden objects revealing “secret passages” for trusted friends.
+Others suggested knowledge-based gates like “secret handshakes” (P21) shared among friends.
+Drawing from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces when correctly summoned.
+P05 compared these features to “little Easter eggs,” remarking “there’s that little kid part of everyone” that makes such features enjoyable, contrasting them with mainstream social media’s “wall of settings” that often feel confusing.
 4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
 Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities.
-P01 questioned the need for phones at all, asking, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting “holographic screens, like Jarvis in Iron Man,” or activating “panels” and “big screens” through spells.
-Others suggested alternative magical devices.
-P15 preferred VR glasses, describing them as “more natural,” “easier,” and “more fun.” Participants also imagined objects like a “makeup mirror” that could serve as a portal, a “watch” that immerses users into SMH, or a “mirror glass” for video calls or summoning friends with spells like accio.
+P01 questioned the need for phones at all, envisioning wands projecting “holographic screens, like Jarvis in Iron Man,” or activating interfaces through spells.
+Participants also imagined alternative magical devices: VR glasses, a “makeup mirror” portal, or a “watch” that immerses users into SMH.
 These magical affordances were not only imagined as novel interfaces but also as ways to make extraordinary experiences more accessible.
@@
 4.7. Expanding Horizons
-Given the speculative nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what interactions on social media platforms could embody.
-Therefore, while they acknowledged potential challenges, such as the platform being over-immersive or distracting, they overwhelmingly viewed these downsides as worthwhile trade-offs for its many benefits.
-Overall, participants described SMH as an inspiring alternative or expansion of existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional.
+Given the speculative FI co-design workshop, SMH embodies youth aspirations and values for social media platform interactions.
+While acknowledging potential challenges like over-immersion or distraction, they viewed these downsides as worthwhile trade-offs for its benefits.
+Participants described SMH as an inspiring alternative to existing social media, offering self-expression opportunities and deeply personal, intentional connection.
 They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
 4.7.1. Breaking Free from Platform Monotony with Novel Interactions
-Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become.
-For example, P06 shared how SMH combines several desirable aspects of social media in a cohesive way, creating an environment focused on meaningful interactions: “You get friends who you can talk to, keep up with, and stay connected to.
+Participants saw SMH as offering better user experiences, departing from existing platforms’ repetitive structures, and inspiring hope for social media’s future.
+P06 appreciated how SMH combines desirable aspects: “You get friends who you can talk to, keep up with, and stay connected to.
 And you get things you can customize.
-You kind of get the whole package there.
-I kind of like all of those here.”
-Participants also appreciated how SMH offered a break from the monotony of existing platforms.
+You kind of get the whole package there.”
+Participants appreciated SMH’s break from existing platform monotony.
 P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns.
 P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions.
-Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram.
-P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “novel,” like a “breath of fresh air.”
-For some participants, SMH’s novelty lay in giving themselves permission to prioritize enjoyment over seriousness.
+P17 envisioned SMH as “more interactive, kind of like if Discord and Instagram were combined,” enjoying Discord’s interactions with Instagram’s profiles and posting ability.
+P22 highlighted limited customization on other platforms, explaining “I just have a profile picture.” On SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” making the experience feel “novel,” like a “breath of fresh air.”
+For some, SMH’s novelty lay in prioritizing enjoyment over seriousness.
 P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun.
-For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
+For example, P14 initially preferred a 2D version but shifted to 3D, reasoning:
 “Yeah, [I would prefer the 3D version] actually 100%.
@@
 P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms.
-P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness.
-P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
+P06 observed that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” compared to “traditional ones.” The platform also evokes nostalgia and happiness.
+P15 remarked “Having this app would actually make them feel happy… with nostalgia,” anticipating SMH would be “a super big hit” with teenagers due to its playful, engaging nature.
 4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
-Through the FI process, participants realized they were not limited by the structures of existing platforms.
-Instead, they felt empowered to imagine social technologies that center joy, creativity, and authentic human connection, qualities they had once believed were beyond the reach of digital spaces.
+Through the FI process, participants realized they weren’t limited by existing platform structures.
+Instead, they felt empowered to imagine social technologies centering joy, creativity, and authentic human connection–qualities they had once believed beyond digital spaces’ reach.
 P10 pointed out that overcoming the design fixation around social media is crucial:
@@
 In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
-Figure 2 captures how participants described their vision for ideal social media as something that pulls benefits from both types of platforms they already use.
-On one side are social networking and broadcasting platforms like Instagram, Facebook, and TikTok, which make it easy to stay socially visible but often feel shallow or performative.
-On the other side are platforms like FaceTime, Minecraft, Discord, and Roblox, where youth experience more organic and engaging connections, but where social interaction is often secondary to other goals like gaming or productivity.
-Our participants imagined “Social Media at Hogwarts (SMH)” as a platform that blends these strengths, creating a space where meaningful social connection effectively surfaces not just a peripheral effect, but a key focus of the platform.
+Our analysis shows that youth imagine future social media not as a rejection of current platforms but as a reconfiguration of their strengths.
+They want the visibility and reach of social networking and broadcasting platforms while also craving the intimacy and authenticity found in spaces like games, voice chat, or small-group apps.
+Figure 2 illustrates this blended vision: an aspirational middle ground where meaningful connection is not an afterthought but the central purpose.
+Building on this insight, the following subsections trace how participants moved from critiquing existing platforms toward articulating design goals, culminating in our proposal of “spatial integrity” as a framework for supporting deeper forms of online connection.
+The relevance of these findings has become increasingly apparent with recent technological developments.
+With the introduction of technologies like Genie 3 and similar platforms, the fictional world medium of remote social connection that participants envisioned is barely fictional.
+The 3D, open-world, and fully customizable platforms for self-expression and organic social connection that emerged from our workshops now represent concrete possibilities for the near future, much nearer than we anticipated when we began this study.
+This technological trajectory underscores the timeliness of our findings and their potential to guide the development of relationship-supporting technologies that leverage these emerging capabilities (DeepMind, 2025).
+5.1. From Reactionary Designs to Aspirational Design Goals
+There are societal concerns about social media’s impact on youth well-being (Oth, 2023).
+Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023).
+This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
+However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality.
+Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023).
+At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024).
+This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connections that youth actually seek.
+Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
+However, in our earlier work with this youth population, we noticed several recurring challenges that shaped how participants thought about social media.
+First, there was a strong design fixation on social media design around existing apps, both in terms of interface, such as small 2D mobile screens, endless scrolling, and uniform templates, and in terms of user experience, where visibility depends on maintaining large networks of friends and followers, often at the expense of authenticity and comfort.
+Second, participants showed a cognitive fixation when imagining the future of social media, struggling to see it as anything beyond its currently perceived role as a convenient but limited tool for remote social connection.
+Third, these fixations were compounded by generally negative experiences with social media, which led many in earlier interviews to propose reactionary fixes, like stronger privacy, the removal of infinite scrolls, or eliminating general media content such as Reels.
+Such cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
+To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting.
+The procedure was intended to serve three purposes: breaking participants free from these biases, learned helplessness, and pessimism by encouraging them to think beyond the “app” or “platform” on mobile devices; equipping them with prompts that licensed thinking beyond perceived technical and practical constraints; and creating conditions where they could explore proactive and aspirational visions of social interaction rather than simply reacting to existing frustrations.
+Our chosen method prioritized uncovering underlying desires and needs that young people associate with meaningful remote social interaction over producing concrete feature lists.
+In line with the goals of co-design processes (Sanders and Stappers, 2008), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most.
+Particularly noteworthy was the fact that the workshops helped participants expand their understanding of what social platforms could be, empowering them to feel more hopeful and articulate desires they might have previously dismissed as impossible within conventional social media paradigms.
+This demonstrates how the design method chosen in our study offered a generative path toward hope in domains where narratives of doom are prevalent.
+Our analysis contributes insights about the types of remote social interactions that youth find most meaningful and fulfilling, examining how these desired interactions align with or diverge from current platform affordances, as we detail in Section 5.2.
+We also identify core aspects of design that can undermine relational quality from which we develop the “spatial integrity” framework introduced in Section 5.3.
+Taken together, we hope this work equips designers with proactive design goals for envisioning new futures of social technologies and provides an initial framework for evaluating design integrity in relationship-supporting aspects.
 5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
@@
 (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
-5.3. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
-There are societal concerns about social media’s impact on youth well-being (Oth, 2023).
-Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023).
-This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
-However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality.
-Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023).
-At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024).
-This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connection that youth actually seek.
-Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
-Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be.
-Youth have come to equate social media with certain limiting characteristics: endless content streams, superficial interactions, and 2D interfaces focused on broadcasting posts.
-Interestingly, while they report meaningful social connections and positive experiences in other online social spaces like Minecraft or Discord, they view mainstream social media as inherently tied to emotionally disconnected experiences and superficial interactions.
-This cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
-To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting.
-The Hogwarts-inspired context provided fresh metaphors for exploring social spaces, privacy, and connection.
-Unlike traditional design workshops where youth typically propose incremental improvements to existing platforms, these sessions enabled participants to imagine entirely new possibilities for meaningful connection.
-More importantly, the workshops helped participants expand their understanding of what social platforms could be, empowering them to articulate desires they might have previously dismissed as impossible within conventional social media paradigms.
-This study demonstrates how design methods can offer a generative path toward hope in domains where narratives of doom are prevalent.
 5.4. Limitation and Future Work
 Our study has several key limitations to consider.
-First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently generalize to youth who are not familiar with such environments.
+First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently transfer to youth who are not familiar with such environments.
 Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations.
@@
 6. Conclusion
-This study explored how 3D gamer youth reimagine social media when given the freedom to think beyond the limits of today’s platforms.
-Through the fictional world of Hogwarts, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption that defines mainstream social media.
-Their visions called for platforms that feel more like 3D games than archetypal ones like Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to the complexity of real relationships and incorporate elements of playfulness and whimsy.
-Building on these insights, we introduced the concept of spatial integrity, a set of foundational design principles that make such visions possible.
-While spatial integrity alone does not guarantee meaningful connection, it provides the structural conditions that support the kinds of relational, embodied, and playful interactions participants described.
-The study underscores that youth are not necessarily seeking less time on social media, but rather more fulfilling and emotionally present interactions within these digital spaces.
-Their visions point towards a future where social media feels more like a dynamic, shared social world like those found in 3D gaming environments.
-While our findings are rooted in the experiences of youth already familiar with 3D gaming, they offer a hopeful design direction for spaces that support meaningful social connections that youth desire.
+This study explored how youth with 3D gaming experience reimagine social media when freed from current platform constraints.
+Through Hogwarts’ fictional world, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption defining mainstream social media.
+Their visions called for platforms that feel more like 3D games than Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to real relationships’ complexity while incorporating playfulness and whimsy.
+Building on these insights, we introduced spatial integrity, foundational design principles enabling such visions.
+While spatial integrity alone cannot guarantee meaningful connection, it provides structural conditions supporting the relational, embodied, and playful interactions participants described.
+The study underscores that youth seek not less time on social media, but more fulfilling and emotionally present interactions within digital spaces.
+Their visions point toward a future where social media feels like the dynamic, shared social worlds found in 3D gaming environments.
+While our findings stem from youth familiar with 3D gaming, they offer a hopeful design direction for spaces supporting the meaningful social connections youth desire.
 Appendix A FI Workshop Protocol

### v4

@@
-Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry
+Social Media Should Feel Like Minecraft, Not Instagram: Youth Visions for Meaningful Social Connections through Fictional Inquiry
 Abstract.
-We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments.
-Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces.
-Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics.
-We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth.
-Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction.
-Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
+We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision ideal remote social connection.
+Using Fictional Inquiry (FI) within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord).
+Participants envisioned an immersive platform prioritizing meaningful social connection through presence and immersion, natural self-expression, intuitive spatial navigation leveraging physical-world norms, and playful, low-stakes opportunities for friendship development.
+We synthesize these visions into six themes articulating relational needs that current platforms systematically marginalize.
+The FI method proved effective in generating innovative ideas while empowering youth by fostering hope and agency over social media’s future.
+Our findings challenge “doom” narratives by reframing social media’s harms as outcomes of specific design choices, demonstrating how design research can reopen space for imagining more supportive forms of mediated connection.
 social media; fictional inquiry; youth; design
@@
 1. Introduction
-Meaningful social connections are critical for young people’s development and well-being (Hawkley and Cacioppo, 2010), and social media has great potential to support these.
-At the individual level, richer and healthier interactions support youth well-being and identity development (Kim et al., 2024c).
-At the societal level, platforms act as critical infrastructures for sustaining friendships and community ties, making their design choices consequential for social capital (Phua et al., 2017a; Ellison et al., 2007).
-In fact, many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n.
-d.]).
-However, social media platforms increasingly shift away from facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025).
-While designers initially created these platforms around sharing personal updates with friends, they now emphasize entertainment, news, and engaging larger, less intimate audiences.
-Meta’s data illustrates this shift: between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025).
-Mark Zuckerberg described this change as aligned with the company’s focus on “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
-This shift raises concerns about meaningful online relationships, particularly for young people who have grown up with social media as central to their social lives.
-Most young people engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023).
-Despite high engagement, many report dissatisfaction with their social interactions, not because they want reduced usage, but because they expect platforms to better support their connections (Kim et al., 2024c).
-Current designs prioritize passive content consumption over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
-Given social connections’ importance for young people and social media’s vital role in facilitating them, we must explore new design possibilities.
-Young people frequently express dissatisfaction with current experiences, revealing a significant gap between their social needs and existing platform interactions.
-Mainstream platforms employ standardized design patterns that constrain creative possibilities and foster pessimism among users about social media’s potential.
-This creates a compelling need for innovative designs that complement existing platforms by emphasizing meaningful interpersonal connections.
-Therefore, we ask the following research question: What types of social interactions might current social media designs be overlooking or inadequately supporting?
-To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) with 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n.
-d.]) and Roblox (Roblox Corporation, [n.
-d.]).
-Participants collaboratively imagined a social media platform designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n.
-d.]).
-This imaginative context allowed participants to step away from existing platform constraints and envision more varied and meaningful social interactions.
-Since researchers recognize 3D gaming environments as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated participants could leverage gaming insights to propose novel social media interactions.
-Specifically, our research addresses: What types of interactions do young people familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
-Our findings reveal that young people with 3D gaming experience envision ideal social media platforms as immersive, navigable environments that designers should create to feel like neighborhoods, community centers, and shared social spaces.
-Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming.
-These visions highlight young people’s desire for social media that feels more relational, meaningful, and personally rewarding.
-We contribute empirical findings on young people’s perceptions of current social media platforms’ shortcomings and their desired experiences.
-We propose the concept of spatial integrity for social interactions in online spaces, encompassing four dimensions of spatial affordances that facilitate meaningful social connections: spatial presence (understanding the space and identifying who is present), spatial composition (the space’s meaning and significance), spatial configuration (the space’s organizational structure), and spatial depth (the extent of perceivable information within the space).
-We demonstrate that the FI method in social media design contexts prompts creative idea generation, enables participants to transcend current design patterns, and empowers them to envision social media’s positive potential, leaving them more hopeful about future possibilities.
+Meaningful social connections are critical for young people’s development and well-being (Hawkley-2010-LonelinessMattersMechanisms-v).
+Early social computing researchers and users celebrated these technologies for their potential to build relationships that transcend physical boundaries (ellison2007benefits).
+Over time, however, that promise has narrowed.
+Between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley-2025-InstagramFacebookProof-l).
+Current designs increasingly prioritize passive consumption over active interaction with close friends (Landesman-2024-IInstagram-j), and are associated with fatigue, mistrust, and perceived harm.
+Platforms once imagined as infrastructures for personal connection now make even small acts of self-disclosure—the basic currency of relationship development—feel high-stakes (Kim2025Trust-EnabledRegulationz) and anxiety-inducing (jaewon-privacy).
+For youth who have grown up with these technologies as central to their social lives, this shift carries consequences beyond user experience.
+While young people report frustration, anxiety, and toxicity, they also describe moments of genuine joy and meaningful connection—and participation remains a practical necessity for staying connected with peers.
+Yet dominant policy, educational, and design responses frame youth engagement in terms of prevention and protection, emphasizing restriction over support for experimentation that might foster agency and resilience (Kim2025DesignChallengese; Kim2024EnvisioningPreventingn).
+Youth consistently report that measures such as outright bans leave them feeling unheard (Kim-2024-AustraliaBarred-y; TheLearningNetwork-2025-WhatTeensMedia-z; Weinstein2022-rh).
+Their relationships with social media are ambivalent rather than uniformly negative, but this complexity is rarely reflected in prevailing approaches to governance or design.
+This disconnect is reinforced by the broader convergence of social media platforms themselves (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m).
+Despite surface-level differences, mainstream platforms increasingly share a common paradigm organized around feeds, mobile interfaces, and vertical profiles, crowding out alternative models of mediated connection.
+As a result, much design research now operates in a reactive mode.
+The guiding question has shifted from what meaningful online connection might look like to how people can be protected from existing platforms.
+This framing treats the dominant paradigm as fixed, focusing on mitigation rather than reimagination.
+Although generative work exists within HCI and related fields, it is often overshadowed by protectionist narratives that position youth as vulnerable users in need of safeguards, prioritizing safety over empowerment and incremental repair over structural alternatives (Kim2025DesignChallengese; Kim2024EnvisioningPreventingn).
+These constraints shape not only policy and design discourse but also participatory research itself.
+When young people are asked to redesign social media, the prompt arrives burdened with existing controversies around addiction, mental health, and toxicity.
+Participants often respond defensively, reproducing adult-driven narratives through social desirability bias even when their lived experiences are more nuanced (Davis2019EverythingsRelationshipsg).
+Without language or precedents beyond existing platforms, participants tend to remix familiar features—a cleaner feed, better moderation, fewer metrics—rather than imagining fundamentally different forms of social interaction.
+The term “social media” itself anchors imagination to platforms like Instagram or TikTok, constraining what can be envisioned before design even begins.
+As a result, participatory design risks generating refinements when its prompts and methods keep participants inside the same conceptual frame.
+To move beyond these constraints, we ask:
+  • RQ: What forms of social interaction do youth envision when freed from current platform constraints?
+Rather than beginning from existing systems or problems, we focus on the relationships, interactions, and forms of presence young people desire, and on the design considerations that emerge from these visions.
+To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (IversenDindler-2007-FictionalInquiry--designSpace-m), a speculative approach that invites participants to imagine within a fictional scenario removed from real-world constraints.
+Twenty-three participants aged 15–24 imagined how students at Hogwarts—the fictional school from Harry Potter(harrypotterfanclub)—might connect with friends beyond their campus using any magical powers they could envision.
+This framing served three purposes.
+First, it functioned as a defamiliarization strategy, breaking the conceptual lock of existing platforms by removing both technical limitations and loaded terminology, starting from desire rather than constraint.
+Second, it centered meaningful social connection as the explicit design goal, reflecting youth developmental needs and interests documented in prior research (Davis2012-bq).
+Third, it created a space of play in which youth are natural experts, enabling them to express values, social logics, and relational priorities difficult to articulate within dominant narratives about social media.
+Participants were not asked to redesign social media; they were asked how wizards might stay close across distance.
+Our findings reveal six themes characterizing how youth envision meaningful remote social interaction: intuitive social navigation modeled on real-world movement and presence; shared experiences through collaborative activities and casual co-presence; communal environments that prioritize close relationships and low-pressure interaction; flexible self-presentation through avatars and contextual identity management; intentional engagement that replaces passive scrolling with purposeful navigation; and playful, low-stakes social mechanics.
+Within these themes, participants described concrete interactions—sitting together in a virtual study room with Lo-Fi music, leaving a small gift at a friend’s door without initiating a full conversation, or using secret knocks to grant trusted friends access to private spaces.
+Although participants developed their visions independently, their accounts of how these imagined systems would feel converged strongly, emphasizing connection, engagement, fulfillment, and agency.
+These patterns suggest shared relational needs not well supported by current platforms.
+We also observed that the FI process itself shifted participants from reactive problem-fixing toward aspirational imagining, with many reporting increased hope and excitement about what mediated connection could become.
+This paper makes three contributions.
+First, we offer a methodological account of FI as a research-through-design approach for reimagining meaningful remote social connection.
+By suspending constraints and removing the term “social media” itself, the fictional framing enabled youth—a population often positioned as subjects to protect rather than agents of imagination—to reason from felt experience rather than inherited platform templates.
+Second, we present an empirical account of the interaction concepts, affordances, and social mechanisms young people proposed when freed from feed-based paradigms.
+Third, we synthesize these concepts into six themes that articulate underlying relational needs and design directions, highlighting what modern social media systematically marginalizes.
+Together, these contributions challenge “doom” narratives around social media, reframing them as outcomes of specific design choices and demonstrating how design research can reopen space for imagining more supportive forms of mediated connection.
 2. Related Work
-2.1. Meaningful Social Connection in Online Environments
-2.1.1. Defining Meaningful Social Interaction
+2.1. Beyond Moral Panic: From Panic Narratives to Design Opportunity
+Widespread societal concerns exist about social media’s impact on youth well-being (OtherOther-2023-SocialMediaAdvisory-u).
+Rather than addressing these constructively, public discourse has framed social media as a societal evil (Wike-2022-2ViewsSociety-y), particularly in the U.S., where it is blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt-2023-SocialMediaEvidence-c).
+This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
+However, treating social media as inherently harmful, or youth as best served by removal, collapses a more complex reality.
+Despite public concern about social media’s role in youth mental health, causal evidence remains inconclusive, and youth report a mix of positive and negative experiences (Vogels-2023-TeensSocialSurveys-n).
+Fear-based policies and alarmist narratives can themselves cause harm, amplifying anxiety, mistrust, and helplessness among youth (kim2024privacysocialnormsystematically; lee2024social).
+Recognizing nuance does not deny that serious harms occur, including high-risk vulnerabilities (e.g., sexual exploitation, grooming, and synthetic media threats), nor ignore structural forces like platform power asymmetries and algorithmic manipulation.
+Instead, it motivates a complementary design question: what kinds of online social environments could better support youth relationships while accounting for both everyday experiences and real risks?
+2.2. Meaningful Social Connection Online, and Why Mainstream Platforms Miss It
+What counts as meaningful connection
 Litt et al.
-(Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself,” involving positive emotions, information exchange, or concrete benefits.
-The study identified three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services).
-Non-meaningful interactions lacked any such impact.
-Strong ties (friends and family), community ties, and shared activities were associated with meaningfulness, but acted as facilitators enabling impact rather than primary sources of meaning.
-Other attributes linked to meaningfulness included advance planning, memorialization (photos or videos), and synchronicity.
-The communication medium itself—online or offline—was less important than these other characteristics.
-We define meaningful social interaction as any interaction that participants would actively seek in an ideal scenario, reflecting their subjective meaningfulness.
-We expected these themes to emerge in participants’ descriptions of their experiences.
-2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
-Though social media platforms are primary connection tools for young people (Boyd, 2014a), many users report feeling lonely or disconnected despite widespread use (Twenge et al., 2019; Primack et al., 2017), highlighting a gap between intended purpose and actual experience.
-While social connection remains a primary motivation for social media use, mainstream platforms like Instagram (All, [n.
-d.]), YouTube (you, [n.
-d.]), and Discord (discord page, [n.
-d.]) are designed for diverse user motivations including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022).
-From a uses and gratifications perspective (Katz et al., 1973), users pursue varied goals from building social capital (Phua et al., 2017b) to media consumption (Menon, 2022).
-Users seeking social connection may encounter experiences prioritizing other engagement forms, leaving social needs unmet.
-This gap underscores the need for a clearer understanding of meaningful social connections in online environments.
-While social media presents risks like social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b) alongside benefits like marginalized group support (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and identity development opportunities (Davis, 2012; Lee et al., 2022), social connection remains central for many users.
-For young people increasingly relying on digital platforms, we must move beyond treating connection as one competing goal toward a focused understanding of how platforms can meaningfully support social relationships.
-2.2. Social Technology Design Explorations in HCI
-2.2.1. Designing for Meaningful Social Connections
-HCI research has extensively explored designs fostering meaningful social interactions online at both platform and feature levels.
-Platforms like BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) introduced unique affordances emphasizing spontaneity, ephemerality, and creative self-expression.
-This research identified ways to build relationships through features prioritizing authenticity and playful interaction.
-However, users report increasingly toxic interactions as platforms gain popularity, with design choices encouraging network expansion and compulsive behaviors like tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), detracting from meaningful interaction goals.
-Feature-level efforts include designs encouraging effortful communication (Liu et al., 2021; Zhang et al., 2022) and genuine connections (Stepanova et al., 2022).
-However, social media experiences are holistic, involving a complex interplay of multiple design aspects and user behaviors.
-Individual features alone cannot fundamentally alter overarching social media narratives.
-Beyond archetypal social media, recent work investigates spatial social technologies emphasizing immersive, embodied interactions.
-Social VR platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide virtual spaces and embodiment for co-presence and shared experiences, facilitating more personal and tangible interactions than 2D environments.
-Platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend gaming and social networking, enabling virtual environment navigation, avatar customization, and dynamic group interactions.
-However, these platforms are not generally recognized as conventional social media.
-What youth envision as ideal social media at the landscape and feature levels remains unclear, especially considering emerging technologies like spatial computing platforms.
-Understanding these perspectives is crucial for reimagining social media’s future to meet users’ evolving needs.
-2.2.2. Video Games as a Source of Meaningful Social Connection
+(Litt-2020-WhatMeaningfulSurvey-h) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself.” They identify three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services).
+Strong ties, community ties, and shared activities facilitate meaningfulness, while attributes like advance planning, memorialization, and synchronicity often support impact.
+Building on this work, we treat meaningful social interaction as any interaction participants would actively seek in an ideal scenario, reflecting their subjective sense of meaningfulness.
+Connection-seeking on platforms optimized for other gratifications
+Though social media platforms are primary connection tools for young people (BoydBoyd-2014-ComplicatedSocialTeens-j), many users report feeling lonely or disconnected despite widespread use (twenge2019less; primack2017social), highlighting a gap between social media’s promise and lived experience.
+Mainstream platforms like Instagram (instagram), YouTube (youtube), and Discord (discord) serve diverse user motivations including entertainment, information seeking, and self-promotion (whiting2013people; Menon-2022-UsesGratificationsInstagram-x).
+From a uses and gratifications perspective (GurevitchKatz-1973-UsesGratificationsResearch-y), users pursue varied goals from building social capital (phua_uses_2017) to media consumption (Menon-2022-UsesGratificationsInstagram-x).
+Users who primarily seek meaningful connection may thus encounter experiences that prioritize other forms of engagement, leaving social needs unmet.
+HCI and CSCW work has explored designs supporting more meaningful interaction, including efforts encouraging effortful communication (LiuFannie2021SOUt; LiuZhang-2022-AuggieEncouragingExperiences-n) and reflective or value-oriented social interaction (Stepanova2022-vh).
+More broadly, recent work articulates design spaces and system-level alternatives beyond dominant templates (e.g., form factors, interaction conventions, and social architectures) (zhang2024form).
+Yet if connection is so central, why do mainstream platforms consistently fail to prioritize it?
+Part of the answer lies in structural forces constraining both platform incentives and design imagination.
+2.3. Recognizing and Overcoming Social Media Design Fixation
+Design fixation (jansson1991design) describes the tendency to remain constrained by conventional thinking or existing solutions, even when divergent exploration is needed.
+Social media design faces similar fixation (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m).
+Ad-based revenue models incentivize increased time on apps (Other-Other-PlatformCapitalism-f), shaping platforms toward engagement metrics and features that prolong attention.
+Platforms deploy compulsive strategies like friend recommendations (HinikerKim-2024-SharingDesign-x) and engagement metrics (e.g., Snap scores (rozgonjuk2021comparing)), which erode trust and foster distrustful user communities (Kim2025Trust-EnabledRegulationz).
+In parallel, design efforts are often dominated by reactive harm mitigation rather than proactive exploration of new approaches (Kim2024EnvisioningPreventingn).
+Together, these forces help explain why users experience mainstream platforms as converging toward similar, feed-driven patterns, even when those patterns do not align with meaningful connection goals.
+These constraints motivate methods that deliberately create distance from current systems, enabling participants to articulate what they would want social media to optimize for without starting from existing platform assumptions.
+2.4. Alternative Paradigms and Methods for Imagining Beyond Platform Templates
+Sociality in games and multi-space online communities
 Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth.
-Large-scale surveys show games serve as major venues for friendship development and maintenance, especially for boys (Lenhart, 2015c).
-Teens play with known friends and make new ones through online gameplay (Lenhart, 2015a), with over half reporting meeting at least one friend online (Lenhart, 2015a, b).
-These interactions extend beyond gameplay, supporting ongoing relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
-Research highlights gaming’s social and emotional benefits, including cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014).
-Certain designs promote prosocial outcomes by encouraging players to help others in-game and beyond (Gentile et al., 2009).
-Games offer deeper, eudaimonic experiences involving meaning-making, personal growth, and connection (Daneels et al., 2023).
-These findings suggest video games can foster meaningful social connections among youth and highlight the potential for game-like features in social media.
-2.3. Recognizing and Overcoming Social Media Design Fixation
-Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented design challenge.
-Though generating divergent ideas is critical for innovation, we often impose unnecessary restrictions.
-This self-limiting process confines design to localized problems rather than exploring broader, optimal solutions.
-Social media design faces similar fixation (Sundaram, [n.
-d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020).
-Ad-based revenue models incentivize increased time on apps (Oth, [n.
-d.]).
-Platforms use compulsive strategies like friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)), creating distrustful user communities (Kim et al., 2025).
-Rather than proactively exploring new paradigms of designs and interactions, much recent work on social media design has disproportionately focused on reactively addressing existing harm (Kim et al., 2024b).
-The domination of mainstream platforms along with this reactive approach to design leaves users dissatisfied and reinforces perceptions that all social media experiences are fundamentally the same (Sundaram, [n.
-d.]; Pardes, 2020).
-Analogical transfer offers one powerful method for overcoming design fixation (Novick, 1988; Buchanan, 1992).
-Therefore, in this paper, we attempt to break out of the design and cognitive fixation that assumes the future of social media is bleak by freeing participants from real-world constraints and beliefs and leveraging elements from fictional settings to serve as sources for analogical transfer.
+Large-scale surveys show games serve as major venues for friendship development and maintenance (Lenhart-2015-Chapter3Boys-v), with teens playing with known friends and making new ones through online gameplay (Lenhart-2015-Chapter1Friendships-i).
+CSCW research shows that gaming communities often extend across multiple spaces, requiring coordination and group management across platforms and contexts (toombs2022we).
+These findings suggest gaming-influenced sociality can illuminate how people sustain relationships through shared activities, low-stakes co-presence, and community infrastructure, while raising questions about what translates and who may be excluded by more spatial or higher-cognitive-load interaction styles.
+Spatial social technologies and place-based interaction
+Beyond archetypal social media, HCI and CSCW have examined spatial social technologies emphasizing immersive and embodied interaction.
+Social VR and virtual worlds support co-presence, shared experiences, and identity play through avatars and navigable spaces (Zamanifard-2023-SurpriseBirthdayDistance-n; Freeman-2024-MyAudiences-u; Freeman-2020-MyBodyReality-l; Zamanifard-2019-TogethernessCraveRelationships-s; deighan2023social; rzeszewski2024social).
+Platforms like Gather.town (duarte2023experience; tu2022meetings) and VRChat provide place-like environments where spatial layout, navigation, and ambient co-presence shape interaction rhythms.
+A longer history of virtual worlds (e.g., Second Life) highlights that spatiality can enable rich social experience while introducing familiar challenges around governance, safety, and uneven participation.
+We draw on these literatures as precedents for treating online social spaces as places, where constructs from place theory, environmental psychology, and presence research (e.g., propinquity, wayfinding, place attachment, social presence, and co-presence) can help interpret how spatial organization relates to social connection.
+Youth-centered design and fiction-based methods for escaping fixation
+Youth-focused HCI has long argued for engaging young people as stakeholders and co-designers rather than solely as subjects of protection.
+HCI has also developed fiction-based and speculative methods that create distance from present-day systems, making it easier to reason about values, tradeoffs, and alternative social arrangements.
+Related approaches include design fiction, speculative design, and pastiche scenarios, which use crafted narratives and artifacts to surface new possibilities and critique existing assumptions (blythe2006pastiche; dunneraby2013speculative).
+Prior HCI work has drawn explicitly on fictional worlds, including Harry Potter, to design for social values and family coordination (brown2007locating).
+Building on this methods lineage, we use FI to structure analogical transfer through a shared fictional narrative; unlike design fiction or pastiche scenarios that often center researcher-authored speculative artifacts, FI treats a well-known fictional world as a common design constraint and supports participant-led world-building of social architectures and norms.
+This approach is especially relevant for domains like social media, where incentive structures and design fixation can make it difficult to surface what youth might find meaningful outside current assumptions.
 3. Method
-We conducted co-design interviews with 23 participants aged 15-24 to examine how we might envision ideal social media platforms free from real-world constraints and design fixation on archetypal platforms.
-All participants had 3D gaming experience and Harry Potter (Rowling, 2015; har, [n.
-d.]) familiarity, a popular young adult fantasy series where some possess magical powers.
-We used the FI method to scaffold participants’ design idea generation.
+We conducted co-design interviews with 23 participants aged 15–24 to examine how youth envision ideal social media platforms when freed from real-world constraints and fixation on archetypal platforms.
+All participants had prior experience with 3D games and were familiar with Harry Potter (rowling2015harry; harrypotterfanclub), a popular fantasy series featuring characters with magical powers.
+We used the FI method to scaffold participants’ design ideation.
 All procedures were approved by the university’s Institutional Review Board (IRB).
-3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
-FI is an exploratory co-design method enabling participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007).
-FI creates environments where people imagine beyond present limitations through carefully crafted scenarios incorporating immersive elements like physical artifacts or role-playing activities.
-Participants take generative roles envisioning futuristic or fictional settings.
-The method effectively overcomes design fixation by encouraging analogies from fictional scenarios and free imagination outside real-world constraints.
-Unlike other ideation techniques, FI emphasizes developing futuristic design materials through collaborative designer-user dialogue (Dindler and Iversen, 2007).
-Dindler and Iversen (Dindler and Iversen, 2007) outline three key implementation steps: defining the inquiry’s purpose (staging design situations, exploring future ideas, or driving organizational change); developing an appropriate narrative that participants find comfortable yet distinct enough from current practices to encourage new thinking; and creating compelling plots introducing tension or conflict to motivate workshop activities.
-Unlike traditional role-playing, FI encourages participants to maintain their identities and expertise while operating in fictional contexts, exploring new possibilities through personal motivations and capabilities while circumventing societal constraints.
-We view video games as rich analogy sources for social media design, offering unique affordances and interaction patterns inspiring new approaches to meaningful social connections.
-By examining how games facilitate social interactions, we can identify adaptable design patterns for social media contexts, breaking free from fixation to explore novel alternatives.
+3.1. FI as a Tool for Imaginative Design Exploration
+FI is an exploratory co-design method that enables participants to move beyond real-world constraints through immersion in semi-fictional contexts (IversenDindler-2007-FictionalInquiry--designSpace-m).
+Through carefully crafted scenarios incorporating immersive elements such as physical artifacts or role-playing activities, participants take on generative roles, envisioning futuristic or fictional settings.
+FI helps teams bypass existing sociocultural constraints by working within a shared, partially fictional narrative, which can make it easier to ask questions that might otherwise feel awkward and to examine specific issues in practice (IversenDindler-2007-FictionalInquiry--designSpace-m).
+It also supports future-oriented thinking by reducing the perceived “constraints of reality” and shifting attention from fixing present-day problems toward imagining desirable alternatives (LeeHiniker-2017-Co-DesigningPreschoolersComicboarding-e).
+More broadly, fictional settings function as a rigorous inquiry tool: they can surface new insights by contrasting fiction with reality, stress-test ideas through counterfactual reasoning, and reveal how imagined worlds shape real-world thinking and action (Schoeneborn2022FictionalInquiryg).
+Dindler and Iversen (IversenDindler-2007-FictionalInquiry--designSpace-m) outline three key implementation steps: defining the inquiry’s purpose (e.g., staging design situations, exploring future ideas, or driving organizational change); developing a narrative that participants find approachable yet sufficiently distinct from current practices to encourage new thinking; and creating compelling plots that introduce tension or conflict to motivate design activity.
+Unlike traditional role-playing, FI encourages participants to retain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities through personal motivations and capabilities while sidestepping societal constraints.
+3.1.1. Addressing Design Fixation Through FI
+We hypothesized that FI could support idea generation in social media design, a design space often perceived as stagnant or doomed.
+Perceptions of inevitability are reinforced by the convergence of mainstream platforms around similar structures and interaction conventions (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m), as well as by economic incentives prioritizing attention capture and sustained time on platform (Other-Other-PlatformCapitalism-f).
+Recent scholarship further suggests that design efforts have become dominated by reactive harm mitigation rather than proactive exploration of alternative paradigms (Kim2024EnvisioningPreventingn).
+Together, these conditions constrain both imagination and inquiry, making it difficult to surface what people might find meaningful outside current platform assumptions.
+By freeing participants from real-world constraints and encouraging them to draw on experiences from other media such as games, movies, VR, stories, and magical or fictional worlds, FI allowed participants to engage with concepts they find meaningful without concern for social desirability, ethical scrutiny, or technical feasibility.
+Given the strong design and cognitive fixation characterizing social media design, this form of analogical transfer may enable conceptual breakthroughs (jansson1991design; novick1988analogical).
+Reframing social media design within a fictional setting also functions as a problem-reframing strategy, a core aspect of design thinking that can help designers move beyond entrenched assumptions (Buchanan-1992-WickedProblemsThinking-x).
+Accordingly, we hypothesized that FI would be effective both in uncovering youths’ underlying needs around remote social connection, free from dominant narratives or fear of judgment, and in enabling design breakthroughs through problem reframing and analogical transfer.
 3.2. Material: The Narrative Framework
-Figure 1. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
-[…]
-This is a visual chart designed to organize relationships and connections in the Harry Potter universe.
-There are three columns of labeled sections.
-In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue).
-Each house box has its crest.
-To the left, there are two smaller boxes.
-The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname.
-Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name.
-To the right, there are two more boxes.
-The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family.
-Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid.
-At the bottom, there are three large sections.
-The leftmost section is labeled “Muggles” with a blank space.
-The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo.
-The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes.
-Each section contains an empty space for names or details to be added.
 We selected the Harry Potter world as our FI narrative framework.
@@
 Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement.
-By 2023, 98% of US youth aged 18-27 identified as casual or avid fans, with half describing themselves as avid fans (Share of Harry Potter fans U.S.
-2023, [n.
-d.]).
-The book series sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K.
-Rowling’s Harry Potter and the Sorcerer’s Stone, [n.
-d.]) and was translated into over 80 languages by 2020 (The Rowling Library, 2020).
+By 2023, 98% of US youth aged 18–27 identified as casual or avid fans, with half describing themselves as avid fans (harrypotterfan).
+The book series sold over 600 million copies (harrypotterscholastic) and was translated into over 80 languages by 2020 (TheRowlingLibrary-2020-BigTranslationsPotter-z).
 Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints.
-Young protagonists and coming-of-age narrative align with our focus on youth identity and relationship building.
+Young protagonists and the coming-of-age narrative align with our focus on youth identity and relationship building.
 Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts.
@@
 Interviews averaged 90 minutes.
-We began by asking about 3D gaming experiences and Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
-We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (mug, [n.
-d.])) connecting across Hogwarts and Muggle worlds.
-We shared a Miro (mir, [n.
-d.]) board (Figure 1) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles).
+We began by clarifying what 3D gaming experience means, asking participants what they like or dislike about 3D environments compared to 2D ones, and discussing Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
+We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (muggle)) connecting across Hogwarts and Muggle worlds.
+We shared a Miro (miro) board (Figure 2) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles).
 This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion.
-We then asked about social media experiences, focusing on meaningful versus meaningless aspects and ideal platform visions.
-During the design phase, participants described their ideal magical social media capabilities and what life aspects they would share.
-We maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines.
-On the other hand, we encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform.
+During the design phase, we asked about social media experiences, focusing on meaningful versus meaningless aspects across existing platforms and what they envision in an ideal platform.
+While developing ideal magical social media capabilities and what life aspects they would share, we maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines.
+We also encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform.
 Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take?
@@
 A secret room?
-A creature?)” We concluded by asking participants to compare their envisioned Hogwarts social media with Muggle social media.
-The semi-structured protocol is in Appendix A.
+A creature?)” We concluded by asking participants to reflect on how using their envisioned platform would feel and to compare it with their experiences on current social media.
+The semi-structured protocol is in Appendix B.
 3.4. Participants and Recruitment
-We used purposive and convenience sampling for recruitment.
+We used purposive and convenience sampling.
 The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
-Participants needed to be 15-24 years old per UN (United Nations, [n.
-d.]) and WHO (who, [n.
-d.]) youth definitions and consent to video/audio recording during 90-minute interviews.
+Participants needed to be 15–24 years old per UN (UnitedNationsUnitedNations-Other-Youth|Nations-t) and WHO (who-youth) youth definitions and consent to video/audio recording during 90-minute interviews.
 Video recording followed FI guidelines, enabling gesture analysis if needed.
 We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
-Participants needed sufficient Harry Potter familiarity (agreeing with “I remember some details about the plot, characters, and magical aspects”) and prior 3D gaming experience (e.g., Minecraft, Roblox).
-We did not require extensive knowledge of Harry Potter because we felt that would be an unnecessary bias that is not particularly helpful in envisioning new social media design.
-While this made it somewhat difficult for a majority of participants to remember specific artifacts from the Harry Potter world, they knew enough to recognize the complete and extensive freeing extent of the magical powers in that setting.
-Though we imposed no social media experience constraints, all participants had some social media engagement.
+Because our FI sessions used a Harry Potter-themed scenario to imagine digital spaces unconstrained by current design conventions, we required participants to have at least moderate familiarity with the franchise (agreeing with “I remember some details about the plot, characters, and magical aspects”).
+We did not require extensive knowledge, as deep familiarity with a specific fictional world seemed unnecessary for envisioning new possibilities for social connection and might introduce selection bias.
+We also required at least 20 hours of prior 3D gaming experience (e.g., Minecraft, Roblox).
+This criterion served as a baseline threshold for participants to draw on experiences with digital interfaces beyond 2D mobile screens when imagining alternative designs.
+While other modalities such as virtual reality could support such imagination, VR experience remains relatively uncommon among youth—only 24% of teens have played games on a VR headset (Gottfried2024TeensGamesx).
+In contrast, video gaming is nearly universal: 85% of U.S.
+teens play video games, with the majority playing on consoles or desktop computers that render 3D environments (Gottfried2024TeensGamesx).
+All but one participant met the gaming experience criterion.
+Although we imposed no constraints on social media experience, all participants reported some social media engagement.
 Table 1.
-Demographics Summary.
-Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”).
-Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace).
-Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
-PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
-P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
-P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
-P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
-P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
-P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
-P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
-P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
-P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
-P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
-P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
-P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
-P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
-P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
-P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
-P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
-P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
-P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
-P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
-P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
-P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
-P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
-P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
-P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
+Participant demographics and envisioned social media features.
+Abbreviations for Harry Potter familiarity: Expert (remembers every detail), Proficient (remembers most details), Familiar (remembers some details).
+Abbreviations for platforms: IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace).
+Rows prefixed with “>” denote envisioned social media features.
+PID Gender Age Race/Ethnicity HP Platforms
+P01 Woman 19 Asian Proficient IG, DC, BR, RD
+> Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates and casual interactions; Public areas for open interaction & interest-based rooms; Area-based access control; Music & mood reflection in rooms.
+P02 Woman 20 Asian Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI
+> Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
+P03 Woman 18 Black Proficient IG, DC, TT, X, RD, TH, NP
+> Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; VR-like navigation; Selective visibility & privacy settings; Shopping district & immersive commerce; Customizable avatars & dynamic appearance.
+P04 Woman 16 Asian Expert IG, DC, SC, TT, BR, X, RD
+> Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
+P05 Woman 18 White Proficient IG, DC, BR, RD, FB, TH
+> Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
+P06 Non-binary 18 White Familiar IG, DC, SC, TT, RD, FB
+> Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers; Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
+P07 Woman 17 Hispanic/White Proficient IG, DC, SC, TT, BR, X, FB
+> Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
+P08 Woman 19 White Proficient IG, DC, SC, RD, FB
+> Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
+P09 Woman 21 Asian/White Proficient IG, DC, TT
+> Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Room of Requirement-inspired group chats.
+P10 Non-binary 17 Hispanic/White Proficient IG, DC, SC, TT, BR, RD, TH
+> Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
+P11 Woman/Non-binary 19 Native American/White Familiar IG, DC, TT, X, RD, FB, TH
+> Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
+P12 Man 19 Asian Familiar IG, DC, SC, TT, X
+> Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; Interest-based connection zones.
+P13 Woman 17 Asian Familiar IG, DC, SC, TT, BR, X, RD, FB, TH
+> Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
+P14 Woman 18 Hispanic Expert IG, DC, SC, TT, BR, X, FB
+> Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Authentic avatars.
+P15 Woman 20 Asian Proficient IG, DC, SC, BR, X, RD
+> 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
+P16 Man 18 Asian Familiar IG, DC, SC, BR, X, RD
+> Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
+P17 Woman/Non-binary 17 Hispanic/White Expert IG, DC, SC, TT, X, RD, TH
+> Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
+P18 Woman 23 Asian Familiar IG, DC, SC, TT, X, FB
+> Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
+P19 GenderFluid/Trans Man 19 White Expert IG, DC, SC, TT, X, RD, FB
+> Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
+P20 Woman 17 Black Proficient IG, DC, SC, TT, BR, X, FB
+> Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
+P21 Woman 17 Asian Expert IG, DC, SC, RD
+> Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking; Immersive social events; Space-based access control.
+P22 Woman 19 White Familiar IG, DC, SC, TT, X, RD
+> Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
+P23 Man 18 Asian/Hispanic Familiar IG, DC, TT
+> Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; “Hall of Fame” memory posts.
 Eighty-four individuals completed the screener survey.
@@
 Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants.
-While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel, 2019), we note that our participant demographics skew toward women.
+While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel2019UsingResearchs), we note that our participant demographics skew toward women.
 Full demographic information appears in Table 1.
-(Lenhart, 2015b, c)
 3.5. Data Analysis
-We analyzed interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides systematic yet adaptable qualitative data interpretation.
+We analyzed interview transcripts using reflexive thematic analysis (BraunClarke-2021-ThematicAnalysisGuide-k), which provides systematic yet adaptable qualitative data interpretation.
 Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
-Since not all participants agreed to have their videos on during interviews, we did not conduct video analysis of gestures that participants sometimes used when describing their ideas.
-Additionally, drawing was not part of our protocol because we have experienced in prior interviews for other studies that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination.
-All data therefore consisted of transcripts from verbal interviews.
+Since not all participants agreed to have their videos on during interviews, we did not conduct gesture analysis.
+Additionally, drawing was not part of our protocol because we have found in prior interviews that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination.
+All data therefore consisted of verbal interview transcripts.
 Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data.
-Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atl, 2024).
+Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atlas).
 After four iterations, we reached code saturation.
@@
 Given J.K.
-Rowling’s controversial transgender statements (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K.
+Rowling’s controversial transgender statements (Other-2020-JKIssues-x; Miller-2024-JKTransantagonism-a), we included a disclaimer in recruitment materials and surveys, reiterating our position during interviews.
+The disclaimer read: We acknowledge that J.K.
 Rowling has made statements regarding transgender individuals that many find harmful and discriminatory.
@@
 Thank you for your understanding and for helping us advance this important research.
-in recruitment materials and surveys, reiterating our position during interviews.
 4. Results
@@
 We refer to this synthesized concept as “Social Media at Hogwarts,” or SMH, throughout this paper.
-4.1. Intuitive Social Navigation
-Youth sought interactions that feel natural, embodied, and socially meaningful, reflecting how people navigate relationships and spaces in everyday life.
-They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy management in intuitive, socially familiar ways.
+In line with the goals of co-design processes (Sanders2008Co-creationDesignh), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most.
+4.1. Theme 1: Intuitive Social Navigation
+Youth sought interactions that feel natural, embodied, and socially meaningful—reflecting how people navigate relationships and spaces in everyday life.
+They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy in intuitive, familiar ways.
 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
-Participants described how embodied, immersive qualities like realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space fostering presence, emotional connection, and meaningful interaction.
-In contrast, current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation, explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
+Embodied, immersive qualities—realistic movement, lifelike avatars, holographic interfaces—could transform SMH into a space fostering presence and emotional connection.
+Current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation: “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
 For others, fulfillment came from embodied physical actions.
-P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing,” explaining “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
-P03 reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
-This sense of presence was felt even in seemingly passive actions. As P01 explained,
+P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing”: “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” Traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
+P03 reflected that simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
+This sense of presence emerged even in seemingly passive actions. As P01 explained,
 “I don’t know, there’s just a feeling about it that makes it different.
@@
 I don’t know if that makes sense but.” (P01)
-For P01, 3D interaction supported more embodied engagement and introduced a felt presence that lingered, even without direct interaction.
+For P01, 3D interaction introduced a felt presence that lingered, even without direct interaction.
 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
-Participants proposed using spatial and visual cues modeled after real-world environments to make managing access and privacy feel more intuitive, socially natural, and contextually appropriate.
+Participants proposed spatial and visual cues modeled after real-world environments to make access and privacy management feel more intuitive and contextually appropriate.
 P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
-P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
+P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” They expected others to recognize and respect spatially cued privacy norms:
 “I don’t think people should be able to request access to areas when you’re not there.… Why would you want to go into my virtual bedroom when I’m not even home?
 If I wouldn’t do it in real life, I probably wouldn’t want to do it online either.” (P22)
-While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive and natural.
-P11 suggested spatial contexts could reduce privacy control efforts, stating “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
-4.2. Shared Experiences
-Participants emphasized that shared experiences, through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
+While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive.
+P11 suggested spatial contexts could reduce privacy control efforts: “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
+4.2. Theme 2: Shared Experiences
+Shared experiences—through active collaboration or simply being together—play a crucial role in forming meaningful relationships and lasting memories.
 4.2.1. Fostering Shared Memories Through Collaborative Activities
-Participants described shared activities as meaningful foundations for building friendships.
+Shared activities served as meaningful foundations for building friendships.
 P06 contrasted casual solo games with more immersive, co-experienced ones.
-While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” they described Roblox as fostering lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
-P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
+While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” Roblox fostered lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
+P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” Working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
 The value of shared experiences extended beyond completing tasks.
@@
 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
-Participants saw low-pressure, ambient social interaction, like co-existing in shared spaces or listening to music together, as natural ways to build connections and shared memories.
-P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual study room with Lo-Fi music.
-These spaces would mirror Discord servers’ 24/7 activity, offering comforting presence even without direct conversation.
-Even “mindless scrolling” became meaningful when shared, as P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
-Music often facilitated these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs,” reflecting on their experience:
+Low-pressure, ambient interaction—co-existing in shared spaces or listening to music together—offered natural ways to build connections.
+P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual room with Lo-Fi music, mirroring Discord servers’ 24/7 activity and offering comforting presence without direct conversation.
+Even “mindless scrolling” became meaningful when shared: “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
+Music often facilitated these interactions. P19 envisioned a “jukebox” where friends could “queue up songs”:
 “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.
@@
 It was just nice.” (P19)
-They added that music could serve as a “conversation starter” while providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music.
-You can just listen to the music.”
-4.3. Communal Ambience
+Music served as both “conversation starter” and passive connection: “you don’t have to connect through talking to someone about music. You can just listen to the music.”
+4.3. Theme 3: Communal Ambience
 SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison.
-P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features fostering organic, intentional, personalized interactions and closer friendships.
-They described several design features–neighborhood structures, shared gathering spaces, and stronger boundary controls–that could cultivate this communal, low-pressure environment.
+P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, SMH would foster organic, personalized interactions and closer friendships through neighborhood structures, shared gathering spaces, and stronger boundary controls.
 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
-Participants wanted SMH’s 3D world to mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster social interaction and relationship-building.
-P22 described in detail:
+SMH’s 3D world would mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster interaction. P22 described this in detail:
 “So, in a virtual world, you might have your house, and then you could go to a community board or town square with postings for different groups… From there, you could travel to those group spaces.
@@
 I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
-Youth envisioned SMH reflecting real-world relationship dynamics, using physical distances within the landscape to represent social tie strength.
-Some participants envisioned SMH resembling a college campus or neighborhood, where physical closeness naturally fosters interaction.
-One participant elaborated:
+Physical distances within the landscape would represent social tie strength. One participant elaborated:
 “And in this magical world, I imagine the people around you in your neighborhood would be from the same house as you, kind of like how people live together in dorms on a college campus.
@@
 P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach.
-This approach of aligning physical landscapes with emotional closeness would streamline interactions and focus on meaningful relationships.
-P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
+This alignment of physical landscape with emotional closeness would streamline interactions.
+As P07 noted, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
-Participants imagined SMH neighborhoods including shared third places where people could casually gather, lowering social barriers and creating connection opportunities.
-P05 described these as “neutral spaces” not tied to specific activities or relationships, where groups could gather for various purposes.
-They imagined “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
-Participants described shared neighborhood spaces like common rooms, courtyards, or clubhouses where residents could casually interact outside personal spaces.
+SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating connection opportunities.
+P05 described these as “neutral spaces” not tied to specific activities or relationships, with “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
+Shared neighborhood spaces—common rooms, courtyards, clubhouses—would allow casual interaction outside personal spaces.
 P01 described gathering in a “courtyard or game room” at the neighborhood center.
-Drawing on Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities.
-P09 imagined a “virtual dining room” allowing casual conversation while eating together.
-Unlike Hogwarts’ segmented house-based accommodation, this dining space was open and inclusive, enabling cross-social group connections.
-Others described public third places supporting casual, serendipitous interactions.
+Drawing on Club Penguin (club-penguin) and Poptropica (FeigenbaumOtherPoptropicaGamesb), they imagined a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities.
+P09 imagined a “virtual dining room” for casual conversation while eating—open and inclusive, enabling cross-social group connections unlike Hogwarts’ segmented house-based accommodation.
+Public third places would also support serendipitous interactions.
 P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
-To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends and trusted communities.
+To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends.
 P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
 We didn’t need Instagram shopping… We don’t need extra things in an app that is only made for one thing—connecting… We need to figure out how to put boundaries in place between connecting with the people we love and connecting with people we don’t know yet… you don’t need to have so many political takers on this one app… It’s just a very stressful situation.” (P19)
-Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference for close connections: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH being “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
+Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” SMH would be “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
-Participants described how themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
+Themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal than MSM.
 “Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread.
@@
 I find it’s way easier to get to know people a lot better.” (P22)
-Participants described how conversational space design could make SMH interactions feel safer, more personal, and more inviting than existing platforms.
 Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences.
-P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations.
-They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
-Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected.
-P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate.
-P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n.
-d.]) where users could create custom rooms on demand to match the “purpose” of their conversation.
-Participants also valued interest-based spaces, similar to Discord (discord page, [n.
-d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
+P22 imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences.
+These would be “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting with MSM’s more exposed feel where “you don’t know everything around it.”
+Personalizing these spaces would make them more emotionally connected.
+P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories.” P09 imagined a “Room of Requirement”-like system (room-of-requirement) where users could create custom rooms on demand to match the “purpose” of their conversation.
+Interest-based spaces, similar to Discord (discord), would provide a “common point” (P13) or “good talking point” (P07) to break the ice.
 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
-To ensure SMH feels safe and personal, participants stressed needing comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools.
-P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (Huang and Vitak, 2022)) to people they wanted to avoid, based on mutual follows.
-They stressed needing stronger, more thorough blocking controls, including blocking “alternate accounts” and fully removing blocked person traces, explaining “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (Bojkov, 2019).
-Participants wanted stronger control over their visibility, imagining features letting them opt out of being seen or discovered entirely.
+SMH would need comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools.
+P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (VitakHuang-2022-FinstaAccounts-t)) to people they wanted to avoid, based on mutual follows.
+They stressed needing stronger blocking controls, including blocking “alternate accounts” and fully removing blocked person traces: “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (story-highlights).
+Participants wanted stronger control over visibility, imagining features letting them opt out of being seen entirely.
 P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
@@
 P22 proposed a Discord-inspired model where only people you’ve interacted with can contact you, while P13 argued people “should only be able to add you if you meet in person and exchange a QR code,” reflecting that “it’s too easy to find people on social media.”
-4.4. Nuanced Self-Presentation
+4.4. Theme 4: Nuanced Self-Presentation
 Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14).
 This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23).
-P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting their full range without curation or performance pressures.
+P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” SMH, by contrast, would allow self-presentation that feels more personal, authentic, and flexible—reflecting users’ full range without curation or performance pressures.
 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
-Participants consistently described customizable avatars and personal spaces as meaningful self-expression tools, offering far more creativity than MSM’s limited settings.
-P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
+Customizable avatars and personal spaces offered meaningful self-expression far beyond MSM’s limited settings.
+P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 wanted to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
 Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality.
-P22 reflected on early Myspace (mys, [n.
-d.]), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames were ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display.
-P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right?
+P22 reflected on early Myspace (myspace), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames could “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display.
+P17 added that even unconscious choices reveal something personal: “your brain conjured that up, right?
 So I think that’s part of you.”
-Beyond visual customization, participants imagined SMH supporting ambient, indirect expression, offering subtle cues about what someone is doing, feeling, or thinking without formal posts.
+Beyond visual customization, SMH would support ambient, indirect expression—subtle cues about what someone is doing, feeling, or thinking without formal posts.
 P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh.
 that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
-Several participants highlighted the contrast between 2D platforms’ flat, compressed feel and 3D environments’ immersive qualities.
+The contrast between 2D platforms’ flat feel and 3D environments’ immersive qualities mattered.
 P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
-Participants described how these choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests.
-As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation.
+These choices feel more authentic than generic posts because they emerge naturally from personal preferences.
+As P01 put it, such “creative expression” feels “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without direct conversation.
 P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
-Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
+Personal spaces like bedrooms carry intimacy and authenticity that translates well to SMH. P13 explained:
 “A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.
@@
 It’s easier than physically changing a space.” (P13)
-Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes.
-P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention.
-They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.
+These ambient, low-pressure forms of expression also carried lower social stakes.
+P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling performative.
+This kind of expression felt “more justified” and less “embarrassing” than posting content just to be seen.
 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
-Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships while reducing pressure to maintain a single, polished identity.
-Rather than being limited to one polished identity, as on MSM, participants wanted ways to represent their full range of interests, identities, and relationships.
-P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things.
-You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are.
-P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
+SMH would support flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships without pressure to maintain a single, polished identity.
+P01 described having a house with rooms “dedicated to each of your interests”: “I’m into multiple things.
+You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” Representing different facets would reduce misjudgment and give others a more complete picture.
+P02 imagined users with multiple “Muggle personalities” organizing them into separate spaces: “Each personality has its own vault.”
 4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
-Participants valued sharing casual, everyday moments and rich, embodied memories, filling gaps between MSM’s typical highlights to create a more natural, meaningful extension of daily life and personal connection.
-P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment.
-For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
-Participants imagined SMH supporting personal, embodied memory-sharing to deepen understanding and reduce misinterpretation.
-P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n.
-d.]) as a way to create stronger emotional connections.
-Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully.
-P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
-Some participants imagined these shared memories including sensory or emotional details beyond visuals.
-P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication.
-Others compared it to live streaming personal experiences.
-P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record.
-P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
+Participants valued sharing casual, everyday moments and rich, embodied memories—filling gaps between MSM’s typical highlights to create a more natural extension of daily life.
+P04 wanted to share small, ordinary moments, like “what I’m doing at the moment.
+For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
+SMH could also support personal, embodied memory-sharing to deepen understanding and reduce misinterpretation.
+P04 wanted to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pensieve).
+Unlike MSM, where communication often feels like “just words,” (P03) memory-sharing could communicate context, emotion, and intention more fully.
+P02 explained this could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
+Some imagined shared memories including sensory or emotional details beyond visuals.
+P01 expressed interest in being able to “transmit smells through social media.” Others compared it to live streaming personal experiences—P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness for capturing fleeting moments, like crossing a state line, without pausing to record.
+P03 imagined instantly sharing moments with loved ones “in the moment,” rather than explaining them after the fact when the emotion has passed.
 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
-Participants emphasized that flexible, contextual privacy is essential for authentic self-presentation, allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented.
-They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life.
-P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
-Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces.
+Flexible, contextual privacy is essential for authentic self-presentation—allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented.
+Current platforms offer rigid, limited controls that fail to account for real social life’s complexity.
+P16 explained that on Instagram, privacy options are reduced to features like “close friends”, adding, “it’s kind of a pain to organize.” P11 noted that as people share “more like their genuine real life and their experiences,” privacy needs become “a little bit more complicated,” requiring finer-grained visibility control.
+SMH could support more seamless transitions between identities and spaces.
 P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts.
-Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation.
+Users could shift how they present themselves based on audience or situation.
 P03 proposed a model where everyone maintains a “base” identity but customizes additional layers.
-P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
-Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors.
-P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time.
-As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on.
+P22 described having different avatars for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
+Contextual privacy would extend beyond people-based access to include temporal and relational factors.
+P14 proposed limiting access to different stages of one’s past identity: “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” P21 suggested gradual openness, where people could selectively reveal more over time: “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on.
 Or turn it off.”
-Participants also highlighted the importance of individual control over these transitions and boundaries.
-P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
-4.5. Intentional Engagement
-Participants described growing frustration with current social media platforms prioritizing passive consumption over meaningful connection.
-In response, they envisioned SMH as a space promoting intentional, socially rewarding interactions.
-This section explores themes reflecting their desire to refocus social media on authentic connection, purposeful engagement, and meaningful time use.
+Individual control over these transitions mattered.
+P14 explained that privacy settings should not be one-size-fits-all: “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy not just as a technical feature, but as a core requirement for the personal, meaningful interactions youth hoped SMH would enable.
+4.5. Theme 5: Intentional Engagement
+Participants described growing frustration with current social media prioritizing passive consumption over meaningful connection.
+SMH would promote intentional, socially rewarding interactions—refocusing social media on authentic connection, purposeful engagement, and meaningful time use.
 4.5.1. Encouraging Deliberate Navigation and Content Discovery
-Participants imagined SMH encouraging deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with.
-P22 described frustration with Instagram, explaining “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage intentional, purposeful engagement when browsing content and navigating the virtual world.
-P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful.
-Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
-Participants imagined features prompting users to actively choose what they wanted to engage with before entering a space.
-P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
+SMH would encourage deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with.
+P22 described frustration with Instagram: “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.”
+P04 described how slowing down to explore someone’s space or story could make interactions feel more personal.
+Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than “taking a quick second to look at it… and then being done.”
+Some imagined features prompting users to actively choose what they wanted before entering a space.
+P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces surfacing content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
 Physical movement through SMH’s world was key to this intentionality.
-P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
+P03 explained that embodied movement makes SMH feel “more connected than regular social media” because it requires active participation rather than passive scrolling.
 P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions.
-I want everything to be very thought through.” While convenient features existed, participants valued walking and exploring, turning simple actions into shared social moments.
-P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them.
-They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it.
-It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
-P05 believed this effortful navigation would naturally discourage superficial engagement: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
+I want everything to be very thought through.”
+P05 imagined SMH as a world where you might “walk through a town to get there”: “It’s not challenging to navigate, but you have to intentionally go there to see it.
+It isn’t something you can do mindlessly.” P14 agreed that visiting someone’s space in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
+This effortful navigation would naturally discourage superficial engagement.
+P05 noted: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
 Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection.
-They expressed frustration with current platforms promoting mindless, passive consumption and hoped SMH would foster meaningful, intentional connections through active social engagement.
-They described frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled.
-P14 captured this sentiment, describing spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
+They expressed frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled.
+P14 captured this, describing time spent on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
 MSM’s infinite scrolling design contributed to this feeling.
@@
 P14 critiqued TikTok for exposing users to “clickbaity” low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
-In contrast, participants described active, social engagement as far more rewarding.
+Active, social engagement felt far more rewarding.
 P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
@@
 P10 explained that while doomscrolling feels unproductive, activities like Minecraft offer “progress” that can be “saved and come back to.” P06 noted they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
-Reflecting on SMH’s potential, P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference.
-They explained SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting meaningful engagement, rather than passive consumption, ultimately mattered.
-4.6. Playfulness and Whimsy
+P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference.
+SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content”—meaningful engagement, rather than passive consumption, ultimately mattered.
+4.6. Theme 6: Playfulness and Whimsy
 Participants described SMH as a platform feeling youthful, playful, and deeply immersive, offering entertainment plus opportunities for creative self-expression, social bonding, and emotional well-being.
@@
 P16 imagined that users could take part in activities they might never otherwise experience, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
-4.7. Expanding Horizons
+4.7. Synthesis: SMH as a Hopeful Reframing of Social Media
 Given the speculative FI co-design workshop, SMH embodies youth aspirations and values for social media platform interactions.
@@
 5. Discussion
-Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
-[…]
+This paper starts from the premise that the central challenge in reimagining social media is not a lack of ideas, but a lack of conditions that allow different ideas to emerge.
+Asking young people how to redesign “social media” is itself a loaded prompt: the term carries accumulated controversy, cynicism, and tightly constrained expectations about what platforms can and should be.
+As our findings show, this framing activates design fixation rather than imagination.
+FI offers an alternative starting point by suspending the assumptions that quietly govern what feels reasonable, feasible, or even worth proposing.
+FI does not magically produce ideas that have never existed.
+Instead, it makes different configurations of priorities and interaction logic available to think with, and legible in participants’ own terms.
+5.1. Defamiliarization, Not Novelty: How FI Expands the Frontier of Social Media Design
+A useful way to interpret participants’ concepts is through defamiliarization: making strange what has been naturalized so we can notice the assumptions embedded in today’s dominant interaction grammar.
+Mainstream social media has converged on a paradigm organized around mobile-first interfaces, algorithmically curated feeds, frictionless consumption, and quantifiable performance.
+Under this convergence, many alternative logics are no longer legible as futures, even if their ingredients have existed before.
+Customizable profiles, spatial exploration, interest-based rooms, ambient co-presence, and playful identity work are not novel components.
+What is striking is how systematically they have been treated as side paths rather than viable organizing principles, crowded out by the priorities that dominate at scale.
+“Newness” here should not mean inventing unprecedented features.
+Many elements participants reach for are historically familiar: MySpace made profile customization central; Second Life organized sociality through spatial exploration; chat rooms and MUDs normalized interest-based rooms and ambient co-presence.
+The problem is not that these ingredients never existed, but that they have been systematically crowded out by a particular configuration that won: the mobile-optimized, algorithmically-curated, vertically-scrolling feed oriented toward frictionless consumption and legible performance.
+The analytic question is less “has anyone ever built X?” than “why did X lose, and what does that loss reveal about what has become hard to imagine as a viable organizing principle?” When prior platforms are invoked today, they are often treated as feature catalogs or nostalgic references, rather than as prompts to reconsider the interaction logics and priorities that once made those logics coherent.
+Seen this way, FI is not a machine for generating ideas no one has ever had; it is a method for making alternative configurations thinkable again.
+Even if every individual component has appeared before, specific arrangements matter enormously: the smartphone, touchscreen, and app store all existed before the iPhone, but a different bundling of affordances and defaults made a new paradigm visible.
+Participants’ proposals are revealing because they reorder emphasis: what is made reversible versus permanent, what requires consent, what counts as presence, and what the system treats as the default social unit.
+This reframing also helps explain why many youth examples draw from spaces not labeled “social media” at all: Discord servers, Minecraft worlds, and game-like environments where connection is organized around shared activity, rooms, and mutual orientation rather than broadcast audiences.
+FI loosens the category boundary itself, expanding what futures feel available to propose, including trajectories previously abandoned for reasons of incentive structures, scale narratives, or technical constraints rather than because their social ideas were exhausted.
+FI matters because it interrupts the subtle constraints that ordinarily shape what participants feel permitted to propose.
+The fictional frame suspends economic, technical, and reputational pressures, but more importantly it removes the category label “social media,” which often triggers cynicism and incremental repair.
+In its place, FI invites reasoning from felt social experience: what presence should feel like, what kinds of disclosure should be low-stakes, what boundaries should be socially intelligible, and what kinds of connection should be effortless versus effortful.
+That shift makes it easier to surface configuration-level commitments: not whether a platform has a “profile” or “messaging,” but what counts as visibility, what is reversible, what requires consent, and what the system treats as the default social unit (an audience, a network, a room, a neighborhood, a party).
+Figure 1. Youth’s Understanding of Social Media Interactions
 A Venn diagram comparing two categories of online platforms.
@@
 In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
-Our analysis shows that youth imagine future social media not as a rejection of current platforms but as a reconfiguration of their strengths.
-They want the visibility and reach of social networking and broadcasting platforms while also craving the intimacy and authenticity found in spaces like games, voice chat, or small-group apps.
-Figure 2 illustrates this blended vision: an aspirational middle ground where meaningful connection is not an afterthought but the central purpose.
-Building on this insight, the following subsections trace how participants moved from critiquing existing platforms toward articulating design goals, culminating in our proposal of “spatial integrity” as a framework for supporting deeper forms of online connection.
-The relevance of these findings has become increasingly apparent with recent technological developments.
-With the introduction of technologies like Genie 3 and similar platforms, the fictional world medium of remote social connection that participants envisioned is barely fictional.
-The 3D, open-world, and fully customizable platforms for self-expression and organic social connection that emerged from our workshops now represent concrete possibilities for the near future, much nearer than we anticipated when we began this study.
-This technological trajectory underscores the timeliness of our findings and their potential to guide the development of relationship-supporting technologies that leverage these emerging capabilities (DeepMind, 2025).
-5.1. From Reactionary Designs to Aspirational Design Goals
-There are societal concerns about social media’s impact on youth well-being (Oth, 2023).
-Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023).
-This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
-However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality.
-Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023).
-At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024).
-This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connections that youth actually seek.
-Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
-However, in our earlier work with this youth population, we noticed several recurring challenges that shaped how participants thought about social media.
-First, there was a strong design fixation on social media design around existing apps, both in terms of interface, such as small 2D mobile screens, endless scrolling, and uniform templates, and in terms of user experience, where visibility depends on maintaining large networks of friends and followers, often at the expense of authenticity and comfort.
-Second, participants showed a cognitive fixation when imagining the future of social media, struggling to see it as anything beyond its currently perceived role as a convenient but limited tool for remote social connection.
-Third, these fixations were compounded by generally negative experiences with social media, which led many in earlier interviews to propose reactionary fixes, like stronger privacy, the removal of infinite scrolls, or eliminating general media content such as Reels.
-Such cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
-To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting.
-The procedure was intended to serve three purposes: breaking participants free from these biases, learned helplessness, and pessimism by encouraging them to think beyond the “app” or “platform” on mobile devices; equipping them with prompts that licensed thinking beyond perceived technical and practical constraints; and creating conditions where they could explore proactive and aspirational visions of social interaction rather than simply reacting to existing frustrations.
-Our chosen method prioritized uncovering underlying desires and needs that young people associate with meaningful remote social interaction over producing concrete feature lists.
-In line with the goals of co-design processes (Sanders and Stappers, 2008), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most.
-Particularly noteworthy was the fact that the workshops helped participants expand their understanding of what social platforms could be, empowering them to feel more hopeful and articulate desires they might have previously dismissed as impossible within conventional social media paradigms.
-This demonstrates how the design method chosen in our study offered a generative path toward hope in domains where narratives of doom are prevalent.
-Our analysis contributes insights about the types of remote social interactions that youth find most meaningful and fulfilling, examining how these desired interactions align with or diverge from current platform affordances, as we detail in Section 5.2.
-We also identify core aspects of design that can undermine relational quality from which we develop the “spatial integrity” framework introduced in Section 5.3.
-Taken together, we hope this work equips designers with proactive design goals for envisioning new futures of social technologies and provides an initial framework for evaluating design integrity in relationship-supporting aspects.
-5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
-Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online.
-Instead, they seek spaces that support deeper, more intentional social interactions.
-Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative.
-The changes in social media platform design that they sought were as follows:
-  1. (1)
-    Immersive Experiences Through Spatial Elements.
-Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences.
-They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed.
-They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied.
-They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures.
-Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
-  2. (2)
-    Organic and Intentional Social Interactions.
-Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions.
-They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally.
-They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out.
-Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people.
-Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
-  3. (3)
-    Expressive and Lower-Stakes Sharing.
-Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts.
-They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars.
-They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection.
-Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
-  4. (4)
-    Granular, Intuitive, and Playful Privacy Mechanisms.
-Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid.
-They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries.
-They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable.
-Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments.
-They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
-  5. (5)
-    Refocusing Interaction Priorities for Meaningful Engagement.
-Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about.
-They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams.
-They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections.
-Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape.
-Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
-These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams.
-Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement.
-While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
-Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions.
-This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
-5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
-Concept Description Supported Interactions
-Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence.
-Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation.
-Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
-Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction.
-Platforms like Discord create stable communal spaces with clear channels and streams.
-Users value this for fostering deeper relationships.
-Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
-Spatial Composition The segmentation of a social network into defined areas and pathways.
-Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces.
-Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
-Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting.
-3D social media limits expression to explicit actions like posting and commenting.
-Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
-Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
-Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
-  1. (1)
-    Spatial Presence: Understanding the Space and Who’s There.
-Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others).
-While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
-  2. (2)
-    Spatial Composition: The Meaning of the Space.
-Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019).
-Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams.
-Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
-  3. (3)
-    Spatial Configuration: How the Space is Organized.
-Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019).
-Clear boundaries and pathways help people recognize territories and manage social norms.
-Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness.
-This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
-  4. (4)
-    Spatial Depth: How Much Information Can Be Perceived.
-Spatial depth determines how much information users can perceive in a space.
-3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001).
-Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression.
-Incorporating spatial depth could enable richer self-expression and more natural interactions.
-According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire.
-The relationship between these dimensions and the interactions they support is laid out in Table 2.
-These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections.
-By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish.
-While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram.
-Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms.
-(It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
-5.4. Limitation and Future Work
-Our study has several key limitations to consider.
-First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently transfer to youth who are not familiar with such environments.
-Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations.
-Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance.
-This suggests opportunities to refine the narrative approach.
-Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces.
-This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
-Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant.
-They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends.
-Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share.
-This is particularly striking given that participants were otherwise skeptical of AI.
-This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative.
-It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would.
-Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
+5.2. Minecraft, Not Instagram: Interaction Logics Beyond the Feed
+Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction.
+This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways they happen to be designed today (OtherOtherTikTokHarmst; noauthor_2023-pn).
+Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections.
+Yet this characterization did not match where they described actually feeling close to friends.
+As illustrated in Figure 1(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections (such as Discord or Minecraft).
+The latter are not typically framed as “social media” in everyday discourse, even though they often serve social functions more directly.
+This boundary is not merely semantic: it shapes what becomes imaginable.
+If “social media” is assumed to mean a 2D feed, a profile, and an attention economy, then even well-intentioned redesign prompts tend to elicit cynicism or incrementalism.
+When participants reasoned from spaces they already associate with social closeness (games, chat servers, shared virtual worlds), the future of social connection expanded beyond the feed metaphor.
+This misalignment between expectations and experiences is problematic because negative perceptions of social media use can diminish psychological well-being (lee2024social; kim2024privacysocialnormsystematically).
+Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce.
+As shown in Figure 1(b), social media serves diverse uses and gratifications, from information seeking and entertainment to status enhancement and network maintenance.
+Rather than maintaining an entirely negative view confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations and to name the kinds of connection each one is organized to support.
+Critiquing Instagram for not facilitating authentic, deep friendships can miss the point: while such connections might occasionally form there, it is not the platform’s central form of engagement.
+However, platforms are not static; their designs evolve, and user engagement patterns shift over time.
+Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts.
+This fluidity underscores the importance of teaching (WilsonWisniewski-2012-FightingMyRegulation-m; Wisniewski-2018-PrivacyParadox-l; TheLearningNetwork-2025-WhatTeensMedia-z) youth to be more discerning about the connections they seek and what a given platform’s interaction model realistically supports, rather than reinforcing anxiety-inducing narratives that treat “social media” as a single, uniformly harmful category.
+Together, these accounts clarify the contrast participants repeatedly drew between platforms they associate with “social media” and the spaces where they actually feel socially connected.
+When describing meaningful connection, they emphasized presence, shared activity, and low-pressure co-presence, drawing on experiences in Minecraft- and Discord-like environments rather than Instagram-style broadcast and feed consumption.
+5.3. From Subjects to Authors: FI as a Translator for Youth Voice
+Youth are typically positioned in social media discourse as subjects to be protected or problems to be solved: screen time to be managed, mental health to be safeguarded, misinformation vulnerability to be addressed.
+This framing, however well-intentioned, reproduces a power imbalance in which young people live with the consequences of platform design earlier and more intensely than any other demographic, yet rarely get to author the underlying rules.
+FI offers a corrective.
+By treating participants as theorists and inventors rather than survey respondents or focus group informants, FI shifts the locus of imagination.
+The result is not merely a collection of wishes, but a disciplined translation pipeline that surfaces interaction primitives (what actions exist, what counts as visibility, what is reversible, what requires consent), normative frames (what should be easy versus hard, what should be private by default, what should be communal), and actionable design directions grounded in participants’ own social experience.
+This repositioning matters because asking young people to imagine within realism often reproduces the very constraints that limit design thinking.
+Realism already encodes corporate incentives, technical path dependencies, and fear-based narratives inherited from parents, teachers, and policymakers.
+When participants are prompted to “fix” social media, they inherit a frame saturated with cynicism and incrementalism.
+FI sidesteps this inheritance by suspending the category label “social media” altogether.
+Crucially, the fictional frame does more than remove constraints; it makes certain desires and boundaries socially permissible to articulate.
+Wishes that might feel naive, awkward, or excessively earnest under a realistic prompt—wanting connection to feel like being in the same room, wanting visibility to be reversible, wanting strangers to remain strangers unless actively invited—become speakable when the task is to describe a magical world rather than repair an existing platform.
+In place of defensiveness, participants reason from felt social experience: what presence should feel like, what kinds of disclosure should be low-stakes, what boundaries should be socially intelligible.
+A practical consequence is that FI generates common language between youth and adults.
+Much of the current tension in public discourse around social media stems from a mismatch in vocabulary: adults speak in terms of harm, addiction, and regulation, while youth often lack the conceptual tools to articulate what they want connection to feel like.
+FI bridges this gap by making youth priorities legible in terms that designers, researchers, and policymakers can act upon.
+When participants describe wanting “rooms” instead of “feeds,” or “neighborhoods” instead of “audiences,” they are not simply expressing preference for a feature; they are naming a different interaction grammar.
+That grammar, once articulated, becomes available for translation into design specifications, policy frameworks, and educational curricula.
+The contribution is not a single new feature, but evidence that a different organizing logic is possible and legible to users when the design frame allows it.
+Finally, FI cultivates something often missing from conversations about social media: hope.
+The shift we observed from “I do not know what else there could be” to expansive, imaginative visions is not incidental.
+It reflects what happens when participants are given back authorship over the future, even temporarily.
+Past platforms were designed by adults with particular business models and assumptions baked in; youth imagining social worlds from scratch surfaces different priorities.
+The deeper problem is not that no one has tried to improve social media, but that most attempts operate within the same paradigm: adjusting privacy settings, adding screen time limits, tweaking algorithmic recommendations.
+These are repairs to a house whose floor plan is assumed fixed.
+FI invites participants to redraw the floor plan itself, supporting not only imagination but critical capacity: when youth can name what they want connection to feel like, and can describe the rules that would make that possible, they are better positioned to evaluate platforms critically without being trapped in cynicism.
+FI supports youth voice not only by eliciting ideas, but by cultivating the sense that different interaction futures are thinkable and therefore arguable.
+5.4. Limitations and Future Work
+This study is exploratory, and the design orientations we synthesize should be understood as generative rather than generalizable.
+Our participant demographics skewed toward women, and future work should examine how robust these orientations are across broader samples and cultural contexts, including youth with different relationships to gaming, spatial interfaces, and platform norms.
+Also, FI intentionally suspends feasibility and market constraints, which is a strength for defamiliarization but a limitation for direct translation into deployable systems.
+Some of the most compelling interaction logics we surfaced (spatially legible boundaries, ambient co-presence, rich memory-sharing) raise practical questions about accessibility, moderation, harassment, surveillance, and inequities in device access.
+Future work should prototype selectively, using low-fidelity and mid-fidelity implementations to evaluate whether the same felt qualities can be achieved under real constraints, and to identify new risks that emerge when imagined systems become operational.
 6. Conclusion
-This study explored how youth with 3D gaming experience reimagine social media when freed from current platform constraints.
-Through Hogwarts’ fictional world, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption defining mainstream social media.
-Their visions called for platforms that feel more like 3D games than Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to real relationships’ complexity while incorporating playfulness and whimsy.
-Building on these insights, we introduced spatial integrity, foundational design principles enabling such visions.
-While spatial integrity alone cannot guarantee meaningful connection, it provides structural conditions supporting the relational, embodied, and playful interactions participants described.
-The study underscores that youth seek not less time on social media, but more fulfilling and emotionally present interactions within digital spaces.
-Their visions point toward a future where social media feels like the dynamic, shared social worlds found in 3D gaming environments.
-While our findings stem from youth familiar with 3D gaming, they offer a hopeful design direction for spaces supporting the meaningful social connections youth desire.
-Appendix A FI Workshop Protocol
+This paper demonstrates that when youth are invited to imagine social connection outside the conceptual frame of “social media,” they articulate desires and interaction logics that diverge sharply from dominant platform paradigms.
+Through FI workshops with 23 participants aged 15–24, we surfaced six themes—intuitive social navigation, shared experiences, communal ambience, nuanced self-presentation, intentional engagement, and playfulness—that collectively describe what meaningful remote connection might feel like if it were organized around presence, place, and relationships rather than feeds, metrics, and passive consumption.
+The findings carry implications for both research and practice.
+For researchers, this study offers evidence that design fixation in social media is partly a methodological problem: the term “social media” itself activates assumptions that constrain imagination.
+FI provides one way to interrupt those assumptions, enabling participants to reason from felt social experience rather than inherited platform templates.
+For designers and policymakers, the themes we identify suggest that the relational shortcomings of mainstream platforms are not inevitable features of mediated connection but outcomes of specific design choices—choices that could be made differently.
+Appendix A Fictional Inquiry Materials
+A.1. Miro Board for Fictional Inquiry Warm-Up
+Figure 2. Miro board mapping real-life relationships to Hogwarts-themed groups, designed to immerse participants in the Hogwarts-inspired FI narrative.
+This is a visual chart designed to organize relationships and connections in the Harry Potter universe.
+There are three columns of labeled sections.
+In the middle, there are four sections, each representing one of the Hogwarts houses: Gryffindor (red), Hufflepuff (yellow), Slytherin (green), and Ravenclaw (blue).
+Each house box has its crest.
+To the left, there are two smaller boxes.
+The top-left box is labeled “Best Friends / Romantic Relationship(s),” featuring small images of Harry Potter, Hermione Granger, and Ron Weasley, with a space for a name or nickname.
+Below this, there is a smaller box labeled “House Elf,” featuring a picture of Dobby holding a glowing object and an area for a name.
+To the right, there are two more boxes.
+The top-right box is labeld “Family,” containing images of Sirius Black, James and Lily Potter, and the Dursley family.
+Below this, another box is labeld “Profs & Other Adults,” displaying images of Severus Snape, Dumbledore, Professor McGonagall, and Hagrid.
+At the bottom, there are three large sections.
+The leftmost section is labeled “Muggles” with a blank space.
+The middle section, labeld “Students from Other magical Schools,” displays crests from Beauxbatons, Durmstrang, Ilvermorny, and Castelobruxo.
+The rightmost section, labeled “Quidditch Team Members,” features an image of a Quidditch team in red robes.
+Each section contains an empty space for names or details to be added.
+Figure 2 shows the Miro board template used during Fictional Inquiry sessions.
+Participants sorted their friends and family into Potter-inspired categories (houses, teams, old friends) to encourage thinking about relationships within the fictional context and enhance immersion in the design space.
+This activity preceded the main design exploration and helped participants transition from discussing their current social media experiences to imagining alternatives within the Hogwarts setting.
+Appendix B FI Workshop Protocol
 Introduction
@@
 Thank you so much for your time and insights. Your contributions are incredibly valuable to our research. Do you have any final thoughts or questions before we conclude?
-Appendix B Participant Interview Summary
-Table 3. Key Features for Envisioned Social Media from Participant Interviews
-PID Envisioned Social Media Features
-P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
-P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
-P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
-P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
-P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
-P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
-P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
-P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
-P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
-P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
-P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
-P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
-P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
-P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
-P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
-P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
-P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
-P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
-P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
-P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
-P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
-P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
-P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.